### v1

# The analysis of paired comparison data in the presence of cyclicality and intransitivity

Abstract
A principled approach to cyclicality and intransitivity in cardinal paired comparison data is developed within the framework of graphical linear models. Fundamental to our developments is a detailed understanding and study of the parameter space which accommodates cyclicality and intransitivity. In particular, the relationships between the reduced, completely transitive model, the full, not necessarily transitive model, and all manner of intermediate models are explored for both complete and incomplete paired comparison graphs. It is shown that identifying cyclicality and intransitivity reduces to a model selection problem and a new method for model selection employing geometrical insights, unique to the problem at hand, is proposed. The large sample properties of the estimators as well as guarantees on the selected model are provided. It is thus shown that in large samples all cyclicalities and intransitivities can be identified. The method is exemplified using simulations and the analysis of an illustrative example.
Key-Words: Cyclicality, Goodness of fit, Intransitivity, Large sample properties, Least square ranking, Model selection.
1 Introduction
There are many situations in which ranking a set of items is desired. Examples include the evaluation of political candidates, marketing, sports, information retrieval, and a variety of modern internet and e-commerce applications. There is a large literature on the theory, methodology, and application of ranking methods from a varied set of perspectives by researchers in the fields of mathematics (Langville and Mayer 2012), social choice theory (Sen 1986, Slutzki and Volij 2005), machine learning (Ailon et al. 2008, Furnkranz and Hullermeier, 2010), psychology (Regenwetter et al 2011, Davis-Stober et al. 2016) as well as many other disciplines including statistics, e.g., the books by Marden (1995) and Alvo and Yu (2014). It is important to emphasize that a global ranking exists only if preferences are transitive (Vriend, 1996). It is well known that the axioms of rationality in economics imply transitivity. It has been argued that a violation of these axioms would lead to a so–called “money–pump”, cf., Gustafsson (2022). Nevertheless, there is no consensus on this issue and numerous authors have advocated for relaxations of transitivity, e.g., Fishburn and LaValle (1988), Schmeidler (1989), Sen (1997) and Klimenko (2015). These papers focus on individual preferences and thus may have limited applicability to the problem at hand. Empirical aspects of intransitivity will be discussed shortly.
A ranking can be constructed using various data types such as ratings and ranked lists. In particular, paired comparison data (PCD) is obtained if all comparisons involve only two items (David, 1988). To fix ideas suppose that there are $K$ items labelled $1,\ldots,K$ which we would like to rank. Let $Y_{ijk}$ denote the outcome of the $k^{th}$ comparison among items $i$ and $j$. The random variable (RV) $Y_{ijk}$ may be binary, ordinal or cardinal. In this paper we focus on cardinal, i.e., continuous, paired comparison data (PCD). However, we emphasize that our key insights and results can be readily adapted to binary and ordinal PCD as outlined in Section 7. Concretely, we shall assume that the observations $Y_{ijk}$ for $1\leq i\neq j\leq K$ and $k=1,\ldots,n_{ij}$ satisfy
$Y_{ijk}=\nu_{ij}+\epsilon_{ijk},$ (1)
where $\nu_{ij}=\mathbb{E}(Y_{ijk})$ and the errors $\epsilon_{ijk}$ are independent zero mean RVs. Since $Y_{ijk}=-Y_{jik}$ we also have $\nu_{ij}=-\nu_{ji}$. Thus model (1) is indexed by a parameter $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ where $\operatorname{\boldsymbol{\nu}}=(\nu_{12},\ldots,\nu_{1K},\nu_{23},\ldots,\nu_{2K},\ldots,\nu_{K-1,K})\in\mathbb{R}^{K(K-1)/2}$ is a vector of means arranged in a lexicographical order. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote the graph whose vertices are the items $1,\ldots,K$ and edges are the pairs $(i,j)$ for which $n_{ij}>1$. With each edge $(i,j)$ we associate a random sample $\mathcal{Y}_{ij}=(Y_{ij1},\ldots,Y_{ijn_{ij}})$ of $n_{ij}$ comparisons; the set of all samples is denoted by $\mathcal{Y}$. The pair
$(\mathcal{G},\mathcal{Y})$ (2)
is referred to as pairwise comparison graph (PCG).
We write $i\succeq j$ when item $i$ is preferred over item $j$, which in the context of (1) is equivalent to $\nu_{ij}\geq 0$. The magnitude of $\nu_{ij}$ measures the degree of preference of item $i$ over item $j$. A strict preference, i.e., $i\succ j$ is associated with a strict inequality $\nu_{ij}>0$. If, for example, $1\succeq 2\succeq\cdots\succeq K$ then all items are uniquely ranked and it follows that $\nu_{ij}\geq 0$ whenever $i<j$. Such an ordering is called a total order or a global ranking. If $i\succeq j$ and $j\succeq k$ implies that $i\succeq k$ then the preference relation between the items $i,j$ and $k$ is transitive. There are many types of transitivity relations, such as weak stochastic transitivity, strong stochastic transitivity and the even stronger linear transitive order also referred to as the total linear order, cf. Fishburn (1973),Oliveira et al. (2018), and the references therein. The total linear order specifies that
$\nu_{ij}=\mu_{i}-\mu_{j}$ (3)
for all $1\leq i\neq j\leq K$ where $\mu_{1},\ldots,\mu_{K}$ are known as the scores or merits. Clearly $i\succeq j$ whenever $\mu_{i}\geq\mu_{j}$. Moreover the vector of scores completely defines the ranking among all items. When the restriction (3) holds
$\operatorname{\boldsymbol{\nu}}^{\top}=(\mu_{1}-\mu_{2},\ldots\mu_{1}-\mu_{K},\mu_{2}-\mu_{3},\ldots,\mu_{2}-\mu_{K},\ldots,\mu_{K-1}-\mu_{K})$ (4)
for some $\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}$. We denote the collection of vectors satisfying (4) by $\mathcal{M}$. Naturally, $\mathcal{M}\subset\mathcal{N}$. Note that the parameterization (3) results in a non–identifiable model. Therefore it is common to assume that the vector of merits satisfies a linear constraint such as $\mu_{1}+\cdots+\mu_{K}=0$. Consequently $\mathcal{M}$ can be further identified with $\mathbb{R}^{K-1}$. It is emphasized that the parametrization (3) is the common, default, parametrization found in the literature. Further note that the least square estimators (LSEs) for model (1) assuming (3) have a long history and have been studied in diverse fields (cf. Mosteller 1951, Kwiesielewicz 1996, Csato 2015). For a recent comprehensive account see Singh et al. (2024).
Transitivity is a fundamental choice axiom underlying decision theory and its numerous applications. Nevertheless, it is well known that not all preference relations are transitive. For example it possible that among three items, $\{i,j,k\}$ say, we may have a cyclical preference relation $i\succeq j\succeq k\succeq i$. This situation arises in some simple games such as rock–paper–scissors. More importantly, as indicated by the survey below, intransitivity and cyclicality are a part and parcel of many observed real–world phenomena. For example, intransitivity has been extensively explored in the literature on social–choice theory. Condorcet’s paradox, also known as the voting paradox, is a well known example of a cyclical, thus intransitive, preference pattern (Gehrlein 1983). Preference relations of the form
$i_{1}\succeq i_{2}\succeq\cdots\succeq i_{r}\succeq i_{1}$ (5)
where $r\geq 3$ are called cycles and they preclude a ranking among the items. There is substantial empirical evidence of intransitivity in some domains. Tversky (1969), using binary choices among monetary lotteries, demonstrated that individual choices may violate transitivity. Many papers extending and further formalizing his work have been put forth, e.g, Birnbaum and Gutierrez (2007), Birnbaum and Schmidt (2008), Regenwetter et al. (2011), Cavagnaro and Davis-Stober (2014), McCausland et al. (2020) and Alós-Ferrer and Garagnani (2021) and Alós-Ferrer et al. (2022) among many others. The aforementioned papers provide both analytical methods for testing for intransitivity and experimental data supporting it. Intransitivity has been also studied in biological systems, e.g., Baillon et al. (2015), Kerr et al. (2002) and Vandermeer and Perfecto (2023) as well as in the psychological literature, e.g., Regenwetter et al. (2011). We note that Iverson and Falmagne (1985), who reanalyzed Tversky (1969) data, argued that the evidence in that data set is compatible with transitivity. Similar criticisms directed at other authors are also known, e.g., Loomes et al. (1992). Birnbaum and Schmidt (2010) argue that one reason for these conflicting views is that researchers in the field were not able to discriminate between a systematic violation of transitivity and violations due to random sampling. We believe that these controversies, which are due in part to faulty statistical analyses, mostly stem from a lack of proper a proper framework and models for analyzing cyclical and intransitive PCD.
A comprehensive discussion of intransitivity and cyclicality in the context of paired comparison data is lacking. Csato (2019) adopts an axiomatic framework and proves some impossibility theorems. He further discusses restrictions on the preference functions which result in transitivity. Jiang et al. (2011) develop a framework for least squares ranking by viewing cardinal PCGs as edge flows on a comparison graph. Using discrete Hodge theory (Lim, 2020) they argue that edge flows, our $\operatorname{\boldsymbol{\nu}}$, can be decomposed into three orthogonal components which they refer to as the gradient flow, the curl flow and the harmonic flow. Saari (2014) and Saari (2021) develop a linear algebraic formulation and discussed an orthogonal decomposition of $\operatorname{\boldsymbol{\nu}}$ into linear and cyclical components. The aforementioned papers, which provide a conceptual foundation for reasoning about cyclicality lay outside the statistical literature. Further note that cyclicality has long been recognized in the statistical literature as well. Kendall and Smith (1940) developed a coefficient of consistency for detecting cyclicality using combinatorial counting arguments. Their statistical ideas were further developed by Moran (1947), Alway (1962), Knezek et al. (1998), Iida (2009) whereas combinatorial aspects were pursued by Bezembinder (1981), Siraj et al. (2015) and Kulkowski (2018) among many others. It is emphasized that these papers provide methods for detecting, but not for addressing, cyclicality and intransitivity. The literature on statistical models explicitly accounting of intransitivity and cyclicality is rather recent and sparse. We note the papers of Causeur and Husson (2005), Chen and Joachims (2016), Makhijani and Ugander (2019) and Spearing et al. (2023) who all developed limited parametric models for possibly intransitive paired comparison data.
The goal of this paper is to fill this gap and to fully address the vexing problems of cyclicality and intransitivity. We capitalize on Jiang et al. (2011) topological view and Saari’s (2014) and Saari (2021) linear algebraic insights, upon which we expand. We further develop a novel statistical methodology which allows for both the detection and the modelling of cyclicality and intransitivity in a unified way. In particular our main contributions are:
  1. 1.
    A detailed study of the structure of the parameter space $\mathcal{N}$ is provided. In particular an orthogonal decomposition of $\mathcal{N}$ is considered. Meaningful spanning sets for the linear and cyclical subspaces are provided yielding a general modelling framework which accommodates cyclicality and intransitivity in a unified manner. The notion of a minimal, i.e., a most parsimonious, model is introduced and studied along with some allied new concepts.
  2. 2.
    Estimators for models in the presence of cyclicality and intrasitivity are derived along with their statistical properties. Tests for lack of fit are proposed and incorporated in the process of model selection.
  3. 3.
    A novel model selection technique which builds on our geometric insights, discussed in (1) above, is proposed. A key tool is the so called tick–table introduced in Section 4.2. It is further shown that minimal models are generally identifiable and can be selected with probability tending to one. Unlike many other methods of model selection our method scales quadratically with the number of items $K$ and is therefore readily applied to large comparison graphs.
The paper is organized in the following way. In Section 2 the reduced model, i.e., $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$, and the full model model, i.e., $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ are reviewed and the relationship among them is explored. Section 3 provides a rigorous study and a decomposition of the parameter space $\mathcal{N}$ resulting in a a modelling framework which accommodates cyclicality and intransitivity. Estimators and tests of lack of fit for models accomodating cyclicality and intransitivity are developed in Section 4 in which a new model selection strategy is also proposed and studied. Simulation results are presented in Section 5 and an illustrative example is discussed in Section 6. We conclude in Section 7 with a brief summary and discussion. All proofs are collected in an Appendix.
2 Preliminary results
This section is comprised of the three parts. In Subsection 2.1 model (1) is fit assuming (3). This model, which we shall refer to as the reduced model, has been well studied in the literature and the results in this short subsection, except for the goodness of fit test, are not new. In Subsection 2.2 the model is fit without any restriction on $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$. We shall refer to this model as the full model. Finally, in Subsection 2.3 the relationship between the reduced and full model are clarified.
Remark 2.1
In this communication vectors, such as $\operatorname{\boldsymbol{\nu}}$, are often indexed by a pair of indices $(i,j)$ whereas matrices may be indexed by two pairs, i.e., $((i,j),(k,l))$ or a pair and triad, i.e., $((i,j),(k,l,m))$. Unless stated otherwise, the elements of these vectors and matrices are ordered lexicogrphically. In addition the cardinality of a set as well as the dimension of vector or a linear space will be denoted by ||$\cdot$|| ⋅ |.
2.1 Fitting and evaluating the reduced model
Suppose first that $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$, i.e., model (1) holds with (3). Consequently we are interested in estimating $\operatorname{\boldsymbol{\mu}}=(\mu_{1},\ldots,\mu_{K})^{\top}$. Following Singh et al. (2024) the least squares estimator (LSE) is defined by
$\widehat{\operatorname{\boldsymbol{\mu}}}=\arg\min\{Q(\operatorname{\boldsymbol{\mu}}):\operatorname{\boldsymbol{v}}^{\top}\operatorname{\boldsymbol{\mu}}=0\}$ (6)
where the objective function is given by
$Q(\operatorname{\boldsymbol{\mu}})=\sum_{1\leq i<j\leq K}\sum_{k=1}^{n_{ij}}(Y_{ijk}-\left(\mu_{i}-\mu_{j}\right))^{2}$
and without any loss of generality we set $\operatorname{\boldsymbol{v}}=\operatorname{\boldsymbol{1}}1=(1,\ldots,1)^{\top}$. If the graph $\mathcal{G}$ is connected then the LSE is unique and given by
$\widehat{\operatorname{\boldsymbol{\mu}}}=\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{S}}$ (7)
where $\operatorname{\boldsymbol{N}}^{+}$ is Moore–Penrose inverse of $\operatorname{\boldsymbol{N}}$, the laplacian of the graph $\mathcal{G}$ (Bapat 2010), i.e., $\operatorname{\boldsymbol{N}}$ is a $K\times K$ matrix with elements $\sum_{j}n_{ij}$ if $i=j$ and $-n_{ij}$ when $i\neq j$. Furthermore, $\operatorname{\boldsymbol{S}}=(S_{1},\ldots,S_{K})^{\top}$, is a vector of sums where $S_{i}=\sum_{j\neq i}S_{ij}$ and $S_{ij}=\sum_{k=1}^{n_{ij}}Y_{ijk}$. The total number of paired comparisons is denoted by $n=\sum_{1\leq i<j\leq K}n_{ij}$. Singh et al. (2024) provide graph–based necessary and sufficient conditions for consistency as well as asymptotic normality. These are listed here for completeness:
Condition 2.1
The errors $\epsilon_{ijk}$ are IID with zero mean and a finite variance $\sigma^{2}$.
Condition 2.2
There exists a spanning tree $\mathcal{T}\subset\mathcal{G}$ such that $\min\{n_{ij}:\left(i,j\right)\in\mathcal{T}\}\rightarrow\infty$ as $n\rightarrow\infty$.
Condition 2.3
There exists a spanning tree $\mathcal{T}\subset\mathcal{G}$ such that
$\min\{n_{ij}:(i,j)\in\mathcal{T}\}/n\to c\in(0,\infty)\text{ as }n\to\infty.$
Briefly, Condition 2.1 guarantees that the LSE has a finite variance, Condition 2.2 guarantees consistency and Condition 2.3 guarantees asymptotic normality. Under the aforementioned conditions
$\sqrt{n}(\widehat{\operatorname{\boldsymbol{\mu}}}-\operatorname{\boldsymbol{\mu}})\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\sigma^{2}\boldsymbol{\Theta}^{+}),$ (8)
where $\Rightarrow$ denotes convergence in distribution and $\boldsymbol{\Theta}=\lim_{n}\operatorname{\boldsymbol{N}}/n$ is the scaled limit of the Laplacian. Condition 2.3 guarantees that the matrices $\boldsymbol{\Theta}$ and $\boldsymbol{\Theta}^{+}$ are of rank $K-1$.
We are now ready to test whether (3) holds. One way of doing so is by employing the statistic
$R_{n}=\sum_{(i,j)\in\mathcal{E}_{n}}n_{ij}(\frac{{S}_{ij}}{n_{ij}}-(\widehat{\mu}_{i}-\widehat{\mu}_{j}))^{2},$
where $\mathcal{E}_{n}$ is the subset of $\mathcal{E}$ on which $n_{ij}/n\rightarrow c>0$. Thus, $R_{n}$ is a squared weighted residual on $(\mathcal{G},\mathcal{Y})$. Jiang et al. (2011) refers to $R_{n}$ as a certificate of reliability without reference to any formal procedure for using it. Many other measures of fit are surveyed by Brunelli (2018); none to the best of our knowledge have been analyzed from a statistical perspective.
Describing the limiting distribution of $R_{n}$ requires the following notation. First, let $\boldsymbol{\Xi}$ be the $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ diagonal matrix whose $(i.j)^{th}$ diagonal element is $\theta_{ij}=\lim{n_{ij}/n}$. Next let $\operatorname{\boldsymbol{B}}$ be a $\binom{K}{2}\times K$ matrix whose columns are $\boldsymbol{b}_{k}$, for $k=1,\ldots,K$. The $(i,j)^{th}$ element of $\boldsymbol{b}_{k}$ is denoted by $b_{k}(i,j)$ and defined by the relation
$b_{k}(i,j)=\mathbb{I}(i=k)-\mathbb{I}(j=k).$ (9)
For example if $K=4$ then $\operatorname{\boldsymbol{B}}$ is $6\times 4$ matrix given by
$\boldsymbol{B}^{\top}=\begin{pmatrix}\boldsymbol{b}_{1}^{{\top}}\\ \boldsymbol{b}_{2}^{\top}\\ \boldsymbol{b}_{3}^{\top}\\ \boldsymbol{b}_{4}^{\top}\end{pmatrix}=\begin{pmatrix}1&1&1&0&0&0\\ -1&0&0&1&1&0\\ 0&-1&0&-1&0&1\\ 0&0&-1&0&-1&-1\end{pmatrix}.$ (10)
Theorem 2.1
Assume that Conditions 2.1 and 2.3 hold. Suppose further that $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$. If so,
$R_{n}\Rightarrow\sum_{1}^{r}\lambda_{i}Z_{i}^{2}$
as $n\to\infty$, where $r=|\mathcal{E}_{n}|-(K-1)$, $Z_{1},\ldots,Z_{r}$ are independent $\mathcal{N}(0,1)$ RVs and $\lambda_{1},\ldots,\lambda_{r}$ are the non–zero eigenvalues of the $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ matrix
$\boldsymbol{\Psi}=\sigma^{2}(\boldsymbol{\Xi}^{+})^{1/2}(\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi})\,\boldsymbol{\Xi}^{+}(\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi})^{\top}(\boldsymbol{\Xi}^{+})^{1/2};$
the matrices $\operatorname{\boldsymbol{B}}$ and $\boldsymbol{\Xi}$ are defined above and $\boldsymbol{I}$ is the $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ identity matrix.
Theorem 2.1 provides a method for assessing the fit of the reduced model over all edges on which data is observed. The statistic $R_{n}$ provides a global test of fit. Local tests of fit based on $R_{n}$ can also be constructed in at least two ways: $(i)$ Fit the model, i.e., using the entire PCG but evaluate $R_{n}$ only over a subset $\mathcal{E}^{\prime}_{n}$ of $\mathcal{E}_{n}$; $(ii)$ Fit and evaluate the model over a sub-graph. The smallest such subgraph is a triangle of the form ($(\{i,j,k\},\{(i,j),(i,k,)(j,k)\})$. In both cases the resulting limiting distribution can be deduced by appropriately modifying Theorem 2.1. Details are omitted. Finally, it is also worth mentioning that if $|\mathcal{E}_{n}|=K-1$, then $R_{n}=0$ with probability one and therefore the fit of the model can not be tested. This happens if and only if the set $\mathcal{E}_{n}$ corresponds to a path graph.
2.2 Fitting the full model
When there are no restrictions on $\operatorname{\boldsymbol{\nu}}$ the LSE of is given by
$\widehat{\operatorname{\boldsymbol{\nu}}}=\arg\min_{\operatorname{\boldsymbol{\nu}}\in\mathcal{N}}Q(\operatorname{\boldsymbol{\nu}})=\arg\min_{\operatorname{\boldsymbol{\nu}}\in\mathcal{N}}\sum_{1\leq i<j\leq K}\sum_{k=1}^{n_{ij}}(Y_{ijk}-\nu_{ij})^{2}$ (11)
and equals
$\widehat{\operatorname{\boldsymbol{\nu}}}=\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL},$
where $\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}$ is a $|\operatorname{\boldsymbol{\nu}}|$ vector whose $(i,j)^{th}$ element is $\overline{S}_{ij}=S_{ij}/n_{ij}$ if $n_{ij}>0$ and set, arbitrarily, to be 00 otherwise. The following, trivial, result describes the asymptotic behaviour of $\widehat{\operatorname{\boldsymbol{\nu}}}$:
Theorem 2.2
If $n_{ij}\to\infty$ for all $1\leq i<j\leq K$ and Condition 2.1 holds then
$\operatorname{\boldsymbol{D}}^{1/2}(\widehat{\operatorname{\boldsymbol{\nu}}}-\operatorname{\boldsymbol{\nu}})\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\sigma^{2}\boldsymbol{I}),$
where $\boldsymbol{D}$ is a $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ diagonal matrix whose $(i,j)^{th}$ diagonal element is $n_{ij}$. Furthermore, if $n_{ij}=O(n)$ for all $1\leq i<j\leq K$ then
$\sqrt{n}(\widehat{\operatorname{\boldsymbol{\nu}}}-\operatorname{\boldsymbol{\nu}})\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\sigma^{2}\,\boldsymbol{\Xi}^{+}).$
Remark 2.2
Note that $\nu_{ij}$ can not be estimated if $n_{ij}=0$ and can not be consistently estimated whenever $n_{ij}\not\to\infty$.
It is interesting to compare Theorem 2.2 with the limit in (8). First note the former requires that all $n_{ij}\rightarrow\infty$ whereas the latter requires much less, i.e., Condition 2.3. The difference between items $i$ and $j$ is estimated by by $\widehat{\nu}_{ij}$, the $(i,j)^{th}$ element of $\widehat{\operatorname{\boldsymbol{\nu}}}$, when the full model is fit, and by $\widehat{\mu}_{i}-\widehat{\mu}_{j}=(e_{i}-e_{j})^{\top}\widehat{\operatorname{\boldsymbol{\mu}}}$ when the reduced models is fit. If the reduced model holds the latter estimator is typically much more accurate. For example, consider a PCG in which $n_{ij}=m$ for all $1\leq i<j\leq K$. It is easy to verify that under the full model $\mathbb{V}(\widehat{\nu}_{ij})=\sigma^{2}/m$ whereas under the reduced model it is $\mathbb{V}(\widehat{\nu}_{ij})=\mathbb{V}(\widehat{\mu}_{i}-\widehat{\mu}_{j})=2\sigma^{2}/mK+o(1/m)$ the latter is approximately $K/2$ times more precise than the former, exemplifying the benefits of fitting the reduced model when it is in fact correct.
2.3 The relationship between the reduced and full model
Next we examine what happens when model (1) is fit assuming $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ in circumstances in which it possibly shouldn’t.
Theorem 2.3
Assume the errors are IID $\mathcal{N}(0,\sigma^{2})$ RVs. Let $\mathrm{KL}(\mathcal{N},\mathcal{M})$ denote the Kullback–Leibler divergence between the distribution of $\operatorname{\boldsymbol{Y}}$ indexed by $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ and $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ respectively. Then:
${\operatorname{\boldsymbol{\mu}}}=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\mathrm{KL}(\mathcal{N},\mathcal{M})=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\sum_{1\leq i<j\leq K}n_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}.$
Moreover, if Conditions 2.1 and 2.2 hold then the LSE given by (6) will converge to
$\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\{\sum_{1\leq i<j\leq K}\theta_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}:\boldsymbol{v}^{\top}\operatorname{\boldsymbol{\mu}}=0\}.$ (12)
The minimiser of (12) is unique if Condition 2.3 holds. Furthermore the solution of (12) is independent of $\boldsymbol{\Theta}$ if and only if $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$. Finally if the constraint $\boldsymbol{v}=\operatorname{\boldsymbol{1}}1$ is imposed then the linear relation
$\operatorname{\boldsymbol{\mu}}=(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}.$ (13)
holds.
Theorem 2.3 explores the relationship between $\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}$ and $\operatorname{\boldsymbol{\nu}}\in\mathbb{R}^{K(K-1)/2}$. One of its consequences is that the LSE, assuming the reduce model, given by (7), will converge to the LHS of (12). Obviously this limit depends very strongly on the limiting value of the Laplacian, as illustrated by the following example.
Example 2.1
Let $K=4$ and suppose that $\operatorname{\boldsymbol{\nu}}=(-2,0,2,0,2,0)$. It is easy to check that $\operatorname{\boldsymbol{\nu}}\notin\mathcal{M}$. By Theorem 2.3 the LSE of $\operatorname{\boldsymbol{\mu}}$, will converge to the RHS of (13). If the comparison graph is complete and balanced, i.e., $n_{ij}=m$ for $1\leq i<j\leq K$, then $\boldsymbol{\Xi}=diag(1/6,1/6,1/6,1/6,1/6,1/6)$; whereas if it is a balanced path graph, i.e., $n_{12}=n_{23}=n_{34}=m$, then $\boldsymbol{\Xi}=diag(1/3,0,0,1/3,0,1/3)$. In both cases $\operatorname{\boldsymbol{B}}$ is given by (10). Using (13) we find that the the LSE will converge to $(0,1,0,-1)$ for the complete graph and to $(-1.5,0.5,0.5,0.5)$ for the path graph. The resulting merits are obviously very different and topology dependent.
It is worth noting that if we substitute the sample versions of $\boldsymbol{\Xi}$ and $\operatorname{\boldsymbol{\nu}}$ on the RHS of (13) then its LHS, i.e., $\operatorname{\boldsymbol{\mu}}$, coincides with the LSE (7). Therefore (13) provides an alternative way of computing the LSE. Further note that using (13) one can find the value of $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$, denoted by $\operatorname{\boldsymbol{\nu}}_{\operatorname{\boldsymbol{\mu}}}$, providing the best fit for the model when $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$. In fact
$\operatorname{\boldsymbol{\nu}}_{\operatorname{\boldsymbol{\mu}}}=\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}=\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}$
so $\|\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{\nu}}_{\operatorname{\boldsymbol{\mu}}}\|\leq\|\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\|_{F}\|\operatorname{\boldsymbol{\nu}}\|$, where $\|\cdot\|_{F}$ denotes the Frobenius norm. It should also be emphasized that $\operatorname{\boldsymbol{\nu}}_{\operatorname{\boldsymbol{\mu}}}$ is a projection onto $\mathcal{M}$ if and only if $\boldsymbol{\Xi}=a\boldsymbol{I}$ for some $a\in\mathbb{R}$, i.e., if and only if the comparison graph is balanced in which case $n_{ij}=m$ for all pairs $(i,j)$.
We conclude this subsection by applying Theorem 2.3 in the context of Theorem 2.1 and deriving the non–null distribution of $R_{n}$ under local alternatives.
Proposition 2.1
Assume that Conditions 2.1 and 2.3 hold. Suppose $\operatorname{\boldsymbol{\nu}}_{n}=\operatorname{\boldsymbol{\nu}}+n^{-1/2}\boldsymbol{\delta}$ where $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ and $\boldsymbol{\delta}\notin\mathcal{M}$ is a fixed vector. If $n\rightarrow\infty$ then
$R_{n}\Rightarrow\sum_{i=1}^{r}\lambda_{i}(Z_{i}+\gamma_{i})^{2},$
where $\lambda_{1},\ldots,\lambda_{r}$ and $Z_{1},\ldots,Z_{r}$ are as in the statement of Theorem 2.1 and $\gamma_{1},\ldots,\gamma_{r}$ are the elements of the vector $\boldsymbol{\gamma}=\boldsymbol{O}(\boldsymbol{\Psi}^{+})^{1/2}\boldsymbol{\Xi}^{1/2}\boldsymbol{\delta}$ which correspond to the nonzero eigenvalues of $\boldsymbol{\Psi}$. Here $\boldsymbol{O}$ is the orthonormal matrix whose columns are the eigenvectors of $\boldsymbol{\Psi}$.
Remark 2.3
There are many natural alternatives to $R_{n}$. Note that we can express $R_{n}=\boldsymbol{U}_{n}^{\top}\boldsymbol{U}_{n}$ where $\boldsymbol{U}_{n}$ is a random vector whose components are $U_{ij}=\sqrt{n_{ij}}({S}_{ij}/n_{ij}-(\widehat{\mu}_{i}-\widehat{\mu}_{j}))$ if $(i,j)\in\mathcal{E}_{n}$ and $U_{ij}=0$ otherwise. Thus, a Wald type test of the form $W_{n}=\boldsymbol{U}_{n}^{\top}\boldsymbol{\Psi}_{n}^{+}\boldsymbol{U}_{n}$, where $\boldsymbol{\Psi}_{n}=\mathbb{V}ar(\boldsymbol{U}_{n})$ is also possible. It can be further shown that $\boldsymbol{\Psi}_{n}\rightarrow\boldsymbol{\Psi}$ in probability so under the null $W_{n}\Rightarrow\chi^{2}(r)$, whereas under the conditions of Proposition 2.1 $W_{n}\Rightarrow\chi^{2}(r,\boldsymbol{\delta}^{\top}\boldsymbol{\Psi}\,\boldsymbol{\delta})$ where $\chi^{2}(a,b)$ is a non–central chi–square RV with $a$ degrees of freedom and non–centrality parameter $b$. The last two limits require the use of Theorem 2 in Moore (1977).
3 Understanding and decomposing $\mathcal{N}$
The goal of this section is to provide an understanding of the structure of the parameter space $\mathcal{N}$. This is achieved by decomposing $\mathcal{N}$ into two orthogonal components referred to as the linear and cyclical components. Meaningful spanning sets for these components are provided. The recognition that these spanning sets enable the modelling and identification of intransitivity and cyclicality is the key to all further developments.
3.1 Complete PCGs
We start by assuming that the comparison graph $\mathcal{G}$ is complete and consequently the full paired preference profile, i.e., the entire vector $\operatorname{\boldsymbol{\nu}}$ is available. Later, see Section 3.2, this assumption is relaxed.
It is easily verified that if $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ it can be expressed as
$\operatorname{\boldsymbol{\nu}}=(\mu_{1}-\mu_{2},\ldots,\mu_{K-1}-\mu_{K})^{\top}=\mu_{1}\boldsymbol{b}_{1}+\ldots+\mu_{K}\boldsymbol{b}_{K}=\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}},$ (14)
where the vectors $\boldsymbol{b}_{1},\ldots,\boldsymbol{b}_{K}$ are the columns of the matrix $\operatorname{\boldsymbol{B}}$ defined in Section 2.1. Clearly $\mathcal{M}=\mathrm{span}(\boldsymbol{b}_{1},\ldots,\boldsymbol{b}_{K})$. Consequently, $\mathcal{M}$ is a linear subspace of $\mathcal{N}$ for which $\{\boldsymbol{b}_{1},\ldots,\boldsymbol{b}_{K}\}$ is a natural spanning set, see Saari (2021). Since $\sum_{k=1}^{K}\boldsymbol{b}_{k}=0$ it follows that the dimension of $\mathcal{M}$ is $K-1$.
Jiang et al. (2011), using graphical combinatorial Hodge theory (cf., Lim, 2020), and Saari (2014) and Saari (2021), using the language of linear algebra, established that:
Fact 3.1
If $\operatorname{\boldsymbol{\nu}}$ satisfies
$\nu_{ij}+\nu_{jk}+\nu_{ki}=0$ (15)
for all $(i,j,k)\in\{1,2,\ldots,K\}^{3}$ then $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$.
We say that $\operatorname{\boldsymbol{\nu}}$ is consistent on the triad, or triangle, $(i,j,k)$ if (15) holds, and globally consistent if (15) holds on every triad. Global consistency is equivalent to a total linear ordering, i.e., the parametrization (3). If (15) does not hold, then a cyclical preference relation among the items $i,j$ and $k$ obtains. For example, if $(\nu_{ij},\nu_{jk},\nu_{ki})=(1,-1,1)$ then (15) is obviously violated, moreover $i\succ j\succ k\succ i$; the latter preference relation is captured by the cyclical component of $\operatorname{\boldsymbol{\nu}}$ which lies outside of $\mathcal{M}$. Next, denote the orthogonal complement of $\mathcal{M}$ in $\mathcal{N}$ by $\mathcal{N}\setminus\mathcal{M}$. It follows that any $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ can be expressed as a sum of two orthogonal vectors, $\operatorname{\boldsymbol{\nu}}_{\rm linear}\in\mathcal{M}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in\mathcal{N}\setminus\mathcal{M}$ for which
$\operatorname{\boldsymbol{\nu}}=\operatorname{\boldsymbol{\nu}}_{\rm linear}+\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ (16)
where $\operatorname{\boldsymbol{\nu}}_{\rm linear}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ capture, respectively, the linear and cyclic aspects of the preference relations (Saari, 2021).
Example 3.1
Suppose that $K=3$ and $\operatorname{\boldsymbol{\nu}}=(1,-1,1)$. Clearly $\nu_{12}+\nu_{23}+\nu_{31}=3$ so $\operatorname{\boldsymbol{\nu}}\notin\mathcal{M}$. It is easy to verify by (9) that $\boldsymbol{b}_{1}=(1,1,0)$, $\boldsymbol{b}_{2}=(-1,0,1)$ and $\boldsymbol{b}_{3}=(0,-1,-1)$. Clearly $\operatorname{\boldsymbol{\nu}}^{\top}\boldsymbol{b}_{i}=0$ for all $i$ so $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}\setminus\mathcal{M}$ consequently $\operatorname{\boldsymbol{\nu}}_{\rm linear}=\operatorname{\boldsymbol{0}}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{\nu}}$. Since $\nu_{12}=1$ we have $1\succ 2$; similarly $\nu_{23}=1$ and $\nu_{13}=-1$ imply that $2\succ 3$ and $3\succ 1$ respectively. Combining these preference relations we find that $1\succ 2\succ 3\succ 1$, i.e., the preference relation is not transitive but cyclical, it is also equivalent to $2\succ 3\succ 1\succ 2$ and $3\succ 1\succ 2\succ 3$, depending on the item is used to anchor the cycle. A cyclic relation with three items is called a cyclic triad.
The following example illuminates the relationship between cyclicality and transitivity. In the context of cardinal PCD a preference relation is linearly transitive if it satisfies (15), strongly stochastically transitive if $i\succeq j$ and $j\succeq k$ implies that $\nu_{ik}\geq\max\{\nu_{ij},\nu_{jk}\}$ and weakly stochastically transitive if $i\succeq j$ and $j\succeq k$ implies $\nu_{ik}\geq 0$, cf.Oliveira et al. (2018) and the references therein for a through discussion of transitivity relations.
Example 3.2
Let $\operatorname{\boldsymbol{\nu}}_{\rm linear}=(1,2,1)\in\mathcal{M}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=(1,-1,1)\in\mathcal{N}\setminus\mathcal{M}$. Clearly $\operatorname{\boldsymbol{\nu}}_{\rm linear}$ is associated with the preference relation $1\succ 2\succ 3$ whereas $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ with the relation $1\succ 2\succ 3\succ 1$ as demonstrated in Example 3.1. Let
$\operatorname{\boldsymbol{\nu}}=\alpha\operatorname{\boldsymbol{\nu}}_{\rm linear}+\beta\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$
where $\alpha,\beta\geq 0$. We avoid negative values as to not flip the preference order. Now, a bit of algebra shows that (1) is linearly transitive if and only if $\beta=0$, strongly transitive if $\alpha>2\beta$ and weakly transitive if $\alpha>\beta/2$. Moreover, if $\alpha=\beta/2$ then $1\succ 2$, $2\succ 3$ but $1\sim 3$, i.e., $\nu_{13}=0$, whereas if $\alpha<\beta/2$ then the cyclical component dominates the linear component, i.e., $1\succ 2\succ 3\succ 1$. Finally the model becomes purely cyclic if and only if $\alpha=0$.
Example 3.1 shows that when $K=3$ the cyclic relation $1\succ 2\succ 3\succ 1$ can be expressed by the vector $(1,-1,1)$ which is orthogonal to $\mathcal{M}$ and completes the basis for $\mathcal{N}$. Similarly, for any $K$ the cyclic relation $i\succeq j\succeq k\succeq i$ can be expressed by the vector $\boldsymbol{c}_{(i,j,k)}\in\mathcal{N}\setminus\mathcal{M}$ whose $(s,t)^{th}$ element is
$c_{(i,j,k)}(s,t)=\mathbb{I}((s,t)\in\{(i,j),(j,k),(k,i)\})-\mathbb{I}((s,t)\in\{(j,i),(k,j),(i,k)\}).$ (17)
Thus the $(i,j)^{th}$ and $(j,k)^{th}$ elements of $\boldsymbol{c}_{(i,j,k)}$ are $1$, the $(i,k)^{th}$ element is $-1$ and all other elements are 00. Equation (17) demonstrates that cyclical relations have a linear algebraic representation. It is easy to verify that
$\boldsymbol{c}_{(i,j,k)}=\boldsymbol{c}_{(j,k,i)}=\boldsymbol{c}_{(k,i,j)}=-\boldsymbol{c}_{(j,i,k)}=-\boldsymbol{c}_{(i,k,j)}=-\boldsymbol{c}_{(k,j,i)},$
hence there are only two types of cycles among any three items; i.e., $i\succeq j\succeq k\succeq i$ or its reverse $k\succeq j\succeq i\succeq k$. Both cycles can be represented by a single vector, i.e., they lie in the same one dimensional linear space. As a consequence only one element among those in the display above is necessary for modelling cyclicality among $\{i,j,k\}$; for convenience we shall choose $\boldsymbol{c}_{(i,j,k)}$ with $i<j<k$. Further note that a cyclic relation on $(i,j,k)$ can be expressed as superposition of the cyclic relations $(s,i,j),(s,j,k)$ and $(s,i,k)$ for some $s\notin\{i,j,k\}$, i.e.,
$\boldsymbol{c}_{(i,j,k)}=\boldsymbol{c}_{(s,i,j)}+\boldsymbol{c}_{(s,j,k)}-\boldsymbol{c}_{(s,i,k)}.$ (18)
Note that the cyclic relation on the right hand side of (18) honor the preference relation in $(i,j,k)$. Equation (18) plays an important role in later developments.
The following was established by Saari (2014) and Saari (2021),
Fact 3.2
The set $\{\boldsymbol{c}_{(i,j,k)}\}_{1\leq i<j<k\leq K}$, whose cardinality is $\binom{K}{3}$, spans $\mathcal{N}\setminus\mathcal{M}$. Furthermore,
$\dim(\mathcal{N}\setminus\,\mathcal{M})=\binom{K}{2}-(K-1)=\binom{K-1}{2}.$
Consequently, for all $K\geq 4$ the spanning set is larger, often much larger, than the dimension of $\mathcal{N}\setminus\,\mathcal{M}$. Let $\operatorname{\boldsymbol{C}}$ be a $\binom{K}{2}\times\binom{K}{3}$ matrix whose columns are $\{\boldsymbol{c}_{(i,j,k)}\}_{1\leq i<j<k\leq K}$. Note that the rows of $\operatorname{\boldsymbol{C}}$ are indexed by the edges $(i,j)$ and the columns are indexed by the triads $(i,j,k)$. It is now clear that if $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}\setminus\mathcal{M}$ we have
$\operatorname{\boldsymbol{\nu}}=\sum_{1\leq i<j<k\leq K}\gamma_{ijk}\,\boldsymbol{c}_{(i,j,k)}=\operatorname{\boldsymbol{C}}\boldsymbol{\gamma},$ (19)
where $\gamma_{ijk}$ is the coefficient associated with the cyclical triad ${(i,j,k)}$ and $\boldsymbol{\gamma}\in\mathbb{R}^{\binom{K}{3}}$. Consequently for any $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ we have
$\operatorname{\boldsymbol{\nu}}=\operatorname{\boldsymbol{\nu}}_{\rm linear}+\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}+\operatorname{\boldsymbol{C}}\boldsymbol{\gamma}.$ (20)
Consequently the linear and cyclical components of $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ are given by
$\operatorname{\boldsymbol{\nu}}_{\rm linear}=\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{\nu}}~{}\mbox{and}~{}~{}\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{C}}(\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{C}})^{+}\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{\nu}}.$ (21)
It follows that we can write
$\operatorname{\boldsymbol{\nu}}=\|\operatorname{\boldsymbol{\nu}}_{\rm linear}\|\frac{\operatorname{\boldsymbol{\nu}}_{\rm linear}}{\|\operatorname{\boldsymbol{\nu}}_{\rm linear}\|}+\|\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\|\frac{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}{\|\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\|},$
so the quantity $\|\operatorname{\boldsymbol{\nu}}_{\rm linear}\|/\|\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\|$ may be used to quantify the degree to which $\operatorname{\boldsymbol{\nu}}$ is linear or cyclical.
We conclude that given $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ we can identify both its linear and cyclical components. The parameter $\operatorname{\boldsymbol{\mu}}$ is identifiable provided a constraint, such as $\sum_{i=1}^{K}\mu_{i}=0$, is imposed. The identifiability of the parameter $\boldsymbol{\gamma}$ is discussed below. In the following, for the purpose of conciseness, the terms model and parameter are used interchangeably.
Definition 3.1
Consider the set $\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}=\{\boldsymbol{\gamma}\in\mathbb{R}^{\binom{K}{3}}:\,\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{C}}\boldsymbol{\gamma}\}$. For $\boldsymbol{\gamma}\in\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ let $s(\boldsymbol{\gamma})=\sum\mathbb{I}(\gamma_{ijk}\neq 0)$ denote the size of the model associated with $\boldsymbol{\gamma}$. A model $\boldsymbol{\gamma}\in\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ is said to be minimal if $s(\boldsymbol{\gamma})\leq s(\boldsymbol{\gamma}^{\prime})$ for all $\boldsymbol{\gamma}^{\prime}\in\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$.
Minimal models are parsimonious, i.e., they include only those cyclical triads which are necessary for explaining the observed preference relations. We shall always assume that the data generating mechanism is associated with a minimal model; consequently non–minimal models, which incorporate some unnecessary triads, are of no interest to us. It is emphasized that the assumption that the observed data is generated by a minimal model is reasonable and useful, but not testable. We also note that the construction of parsimonious models requires that, at least initially, all possible cyclical triads in the set $\{\boldsymbol{c}_{(i,j,k)}\}_{1\leq i<j<k\leq K}$ be considered. We associate each minimal model with the set $\mathcal{C}_{\boldsymbol{\gamma}}$ of columns of $\operatorname{\boldsymbol{C}}$ corresponding to the non–zero coefficients of $\boldsymbol{\gamma}$.
Definition 3.2
The span and support of a minimal model are
${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}})={\rm span}(\{\boldsymbol{c}_{(i,j,k)}:\,\gamma_{ijk}\neq 0\})~{}~{}\mbox{and}~{}~{}{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})=\{(i,j):\boldsymbol{c}_{(i,j,k)}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}})\}.$
Ideally minimal models would have a unique support and be identifiable from their span. Unfortunately this is not always so.
Example 3.3
For any $K\geq 6$ let $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,2,3)}-\boldsymbol{c}_{(1,2,4)}+\boldsymbol{c}_{(1,3,6)}-\boldsymbol{c}_{(2,3,5)}$. We refer to this model as $\mathcal{C}_{\boldsymbol{\gamma}_{1}}$. Using (18) we also express $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ as $-\boldsymbol{c}_{(1,3,4)}+\boldsymbol{c}_{(1,3,6)}+\boldsymbol{c}_{(2,3,4)}-\boldsymbol{c}_{(2,3,5)}$ which we refer to as $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$. Thus a minimal model may not be unique. Moreover, ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\neq{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ so minimal models may be associated with different linear subspaces. Observe that $(1,2)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$ and $(1,2)\notin{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Thus ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\neq{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$.
Example 3.3 seems to suggest that minimal models, as well as the associated linear spaces and supports, can not be identified. This view, however, is overly pessimistic view. Example 3.3 is possible due to two special features: $(i)$ the vector $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ is a linear combination of triads with high degree of overlap, such as $(1,2,3)$ and $(1,2,4)$ and; $(ii)$ the coefficients of the linear combinations generating $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ are all the same (and equal to unity). Consequently, various elements in $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ are “zeroed out” precluding a unique minimal model. In practice, these conditions are not expected to hold simultaneously.
Theorem 3.1
For any minimal models $\boldsymbol{\gamma}_{1}$ and $\boldsymbol{\gamma}_{2}$ we have $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\cap{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Moreover, if ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ then ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Furthermore, if: $(i)$ $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ implies that $\nu_{{\rm cyclic},ij}\neq 0$; and $(ii)$ ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ implies that ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$; then $\boldsymbol{\gamma}_{1}$ and $\boldsymbol{\gamma}_{2}$ span the same space. Finally, if any two triads in a minimal model share at most one index then the minimal model is unique.
Theorem 3.1 indicates that in practice minimal models are likely to be unique and hence identifiable from their support. Moreover, when a unique minimal model does not exist then the span of all possible minimal models is often the same. For example,
Example 3.4
Let $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\gamma_{1}\boldsymbol{c}_{(1,2,3)}+\gamma_{2}\boldsymbol{c}_{(1,2,4)}+\gamma_{3}\boldsymbol{c}_{(1,3,4)}$. Using (18) we may rewrite it $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\gamma_{1}(\boldsymbol{c}_{(1,2,4)}-\boldsymbol{c}_{(1,3,4)}+\boldsymbol{c}_{(2,3,4)})+\gamma_{2}\boldsymbol{c}_{(1,2,4)}+\gamma_{3}\boldsymbol{c}_{(1,3,4)}=(\gamma_{1}+\gamma_{2})\boldsymbol{c}_{(1,2,4)}+(\gamma_{3}-\gamma_{1})\boldsymbol{c}_{(1,3,4)}+\gamma_{1}\boldsymbol{c}_{(2,3,4)}$. There are many other possible minimal models all of which have the same span.
Note that $\nu_{{\rm cyclic},ij}$, the $(i,j)^{th}$ component of $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$, is given by
$\nu_{{\rm cyclic},ij}=\sum_{1\leq s<t<u\leq K}c_{(s,t,u),ij}\gamma_{stu}.$ (22)
The condition that $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ implies that $\nu_{{\rm cyclic},ij}\neq 0$ and means that the sum in (22) is non–zero for $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$. This will happen if the absolute values of the non–zero elements of $\boldsymbol{\gamma}$ are all distinct as are all of their partial sums. Such a condition is very likely to hold in practice.
Next, we explore the relationship between $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ and minimal models. It is easily verified that for any triad $(i,j,k)$
$\nu_{ij}+\nu_{jk}+\nu_{ki}=\nu_{{\rm cyclic},ij}+\nu_{{\rm cyclic},jk}+\nu_{{\rm cyclic},ki}.$
Thus cyclic triads, i.e., triads violating (15), can also be identified using $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$. It is emphasized that some triads violating (15) may not be part of any minimal model. For example if $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,2,3)}$, then any triad $(i,j,k)$ sharing two indices with triad the $(1,2,3)$, e.g., $(1,2,4)$, will be a cyclical triad, i.e., will not satisfy (15), but will not be part of any minimal model.
Let $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ denote a table with rows and columns indexed by the edges and triads as the matrix $\operatorname{\boldsymbol{C}}$. If the $(s,t)^{th}$ element of $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ is non–zero we place a tick mark in every column in which the combination $(s,t)$ occurs, i.e., in any column indexed by $(s,t,u)$, for some $u\notin\{s,t\}$, or possible permutations thereof. Consequently the columns of the table $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$, referred to as the tick–table, may have $0,1,2$ or $3$ ticks, cf. Example 3.5 below.
Theorem 3.2
If there is a minimal model $\boldsymbol{\gamma}$ with some 00-tick triads then there exists another minimal model $\boldsymbol{\gamma}^{\prime}$ for which no triad has 00-ticks. Additionally, if any two triads in a minimal model share at most one index then only triads in the minimal model will have $3$–ticks. However not all $3$–tick triads are members of a minimal model. Finally, if the minimal model is unique and some of its triads share two indices then some of the triads may have less than $3$–ticks.
Theorem 3.2 shows that when the minimal model is unique it can be identified using the tick–table $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$, whereas if the minimal model is not unique then some minimal model can always be identified. The following example presents both pessimistic and realistic scenarios.
Example 3.5
Set $K=6$ and consider $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ as in Example 3.3. A bit of work show that $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ is given by:
123 124 125 126 134 135 136 145 146 156 234 235 236 245 246 256 345 346 356 456
12
13
14
15
16
23
24
25
26
34
35
36
45
46
56
Note that the $123$ column of $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ shows 00 ticks while $\boldsymbol{c}_{(1,2,3)}\in\boldsymbol{\gamma}_{1}$. Another minimal model is $-\boldsymbol{c}_{(1,3,4)}+\boldsymbol{c}_{(1,3,6)}+\boldsymbol{c}_{(2,3,4)}-\boldsymbol{c}_{(2,3,5)}$, in which all triads have $1$ or $2$ ticks. Thus if the minimal model is not unique then there will be many columns with one or two ticks, and it will be difficult to identify a minimal model by using $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$. However, if
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\gamma_{123}\boldsymbol{c}_{(1,2,3)}+\gamma_{124}\boldsymbol{c}_{(1,2,4)}+\gamma_{136}\boldsymbol{c}_{(1,3,6)}+\gamma_{235}\boldsymbol{c}_{(2,3,5)}$
and $\gamma_{123},\ldots,\gamma_{235}$ are all different then the minimal model is unique. Moreover, the columns corresponding to the triads $123^{th}$, $124^{th}$, $235^{th}$ and $136^{th}$ will be associated with three ticks. All other columns will have less than $3$ ticks. Thus it is straightforward to identify the minimal model in this case.
3.2 Incomplete PCGs
A PCG is referred to as incomplete if some edges are not present in $\mathcal{G}$. Clearly, if the edge $(s,t)$ is not in $\mathcal{G}$ then the items $s$ and $t$ are not directly compared and consequently $\nu_{st}$ is not accessible. Furthermore all triads indexed by $s,t$ and any other item are not present either, consequently cyclicalities associated with the pair $(s,t)$ can not be assessed. As an extreme example consider a PCG in which $\mathcal{G}$ is a path graph. Thus, only $K-1$ edges are present, ensuring connectivity, but no triad is observed. Moreover, under some labeling of the items, we are only privy to the comparisons $\nu_{12},\ldots,\nu_{K-1K}$ and the complete preference profile $\operatorname{\boldsymbol{\nu}}$ is not available. Consequently, (21) can not be applied and $\operatorname{\boldsymbol{\nu}}_{\rm linear}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclical}$ can not, in general, be identified. However, assuming $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ it is possible to reconstruct $\operatorname{\boldsymbol{\nu}}$ as $\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}$ where $\operatorname{\boldsymbol{\mu}}$ is given in (13). Thus, it is clear that incomplete PCGs may hamper the ability to detect all cyclicalities and thus reconstruct the true preference profile $\operatorname{\boldsymbol{\nu}}$.
Let $\mathcal{G}$ denote an incomplete PCG. Clearly $\mathcal{G}$ is a subgraph of the full comparison graph. Let $\boldsymbol{E}$ be the $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ diagonal matrix whose $(i,j)^{th}$ diagonal element is $1$ if comparison $(i,j)\in\mathcal{E}$ and 00 otherwise. Define $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}}=\boldsymbol{E}\operatorname{\boldsymbol{\nu}}$, i.e., all unavailable elements in $\operatorname{\boldsymbol{\nu}}$ are “zeroed out” in $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}}$. The substitution of a 00 for a missing value is arbitrary. Further let ${\rm Tri}(\mathcal{G})$ denote the triads associated with the graph $\mathcal{G}$. Similarly let $\operatorname{\boldsymbol{C}}_{\mathcal{G}}$ denote the matrix whose columns are the elements of the set ${\rm Tri}(\mathcal{G})$ and let $\boldsymbol{\gamma}_{\mathcal{G}}$ denote the corresponding coefficients. As in Equation (21) the following orthogonal decomposition of $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}}$ holds:
$\operatorname{\boldsymbol{\nu}}_{\mathcal{G},\,\rm linear}=\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{E}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{\nu}}_{\mathcal{G}},~{}~{}\operatorname{\boldsymbol{\nu}}_{\mathcal{G},\,\rm cyclic}=\operatorname{\boldsymbol{C}}_{\mathcal{G}}(\operatorname{\boldsymbol{C}}_{\mathcal{G}}^{\top}\operatorname{\boldsymbol{C}}_{\mathcal{G}})^{+}\operatorname{\boldsymbol{C}}_{\mathcal{G}}^{\top}\operatorname{\boldsymbol{\nu}}_{\mathcal{G}}.$ (23)
Observe that for the complete graph $\boldsymbol{E}=\boldsymbol{I}$ and the decomposition above is exactly as in (21).
If $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\notin{\rm span}({\rm Tri}(\mathcal{G}))$ then we can not access $\operatorname{\boldsymbol{\nu}}$ or its components $\operatorname{\boldsymbol{\nu}}_{\rm linear}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$. For example, suppose $K\geq 4$ and set $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,2,3)}$. It is clear that any partial PCG in which one of the edges $(1,2),(1,3)$ or $(2,3)$ is not present, can not identify all components of $\operatorname{\boldsymbol{\nu}}$. On the other hand any connected graph which includes the latter edges completely identifies $\operatorname{\boldsymbol{\nu}}$, and thus its components. Thus:
Theorem 3.3
If $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}({\rm Tri}(\mathcal{G}))$ then there exist coefficient $\boldsymbol{\gamma}_{\mathcal{G}}\in\mathbb{R}^{{\rm rank}({\rm span}({\rm Tri}(\mathcal{G})))}$ such that
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{C}}_{\mathcal{G}}\boldsymbol{\gamma}_{\mathcal{G}}.$
Furthermore, if the conditions of Theorem 3.1 hold for $\boldsymbol{\gamma}_{\mathcal{G}}$ then there exists a unique minimal model for $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}({\rm Tri}(\mathcal{G}))$.
Theorem 3.3 says that when the minimal model for $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ is in the span of ${\rm Tri}(\mathcal{G})$ then the components of $\operatorname{\boldsymbol{\nu}}$ can be identified by $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}}$ and given by (23). If, however, $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}}\neq\operatorname{\boldsymbol{\nu}}_{\mathcal{G},\,\rm linear}+\operatorname{\boldsymbol{\nu}}_{\mathcal{G},\,\rm cyclic}$ then $\operatorname{\boldsymbol{\nu}}_{\mathcal{G},\,\rm linear}\neq\operatorname{\boldsymbol{\nu}}_{\rm linear}$, $\operatorname{\boldsymbol{\nu}}_{\mathcal{G},\,\rm cyclic}\neq\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ and, moreover, $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}}$ does not identify all cyclicalities.
In such situations the unavailable complete preference profile can be expressed as:
$\operatorname{\boldsymbol{\nu}}=\operatorname{\boldsymbol{\nu}}_{\mathcal{G},\,\rm linear}+\operatorname{\boldsymbol{\nu}}_{\mathcal{G},\,\rm cyclic}+\operatorname{\boldsymbol{\nu}}^{*}$ (24)
where $\operatorname{\boldsymbol{\nu}}^{*}$, which is not available, can be further partitioned as $\operatorname{\boldsymbol{\nu}}^{*}=\operatorname{\boldsymbol{\nu}}^{*}_{\rm linear}+\operatorname{\boldsymbol{\nu}}^{*}_{\rm cyclic}+\operatorname{\boldsymbol{\nu}}^{*}_{\rm harmonic}$, where $\operatorname{\boldsymbol{\nu}}^{*}_{\rm linear}\in\mathcal{M}$, $\operatorname{\boldsymbol{\nu}}^{*}_{\rm cyclic}\in{\rm span}({\rm Tri}(\mathcal{G}))$ and $\operatorname{\boldsymbol{\nu}}^{*}_{\rm harmonic}\in\mathcal{N}\,\setminus(\mathcal{M}\cup\,{\rm span}({\rm Tri}(\mathcal{G})))$ represents all unobserved cyclicalities. The subscript ${\rm harmonic}$ appears in Jiang et al. (2011) and used, imprecisely, to model cycles of length $>3$. Such cycles will be briefly discussed in Section 7. Here $\operatorname{\boldsymbol{\nu}}^{*}_{\rm harmonic}$ is the residual of $\operatorname{\boldsymbol{\nu}}^{*}$ inaccessible to $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}}$.
Example 3.6
Let $K=4$ and suppose that $\operatorname{\boldsymbol{\nu}}=(4,1,1,2,4,1)$. Clearly $\operatorname{\boldsymbol{\nu}}_{\rm linear}=(1,2,3,1,2,1)$ and $\operatorname{\boldsymbol{\nu}}_{\rm cylcic}=(3,-1,-2,1,2,0)=\boldsymbol{c}_{(1,2,3)}+2\boldsymbol{c}_{(1,2,4)}$. Consider three situations: (i) a complete PCG which we denote by $\mathcal{G}_{1}$; (ii) an incomplete PCG with triads $\{(1,2,3),(1,2,4)\}$ which we denote by $\mathcal{G}_{2}$; and (iii) an incomplete PCG with a single triad $\{(1,2,3)\}$ and an additional edge $(2,4)$, which we denote by $\mathcal{G}_{3}$. By Theorem 3.1 $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{1},\,\rm linear}=\operatorname{\boldsymbol{\nu}}_{\rm linear}$ and $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{1},\,\rm cylcic}=\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$. If the incomplete PCG $\mathcal{G}_{2}$ is given then $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{2}}=(4,1,1,2,4,0)$ where the last component is set to 00 since the edge $(3,4)\notin\mathcal{G}_{2}$. Using (23) we find that $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{2},\,\rm linear}=(1,2,3,1,2,1)$ and $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{2},\,\rm cylcic}=(3,-1,-2,1,2,0)$. Observe that $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{2},\,\rm linear}=\operatorname{\boldsymbol{\nu}}_{\rm linear}$ and $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{2},\,\rm cylcic}=\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ as expected by Theorem 3.3. Finally, when $\mathcal{G}_{3}$ is given then $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{3}}=(4,1,0,2,4,0)$ and by (23) $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}3,\,\rm linear}=(7,8,19,1,12,11)/3$ and $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{3},\,\rm cylcic}=5\boldsymbol{c}_{(1,2,3)}/3$. However, $\mathcal{G}_{3}$ does not satisfy the conditions of Theorem 3.3, so the linear and cyclic components can not be correctly identified; i.e., $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{3},\,\rm linear}\neq\operatorname{\boldsymbol{\nu}}_{\rm linear}$ and $\operatorname{\boldsymbol{\nu}}_{\mathcal{G}_{3},\,\rm cylcic}\neq\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$. Furthermore by (24)
$\operatorname{\boldsymbol{\nu}}^{*}=\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{\nu}}_{\mathcal{G}3,\,\rm linear}-\operatorname{\boldsymbol{\nu}}_{\mathcal{G}3,\,\rm cyclic}=(0,0,-16/3,0,0,-8/3),$
$\operatorname{\boldsymbol{\nu}}^{*}_{\rm linear}=(-40,-8,-112,32,-72,-104)/9$, $\operatorname{\boldsymbol{\nu}}^{*}_{\rm cyclic}=\operatorname{\boldsymbol{0}}$ and $\operatorname{\boldsymbol{\nu}}^{*}_{\rm harmonic}=(40,8,64,-32,72,80)/9$.
3.3 Ranking
If there are no cyclicalities then $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ then ranking is based solely on the vector of merits ${\operatorname{\boldsymbol{\mu}}}$ given in (13), i.e., $i\succeq j$ if and only if $\mu_{i}\geq\mu_{j}$. As expected ranking is more complicated when $\operatorname{\boldsymbol{\nu}}\notin\mathcal{M}$.
Theorem 3.4
Let $\operatorname{\boldsymbol{\nu}}$ be transitive. Then $i\succeq j$ implies that $\mu_{i}\geq\mu_{j}$ if $\operatorname{\boldsymbol{\nu}}$ is strongly transitive but not if it is only weakly transitive.
We conclude that if $\operatorname{\boldsymbol{\nu}}$ is strongly transitive then the merits can be used for ranking; but not so under weak stochastic transitivity. Nevertheless items can always be ranked even under weak transitivity.
Next consider intransitive models. Saari (2014) and Saari (2021) addressed this issue by treating cyclicalities as noise. For $i=1,\ldots,K$ he proposed to compute the pseudo merit of the $i^{th}$ merit by
$\mu_{i}^{*}=\sum_{j=1}^{K}\nu_{ij}$ (25)
and rank $i$ above $j$ if the pseudo merits satisfy $\mu_{i}^{*}\geq\mu_{j}^{*}$. This approach is a relative of well known row–sum method (Huber, 1963) and is appropriate assuming that the comparison graph is complete and balanced. Further note that for a complete graph the ranking based on $\mu_{i}$ and $\mu_{i}^{*}$ are identical. For more general comparison graphs one may recompute the pseudo merits as $\sum_{j=1}^{K}\theta_{ij}\nu_{ij}$. In the same spirit one can compute
$\mu_{i}^{**}=\sum_{j=1}^{K}\mathbb{I}(\nu_{ij}>0)$ (26)
which we refer to as the dominance score. Notice that the dominance score is reminiscent of the well known Borda–Count advocated among some voting theorists (Saari 2023). Using the dominance score rank $i$ above $j$ if $\mu_{i}^{**}>\mu_{j}^{**}$. Ties among the dominance score can be further broken by repeatedly computing (26) on sets of the form $\{k:\mu_{i}^{**}=\mu_{j}^{**}\}$. The following proposition suggests that dominance scores can be used for ranking.
Proposition 3.1
If $\mu_{i}^{**}\neq\mu_{j}^{**}$ for all $i\neq j$ then $\operatorname{\boldsymbol{\nu}}$ is weakly transitive.
Equations (25) and (26) map $\mathcal{N}\mapsto P_{n}$, the space of all permutations of $\{1,\ldots,n\}$. Such mappings are of course not completely satisfactory. In fact, as noted in Section 1, cyclicalities are often an important aspect of the data and should not be averaged out and discarded.
Example 3.7
Consider a PCG with $K=4$, $\operatorname{\boldsymbol{\nu}}_{\rm linear}=(a,a+b,a+b+c,b,b+c,c)$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\lambda\,\boldsymbol{c}_{(1,2,3)}$, where $a,b,c,\lambda>0$. Then $\operatorname{\boldsymbol{\nu}}=(a+\lambda,a+b-\lambda,a+b+c,b+\lambda,b+c,c)$. The PCG is weakly transitive if $a+b>\lambda$ and the ranking is $1\succ 2\succ 3\succ 4$. If $a+b\leq\lambda$ a full ranking of the items is not possible as for $i\leq 3$ we have $i\succ 4$ while simultaneously $1\succ 2\succ 3\succ 1$. We denote this preference relation by $\langle 1,2,3\rangle\succ 4$.
Thus, global non–transitivity may allow in some settings for partially transitivity, i.e., the set $\{1,\ldots,K\}$ may be partitioned into $g$ groups satisfying
$\langle i_{1,1},\ldots,i_{1,K_{1}}\rangle\succeq\langle i_{2,1},\ldots,i_{2,K_{2}}\rangle\succeq\ldots\succeq\langle i_{g,1},\ldots,i_{g,K_{g}}\rangle,$ (27)
where $\langle i_{1,1},\ldots,i_{1,K_{1}}\rangle\succeq\langle i_{2,1},\ldots,i_{2,K_{2}}\rangle$ means $i_{1,l_{1}}\succeq i_{2,l_{2}}$ for any $1\leq l_{1}\leq K_{1}$ and $1\leq l_{2}\leq K_{2}$. If strong stochastic transitivity holds among any three items inhabiting three different groups then by Theorem 3.4 the merits of these items can be used to rank them. Finally if (27) holds with $G=1$ then $\operatorname{\boldsymbol{\nu}}$ is purely cyclical.
The preference relations in (27) are rather simple to understand and depict. However much more complicated preference relations are possible, see Section 6 for a concrete real example.
4 Estimation and model selection
In this section a statistical methodology accommodating intransitivty and cyclicality is developed. It is assume that for all pairs $1\leq i\neq j\leq K$ we have $n_{ij}>0$, i.e., the PCG is complete.
4.1 Estimation
Let $\operatorname{\boldsymbol{Y}}$ be the $n\times 1$ vector of outcomes arranged lexicographically, i.e.,
$\operatorname{\boldsymbol{Y}}=(Y_{121},\ldots,Y_{12n_{12}},Y_{131},\ldots,Y_{13n_{13}},\ldots,Y_{K-1,K1},\ldots,Y_{K-1,Kn_{K-1,K}})^{\top}.$
Let $\operatorname{\boldsymbol{L}}$ be the $n\times|\operatorname{\boldsymbol{\nu}}|$ matrix whose elements $L_{ij}\in\{0,1\}$ are such that if the $s^{th}$ element of $\operatorname{\boldsymbol{Y}}$ is $Y_{ijk}$ and $t^{th}$ element of $\operatorname{\boldsymbol{\nu}}$ is $\nu_{ij}$ then $L_{st}=1$ and $L_{su}=0$ for all $u\neq t$. It is easy to verify that $\mathbb{E}(\operatorname{\boldsymbol{Y}})=\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}$. It follows that we can rewrite (1) using (20) as
$\operatorname{\boldsymbol{Y}}=(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}},\,\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}})\begin{pmatrix}\operatorname{\boldsymbol{\mu}}\\ \boldsymbol{\gamma}\end{pmatrix}+\boldsymbol{\epsilon}.$ (28)
Let $Q(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma})=(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{\nu}}(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}))^{\top}(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{\nu}}(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}))$ denote the sum of squares associated with (28). As noted the model (28) is not in general identifiable with respect to $(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma})$ precluding a unique LSE. We start by circumventing this problem by defining
$(\widehat{\operatorname{\boldsymbol{\mu}}},\widehat{\boldsymbol{\gamma}})=\arg\min_{\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}}\{Q(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}):\min\|(\operatorname{\boldsymbol{\mu}}^{\top},\boldsymbol{\gamma}^{\top})^{\top}\|\},$ (29)
i.e., $(\widehat{\operatorname{\boldsymbol{\mu}}},\widehat{\boldsymbol{\gamma}})$ are the minimum–norm solution to (29).
Theorem 4.1
Suppose that $n_{ij}>0$ for all $1\leq i<j\leq K$. Then the LSE in (29) is unique and given by
$\widehat{\operatorname{\boldsymbol{\mu}}}=\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top}(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}\widehat{\boldsymbol{\gamma}})$ (30)
and
$\widehat{\boldsymbol{\gamma}}=(\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{L}}^{\top}(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}})^{+}\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{L}}^{\top}(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{Y}}.$ (31)
The LSE for $\operatorname{\boldsymbol{\mu}}$ given above automatically satisfies the constraint $\boldsymbol{1}^{\top}\operatorname{\boldsymbol{\mu}}=0$ as it does for the reduced model. The estimators given in Theorem 4.1 are computed assuming all columns of $\boldsymbol{C}$ are incorporated into the design matrix. Consequently the value of (31) does not identify the minimal model.
Example 4.1
Consider a PCG with $K=5$, $\operatorname{\boldsymbol{\nu}}=\operatorname{\boldsymbol{\nu}}_{\rm linear}+\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ where
$\operatorname{\boldsymbol{\nu}}_{\rm linear}=(1,2,3,4,1,2,3,1,2,1)^{\top}\mbox{ and }\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=(1,-1,0,0,1,0,0,0,0,0)^{\top},$
i.e., the merits are $(-2,-1,0,1,2)^{\top}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ is comprised of a single vector $\boldsymbol{c}_{(1,2,3)}$ with coefficient $\gamma_{123}=1$. Here $|\mathcal{M}|=4$ and $|\mathcal{N}\setminus\mathcal{M}|=6$. If we choose $n_{ij}=m$ for all $1\leq i,j\leq K$ and let $m\to\infty$. Then, it can be shown, that $\widehat{\boldsymbol{\gamma}}$ will converge to $\to(0.6,0.2,0.2,-0.2,-0.2,0,0.2,0.2,0,0)$, which is different from the minimal model $\boldsymbol{\gamma}$ in which the first component is one and all the rest are zero.
The fact that the minimum norm estimator (29) can not recover minimal model is expected. In the example above $\boldsymbol{\gamma}$ has norm $1$. Using (18) we may reexpress $\boldsymbol{c}_{(1,2,3)}$, the only true cyclical element, as $\alpha_{1}\boldsymbol{c}_{(1,2,3)}+\alpha_{2}\boldsymbol{c}_{(k,1,2)}+\alpha_{3}\boldsymbol{c}_{(k,2,3)}-\alpha_{4}\boldsymbol{c}_{(k,1,3)}$ with $\alpha_{2}=\alpha_{3}=\alpha_{4}$ and $\alpha_{1}+\alpha_{2}=1$. If we set $\alpha_{1}=0.7$ and $\alpha_{2}=0.3$ then the associated norm is $\sqrt{0.76}<1$. Thus the minimum norm solution will tend to assign a non–zero value to too many elements of $\boldsymbol{\gamma}$
Next, consider the situation in which the minimal model is known in advance and is comprised of the triads $(i_{1},j_{1},k_{1}),\ldots,(i_{r},j_{r},k_{r})$ with $r\leq\binom{K-1}{2}$. Let $\operatorname{\boldsymbol{C}}_{s}$ denote the matrix whose columns are $\boldsymbol{c}_{(i_{1},j_{1},k_{1})},\ldots,\boldsymbol{c}_{(i_{r},j_{r},k_{r})}$ and let $\boldsymbol{\gamma}_{s}$ be the corresponding vector of coefficients. Models in which $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$ where $\mathcal{M}\subsetneq\mathcal{S}\subsetneq\mathcal{N}$ will be referred to as intermediate models. In this case we fit the model (28) replacing $\operatorname{\boldsymbol{C}}$ with $\operatorname{\boldsymbol{C}}_{s}$ and $\boldsymbol{\gamma}$ with $\boldsymbol{\gamma}_{s}$. The LSE is given by
$(\widehat{\operatorname{\boldsymbol{\mu}}},\widehat{\boldsymbol{\gamma}}_{s})=\arg\min_{\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}_{s}}\{Q(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}_{s}):\operatorname{\boldsymbol{1}}1^{\top}\operatorname{\boldsymbol{\mu}}=0\}.$ (32)
Theorem 4.2
We have:
  1. 1.
    If $n_{ij}>0$ for all $1\leq i<j\leq K$ then the LSEs for $\operatorname{\boldsymbol{\mu}}$ and $\boldsymbol{\gamma}_{s}$ are given by substituting $\operatorname{\boldsymbol{C}}_{s}$ for $\operatorname{\boldsymbol{C}}$ in (30) and (31). These estimators are unbiased.
  2. 2.
    If Conditions 2.1 holds and $n_{ij}\to\infty$ for $1\leq i<j\leq K$, then $(\widehat{\operatorname{\boldsymbol{\mu}}},\widehat{\boldsymbol{\gamma}}_{s})\to(\operatorname{\boldsymbol{\mu}},{\boldsymbol{\gamma}}_{s})$ in probability as $n\to\infty$.
  3. 3.
    If Conditions 2.1 holds, $n_{ij}\to\infty$ and $n_{ij}=O(n)$ for all $1\leq i<j\leq K$ then
    $\sqrt{n}\begin{pmatrix}\widehat{\operatorname{\boldsymbol{\mu}}}-\operatorname{\boldsymbol{\mu}}\\ \widehat{\boldsymbol{\gamma}}_{s}-\boldsymbol{\gamma}_{s}\end{pmatrix}\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\boldsymbol{\Sigma}_{\mu,\gamma_{s}}),$
    where
    $\boldsymbol{\Sigma}_{\mu,\gamma_{s}}=\sigma^{2}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\end{pmatrix}\boldsymbol{\Xi}^{3}(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}.$
Theorem 4.2 shows that when the minimal model is known the LSE $(\widehat{\operatorname{\boldsymbol{\mu}}},\widehat{\boldsymbol{\gamma}}_{s})$ is unbiased, consistent and asymptotically normal. Next, a test of lack of fit is investigated. Let $\mathcal{S}={\rm span}({\operatorname{\boldsymbol{B}}},{\operatorname{\boldsymbol{C}}}_{s})$ and consider testing $H_{0}:~{}\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$ against $H_{1}:~{}\operatorname{\boldsymbol{\nu}}\notin\mathcal{S}$. Under $H_{0}$, $\operatorname{\boldsymbol{\nu}}$ is estimated by $\widehat{\operatorname{\boldsymbol{\nu}}}_{s}=\operatorname{\boldsymbol{B}}\widehat{\operatorname{\boldsymbol{\mu}}}+\operatorname{\boldsymbol{C}}_{s}\widehat{\boldsymbol{\gamma}}_{s}$. Define
$R_{n,\mathcal{S}}=\boldsymbol{U}_{n,\mathcal{S}}^{\top}\boldsymbol{U}_{n,\mathcal{S}},$
where $\boldsymbol{U}_{n,\mathcal{S}}=\operatorname{\boldsymbol{D}}^{1/2}(\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}-\widehat{\operatorname{\boldsymbol{\nu}}}_{s})$.
Theorem 4.3
Assume Condition 2.1 holds and $n_{ij}=O(n)$ for $1\leq i<j\leq K$. Suppose further that $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$. If so as $n\to\infty$
$R_{n,\mathcal{S}}\Rightarrow\sum_{i=1}^{t}{\lambda}_{i}Z_{i}^{2},$
where $t={\rm rank}({\boldsymbol{\Psi}}_{\mathcal{S}})=\binom{K}{2}-(K+r-1)$, $r$ is the number of columns in $\operatorname{\boldsymbol{C}}_{s}$, $Z_{1},\ldots,Z_{t}$ are independent $\mathcal{N}(0,1)$ RVs and ${\lambda}_{1},\ldots,{\lambda}_{t}$ are the non–zero eigenvalues of the $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ matrix ${\boldsymbol{\Psi}}_{\mathcal{S}}=\sigma^{2}\,\boldsymbol{\Xi}^{1/2}\boldsymbol{M}\,\boldsymbol{\Xi}^{+}\boldsymbol{M}^{\top}\boldsymbol{\Xi}^{1/2}$, where
$\boldsymbol{M}=\boldsymbol{I}-(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\end{pmatrix}\boldsymbol{\Xi}.$
Clearly $R_{n}=R_{n,\mathcal{M}}$ so Theorem 4.3 extends Theorem 2.1 to intermediate models. Furthermore, as noted in Remark 2.3 Wald-type tests of the form $W_{n,\mathcal{S}}=\boldsymbol{U}_{n,\mathcal{S}}^{\top}\boldsymbol{\Psi}_{n,\mathcal{S}}^{+}\boldsymbol{U}_{n,\mathcal{S}}\Rightarrow\chi^{2}(\binom{K}{2}-(K+r-1))$, can also be employed. It can be further shown that under local alternatives, cf. Proposition 2.1, we have
$R_{n,\mathcal{S}}\Rightarrow\sum_{i=1}^{t}{\lambda}_{i}(Z_{i}+\gamma_{i})^{2},$
where $\gamma_{1},\ldots,\gamma_{t}$ are the elements of the vector $\boldsymbol{\gamma}_{s}=\boldsymbol{O}_{s}(\boldsymbol{\Psi}_{s}^{+})^{1/2}\boldsymbol{\Xi}^{1/2}\boldsymbol{\delta}$ which correspond to the nonzero eigenvalues of $\boldsymbol{\Psi}$. Here $\boldsymbol{O}_{s}$ is the orthonormal matrix whose columns are the eigenvectors of $\boldsymbol{\Psi}_{s}$.
4.2 Model selection
Methods for model selection have been extensively studied in the literature. Specifically, in the context of linear models, Hastie et al. (2009) discuss a host of approaches including best subset selection, step–wise regression (backwards and forwards) and approaches based on regularization such as the LASSO (Tibshiriani 1996, Freijeiro‐-González et al. 2022). Most methods of model selection incorporate a screening or evaluation criteria such as AIC, BIC or the adjusted-$R^{2}$, c.f., Chen et al. (2013) for a focused review in the context of linear models and Claeskens and Hjort (2008) for a more general approach and a variety of evaluation criteria.
In this section two new methods for model selection, tailored to the problem at hand, are proposed. The first is motivated by Fact 3.1. First observe that cyclical triads can be identified by testing
$H_{0}:\nu_{ij}+\nu_{jk}+\nu_{ki}=0\text{ against }H_{1}:\nu_{ij}+\nu_{jk}+\nu_{ki}\neq 0.$ (33)
A natural test statistic for (33) is
$T_{n}=T_{n}(i,j,k)=\frac{1}{\sqrt{1/n_{ij}+1/n_{jk}+1/n_{ki}}}(\widehat{\nu}_{ij}+\widehat{\nu}_{jk}+\widehat{\nu}_{ki}),$ (34)
for which:
Proposition 4.1
Let $n_{ij},n_{jk},n_{ik}\to\infty$ at the same rate. Then, under $H_{0}$, $T_{n}\Rightarrow\mathcal{N}(0,\sigma^{2})$ where $\sigma^{2}$ can be consistently estimated by $\sum_{1\leq i<j\leq K}\sum_{k=1}^{n_{ij}}(Y_{ijk}-\widehat{\nu}_{ij})^{2}/n$. Under local alternatives of the form $\nu_{ij}+\nu_{jk}+\nu_{ki}=\delta\sqrt{1/n_{ij}+1/n_{jk}+1/n_{ki}}$ we have $T_{n}\Rightarrow\mathcal{N}(\delta,\sigma^{2})$.
Let $P_{(i,j,k)}$ denote the $p$-value associated with the statistic $T_{n}(i,j,k)$. Apply the test (34) to all triads and arrange these $p$-value according to their significance. Suppose the order is $(i_{1},j_{1},k_{1}),\ldots,(i_{r},j_{r},k_{r})$. Let $\mathcal{S}_{1}=\mathcal{M}$, and for $a=2,\ldots,r$ let $\mathcal{S}_{a+1}=\{(i_{1},j_{1},k_{1}),\ldots,(i_{a},j_{a},k_{a})\}$. We now introduce a model selection method referred to as forward step–wise triad selection (FSTS).
Forward Stepwise Triad Selection ($\rm FSTS$)
1:Test whether $H_{0}:\mathcal{S}=\mathcal{S}_{1}$ holds using Theorem 2.1. If the null is not rejected set $\widehat{\mathcal{S}}=\mathcal{S}_{1}$ and stop. Otherwise, set $a=2$.
2:Test whether $H_{0}:\mathcal{S}=\mathcal{S}_{a}$ holds using Theorem 4.3. If the null is not rejected set $\widehat{\mathcal{S}}=\mathcal{S}_{a}$ and stop. Otherwise,
3:Set $a=a+1$ and repeat Step 2.
FSTS requires $\binom{K}{3}$ simple tests. In addition at each stage of the model fitting process an additional test for lack of fit as prescribed in Theorem 4.3 is performed. Clearly FSTS is relative of forward step wise regression (FSR); the difference being that the order by which covariates are added to the model is determined by the $p$–values $P_{(i,j,k)}$ rather than by a sequence of F–test as in FSR.
Numerical experiments indicate that a small p-value $P_{(i,j,k)}$ does not necessarily imply that the triad $(i,j,k)$ is associated with a minimal model even in large samples. This observation motivates the introduction of a second method which builds on the insights of Section 3, specifically Theorems 3.1 and 3.2. The proposed method is, like FSTS, a relative of FSR, and referred to as forward tick–based selection (FTBS). In contrast with standard methods it does not directly utilize features of the design matrix but instead relies on an estimated tick table. To fix ideas, let $\mathcal{S}\subseteq\mathcal{N}$ denote the smallest subspace which contains all possible minimal models and let $\widehat{\mathcal{S}}$ denote the selected model. To find $\widehat{\mathcal{S}}$ first compute $\widehat{\operatorname{\boldsymbol{\mu}}}$ and $\widehat{\operatorname{\boldsymbol{\nu}}}$ as described in (7) and (11) respectively and $\widehat{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}$ as described in (23). Let $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}}$ be the tick–table corresponding to $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}$ where
$\nu_{{\rm cyclic},ij}^{*}=\begin{cases}\widehat{\nu}_{{\rm cyclic},ij}&\text{ if }\phi_{ij}(\widehat{\operatorname{\boldsymbol{\nu}}}_{{\rm cyclic}})=1\\ 0&\text{ if }\phi_{ij}(\widehat{\operatorname{\boldsymbol{\nu}}}_{{\rm cyclic}})=0.\\ \end{cases}$ (35)
Here $\phi_{ij}(\widehat{\operatorname{\boldsymbol{\nu}}}_{{\rm cyclic}})=1$ if $H_{0}^{(i,j)}:\nu_{{\rm cyclic},ij}=0$ is rejected 00 otherwise. The family of hypotheses $\{H_{0}^{(i,j)}\}_{1\leq i<j\leq K}$ is tested at the level $\alpha$ while controlling either the family–wise error rate (FWER) or the false discovery rate (FDR), see Stoica and Babu (2022) for more details on testing based approachs to model selection.
Remark 4.1
[…]
Let $\mathcal{C}_{0},\ldots,\mathcal{C}_{3}$ denote the collection of columns of $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ with $0,\ldots,3$ ticks respectively. Further let $\mathcal{C}_{1,\rm lin}$, $\mathcal{C}_{2,\rm lin}$ and $\mathcal{C}_{3,\rm lin}$ be sets of linearly independent elements in $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ respectively. Next let $\mathcal{C}_{2,\rm lin}^{+}$ be the set of elements in $\mathcal{C}_{2,\rm lin}$ which are not spanned by $\mathcal{C}_{3,\rm lin}$ and similarly let $\mathcal{C}_{1,\rm lin}^{+}$ be the set of elements in $\mathcal{C}_{1,\rm lin}$ which are not spanned by $\mathcal{C}_{3,\rm lin}\cup\mathcal{C}_{2,\rm lin}^{+}$. Further define a nested sequence of models $\mathcal{S}_{1}\subset\ldots\subset\mathcal{S}_{4}$ where $\mathcal{S}_{1}=\mathcal{M}$, $\mathcal{S}_{2}=\mathcal{M}\cup\mathcal{C}_{3,\rm lin}$, $\mathcal{S}_{3}=\mathcal{M}\cup\mathcal{C}_{3,\rm lin}\cup\mathcal{C}_{2,\rm lin}^{+}$, $\mathcal{S}_{4}=\mathcal{M}\cup\mathcal{C}_{3,\rm lin}\cup\mathcal{C}_{2,\rm lin}^{+}\cup\mathcal{C}_{1,\rm lin}^{+}$. Let $\mathcal{S}_{1}^{*},\ldots,\mathcal{S}_{4}^{*}$ denote the corresponding random sets constructed from $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}}$. Now:
Forward Tickwise Selection ($\rm FTBS$)
1:Test whether $H_{0}:\mathcal{S}=\mathcal{S}_{1}^{*}$ holds using Theorem 2.1. If the null is not rejected set $\widehat{\mathcal{S}}=\mathcal{S}_{1}^{*}$ and stop. Otherwise, set $a=2$.
2:Test whether $H_{0}:\mathcal{S}=\mathcal{S}_{a}^{*}$ holds using Theorem 4.3. If the null is not rejected set $\widehat{\mathcal{S}}=\mathcal{S}_{a}^{*}$ and stop. Otherwise,
3:Set $a=a+1$ and repeat Step 2.
Algorithm FTBS is computationally simple. It requires $\binom{K}{2}$ simple tests, cf. (35), followed by the construction of the linearly independent sets $\mathcal{S}_{1}^{*},\ldots,\mathcal{S}_{4}^{*}$ which is computationally a simple task. In addition FTBS requires the fitting and consequent evaluation, by testing for lack of fit, of at most four linear models, cf. Theorems 2.1 and 4.3.
Remark 4.2
The construction of the sets $\mathcal{S}_{1}^{*},\ldots,\mathcal{S}_{4}^{*}$ defined above and the structure of Algorithm FTBS are motivated by Theorems 2.1, 3.1 and 3.2. The key idea is that the more ticks the more likely a column is associated with a minimal model. Further observe that the starred sets are linearly independent from their predecessors.
The properties Algorithm $\rm FTBS$ are explored in Theorem 4.4 below.
Theorem 4.4
Fix $\alpha$ and let $\widehat{\mathcal{S}}$ denote the model selected by Algorithm $\rm FTBS$. If $n_{ij}\to\infty$, ${1\leq i<j\leq K}$ and $n_{ij}=O(n)$ then $\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))\to 1$. Moreover if the errors are sub–Gaussian RVs, then for some $C_{1},C_{2}\geq 0$ we have
$\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))\geq 1-C_{1}e^{(-C_{2}n)}.$
Additionally, if: $(i)$ all tests within $\rm FTBS$ are performed at a level $\alpha_{n}\rightarrow 0$; (ii) the power of rejecting any of the false nulls is $\pi_{n}\rightarrow 1$ when $(i)$ holds; and if $(iii)$ triads in the minimal model share at most index. Then,
$\mathbb{P}(\mathcal{S}=\widehat{\mathcal{S}})\to 1$ (36)
Theorem 4.4 show that Algorithm FTBS outputs a model $\widehat{\mathcal{S}}$ which is a superset of the true model $\mathcal{S}$. Moreover, if the minimal is unique and all test are applied with a small significance level then in large samples the correct model will be selected with probability one. Theorem 4.4 is established by showing that for large enough samples $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}}\supset\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ with high probability. Thus the sets $\mathcal{S}_{a}^{*}\supset\mathcal{S}_{a}$ which implies that all false null hypotheses tested in Step 2 of FTBS will be rejected. When the errors are subguassian the probability of not selecting $\widehat{\mathcal{S}}\supset\mathcal{S}$ is exponentially small. Incidentally, the assumption of subguassian errors can be replaced with subexponential errors. In the third part of Theorem 4.4 it is assumed that triads in the minimal model share at most index. This implies a unique minimal model and there is a one to one correspondence between the support and the span and guarantees that
$\mathbb{P}(\mathcal{S}=\widehat{\mathcal{S}})\to 1.$ (37)
If the triads in the minimal model share more than one index then (37) can not be guaranteed. For example, suppose that the unique minimal model is $\mathcal{C}_{\boldsymbol{\gamma}_{1}}=\{\boldsymbol{c}_{(1,4,5)},\boldsymbol{c}_{(2,5,6)},\boldsymbol{c}_{(3,4,6)}\}$. The model $\mathcal{C}_{\boldsymbol{\gamma}_{2}}=\{\boldsymbol{c}_{(1,4,5)},\boldsymbol{c}_{(2,5,6)},\boldsymbol{c}_{(3,4,6)},\boldsymbol{c}_{(4,5,6)}\}$ is such that ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$, but ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\neq{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. It can be verified that all triads in $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$ have $3$-ticks and therefore it is clear that FTBS will select $\boldsymbol{\gamma}_{2}$ and whence $\mathcal{S}\neq\widehat{\mathcal{S}}$.
Remark 4.3
An analogue of Theorem 4.4 applies also to FSTS except for the relation expressed in (36).
5 Simulations
In this section we compare four methods of model selection; two general purpose methods, namely forward stepwise regression (FSR) and the LASSO, cf. Hastie et al. (2009), and two specialized methods designed for the problem at hand, namely forward tickwise selection (FTBS) and step wise triad based selection (FSTS) described in Section 4.2. LASSO is implemented using cross validation, see the R-package glmnet. All other methods use multi–step testing based methods, in which all tests are carried out at the $0.05$ level. In FTBS, a Bonferroni corrected FWER is used for selecting the support.
Three experimental scenarios are considered, see Table 1. In Scenario I, the minimal model is unique, moreover the triads in the minimal model have three ticks whereas all other triads have either 00 or $1$ ticks in $\mathcal{C}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$. In Scenario II the minimal model is again unique, all triads in the minimal model have three ticks as does the triad $(4,5,6)$ which is not present in the minimal model. Finally, in Scenario III, the minimal model is not unique as $\boldsymbol{c}_{(1,2,3)}-\boldsymbol{c}_{(1,2,4)}=-\boldsymbol{c}_{(1,3,4)}+\boldsymbol{c}_{(2,3,4)}$, and thus no triad has three ticks.
Table 1: Simulation scenarios
Scenario $K$ $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$
I 6 $\boldsymbol{c}_{(1,2,3)}-\boldsymbol{c}_{(1,4,5)}$
II 6 $\boldsymbol{c}_{(1,4,5)}-\boldsymbol{c}_{(2,5,6)}+\boldsymbol{c}_{(3,4,6)}$
III 6 $\boldsymbol{c}_{(1,2,3)}-\boldsymbol{c}_{(1,2,4)}$
We consider only complete comparison graphs in which $n_{ij}=m$ for various values of $m$. For Scenarios I and II we report on ^$\widehat{\mathbb{P}}(\mathcal{S}\subset\widehat{\mathcal{S}})$, the empirical probability that the correct triads were selected, and on $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|\mathcal{S}|)$ the standardized mean size of the selected model. The closer $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|\mathcal{S}|)$ is to unity the better. For Scenario III we report on ^$\widehat{\mathbb{P}}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))$, the empirical probability that the space spanned by selected triads contains true $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$, and on $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|\mathcal{S}|)$ described above. All results are based on $1000$ replications and displayed in Tables 24.
Table 2: Outcomes of Scenario I and II with $K=6$
$m=5$ $m=10$ $m=20$
^$\widehat{\mathbb{P}}(\mathcal{S}\subset\widehat{\mathcal{S}})$ $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$$\widehat{\mathbb{P}}(\mathcal{S}\subset\widehat{\mathcal{S}})$ $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$$\widehat{\mathbb{P}}(\mathcal{S}\subset\widehat{\mathcal{S}})$ $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$
Scenario I FSR 0.842 1.30 0.987 1.33 1 1.31
LASSO 0.976 3.20 1 3.21 1 3.21
[…]
FSTS 0.209 0.66 0.764 0.97 0.993 1.03
Scenario II FSR 0.738 1.17 0.973 1.17 0.998 1.18
LASSO 0.971 2.77 1 2.63 1 2.55
[…]
FSTS 0.066 0.71 0.568 1.01 0.986 1.08
Table 3: Outcomes Scenario II with $K=10$
$m=5$ $m=10$ $m=20$
^$\widehat{\mathbb{P}}(\mathcal{S}\subset\widehat{\mathcal{S}})$ $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$$\widehat{\mathbb{P}}(\mathcal{S}\subset\widehat{\mathcal{S}})$ $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$$\widehat{\mathbb{P}}(\mathcal{S}\subset\widehat{\mathcal{S}})$ $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$
Scenario II FSR 0.607 1.97 0.962 1.96 0.999 1.90
LASSO 0.855 4.42 0.997 4.63 1 4.43
[…]
FSTS 0.006 0.49 0.202 0.81 0.896 1.03
Table 4: Outcomes Scenario III with $K=6$
$m=5$ $m=10$ $m=20$
^$\widehat{\mathbb{P}}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))$ $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$$\widehat{\mathbb{P}}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))$ $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$$\widehat{\mathbb{P}}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))$ $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$
FSR 0.743 1.24 0.982 1.27 1 1.26
LASSO 0.784 3.99 0.987 4.44 1 4.54
[…]
FSTS 0.022 0.61 0.29 1.08 0.926 1.53
First observe that when $m$ is large $\widehat{\mathcal{S}}\supset\mathcal{S}$ with high probability. The coverage probability of the LASSO, i.e., ^$\widehat{\mathbb{P}}(\mathcal{S}\subset\widehat{\mathcal{S}})$, is generally higher than all other methods, however it performs very poorly in terms of the average size of the selected model. For example under Scenario I and $m=20$ the LASSO outputs models of average size $3.21$ when the true minimal model is of size $1$. Other methods output models of size close to $1$. This is not surprising as the LASSO shrinks towards a minimum $L_{1}$ model. Although FSR has lower coverage than the LASSO, considering both the coverage and model size, FSR’s overall performance seems better than that of LASSO. This is probably due to the fact that FSR selects one covariate at a time while the LASSO selects all covariates simultaneously. Among the methods proposed in Section 4.2, the coverage of FSTS as well as its expected model size are typically lower than those of FTBS. However, the gain in model size is not balanced by the loss in coverage in particular for large values of $m$. For example in Table 4 for $m=20$ the coverage of FTBS is $0.993$ whereas its model size is $1.57$, the corresponding figures for FSTS are $0.926$ and $1.53$. Thus a loss of $7\%$ in coverage is offset be $0.04\%$ loss is relative model length. Thus we prefer FTBS, whose theoretical properties are explored in Theorem 4.4, to FSTS. Finally we compare FSR and FTBS. For a unique minimal model they perform similarly. For lower values of $K$ and non–unique minimal models FTBS tends to have larger model size. However, for large values of $K$, the model size of FTBS is smaller. For example, see Table 3, with $K=10$ and $m=20$ FTBS relative model size is $1.54$ while that of FSR is $1.90$. Furthermore the computation costs associated with FTBS are usually lower than those of FSR. Therefore it seems that FTBS performs better than FSR and all other methods.
Remark 5.1
In some settings FTBS can be used as a screening tool to which one can apply either FSR or FSTS. Concretely, one may use the tick–table to identify potential triads to be included in the selected model and then apply either FSR or FSTS to this reduced collection of triads. This approach is exemplified in the following section.
6 Illustrative example
Sporting competitions are an important source of interest, study, and entertainment for the general public as well an active arena to which statistical analyses and ranking methods are passionately applied. To date, methods employed for ranking sports teams have assumed linear transitive models (Sinuany-Stern 1988, Cassady et al. 2005, Langville and Meyer 2012, Barrow et al. 2013). Alternatives allowing for cyclicalities and intransitivities were simply not available.
We apply the framework developed in this communication to analyze data from the England Premier League (EPL), one of the foremost soccer/football leagues in the world. We focus on the 2022–23 season in which $20$ teams competed. The EPL operates on a double round-robin format, i.e., each team plays every other team twice, once on each home field. Thus $38$ matches are played by each team resulting in a total of $380$ matches. In the following, the outcome of each game is the difference between the expected number of goals (xG) scored, c.f., Carey and Worville (2024). The second and third columns of Table 5 report team names and their positions, i.e., their official ranking as it appears on the EPL website www.premierleague.com. Teams are listed alphabetically and the first column of Table 5 assigns each team a index number by which we shall refer to them.
Table 5: Names, positions, merits ($\widehat{\mu}_{s,i}$), Merit–Rank ($r(\widehat{\mu}_{s,i})$), Dominance–Score ($\mu_{i}^{**}$) and Dominance–Rank($r(\mu_{i}^{**})$) of teams in England Premier League (EPL) season 2022-23
Index Name Position $\widehat{\mu}_{s,i}$ $r(\widehat{\mu}_{s,i})$ $\mu_{i}^{**}$ $r(\mu_{i}^{**})$
1 Arsenal 2 0.7325 3 17 2
2 Aston Villa 7 -0.0575 9 10 11
3 Bournemouth 15 -0.635 20 4 19
4 Brentford 9 0.165 8 13 6
5 Brighton 6 0.585 4 15 5
6 Chelsea 12 -0.07 10 9 12
7 Crystal Palace 11 -0.2225 12 6 14
8 Everton 17 -0.505 16 5 15
9 Fulham 10 -0.4325 14 7 13
10 Leeds United 19 -0.49 15 4 17
11 Leicester City 18 -0.3225 13 10 8.5
12 Liverpool 5 0.54 5 16 3
13 Manchester City 1 1.16 1 18 1
14 Manchester Utd 3 0.43 6 12 7
15 Newcastle Utd 4 0.805 2 15 4
16 Nott’ham Forest 16 -0.6175 19 4 18
17 Southampton 20 -0.5825 17 5 16
18 Tottenham 8 0.195 7 10 8.5
19 West Ham 14 -0.095 11 10 10
20 Wolves 13 -0.5825 18 3 20
First, ignoring possible cyclicalities, the data is analyzed assuming the reduced, linearly transitive model, i.e., assuming the relation (3) holds. The fourth column of Table 5 reports the estimated merits and the fifth column on the merit based rankings. The rank of the $i^{th}$ team is $r(\widehat{\mu}_{i})=\sum_{j\neq i}\mathbb{I}(\widehat{\mu}_{i}>\widehat{\mu}_{j})$. For example Manchester City is ranked first with merit of $2.32$. Further note that the merit–based ranking coincides with both the row–sum ranking method and the method proposed by Saari (2014) and Saari (2021), cf. Equation (25).
Next, we apply a test for lack of fit, cf. Theorem 2.1, to assess whether the reduced model fits the data. We find that the p-value associated with this test is $<10^{-3}$ indicating a strong deviation from the null.
Remark 6.1
It is interesting to note that $\|\widehat{\operatorname{\boldsymbol{\nu}}}_{\rm linear}\|_{2}=10.70$ and $\|\widehat{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}\|_{2}=8.59$, i.e., the norms of the cyclical and linear components are comparable. This seem to suggest that the cyclic component is significant. To investigate further, consider the following Monte–Carlo experiment in which the outcomes of the 2022-23 EPL season were simulated $10^{5}$ times assuming transitivity, i.e., by setting $\operatorname{\boldsymbol{\nu}}=\widehat{\operatorname{\boldsymbol{\nu}}}_{\rm linear}$ and assuming IID $\mathcal{N}(0,\sigma^{2})$ errors where $\sigma^{2}=\widehat{\sigma}^{2}$ is defined in Proposition 4.1. In each run ${\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}$ is estimated. The mean norm of the estimated cyclical component is $5.65$ with a standard deviation of $0.31$. The estimated value from the data, i.e., $8.59$, is larger than the largest simulated value. This fact further suggests that the reduced model does not fit these data.
Next we compute and analyze $\widehat{\operatorname{\boldsymbol{\nu}}}$ with the goal of selecting the best fitting model for the EPL. First, a test for cyclicality, cf. Proposition 4.1, is applied to each of the $1140$ triads. Under the null, $57$ triads are expected to be associated with a p–value smaller than $0.05$; in fact there are $195$ such triads, providing additional evidence for a departure from linear transitivity. Table 6 appearing in the Supplementary Materials lists the $100$ most significant triads. The most significant triad with a p-value of $10^{-4}$ is $(2,5,19)$ for which $(\widehat{\nu}_{2,5},\widehat{\nu}_{2,19},\widehat{\nu}_{5,19})=(1.15,-0.35,2.05)$ so $2\succ 5\succ 19\succ 2$.
Moving on to model selection we first directly apply the methods compared in Section 5. We find that: (a) LASSO, which uses cross–validation, does not select any triad; (b) FSR (with $\alpha=0.05$) selects a model with $16$ triads; (c) FSTS selects the model with the two most significant triads $(2,5,19)$ and $(6,11,12)$; $(d)$ Applying FTBS with a Bonferonni–corrected family wise error rate of $0.1$ we find that the only element of $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ which is different from 00 is $\nu_{{\rm cyclic},25}$. Consequently the tick–table $\operatorname{\boldsymbol{C}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}}$ has no $3$ or $2$-tick columns and $18$ triads have $1$-tick. Since FTBS selects variables in blocks, a naive application of FTBS will select a model with $18$ triads.
Clearly FTBS does not operate satisfactorily in this example. This is likely due to the fact that $m=2$ so the tests in (35) are under–powered and therefore do not detect some pairs in the support of the true minimal model. The correct estimation of the support would have likely lead to the detection of the true three tick triads. The performance of FTBS can be modified and enhanced in various ways. In essence, in order to lower the expected model size of FTBS the blocks (of columns) added at each stage of FTBS need to be reduced. One way of doing so is to use FTBS as a screening tool which identifies potential triads to be further selected. Methods which select one triad at a time can be then applied to this collection of triads. Thus FTBS–FSTS or FTBS–FSR hybrids are possible.
When applied to the EPL2023 data set FTBS–FSR selects a model with a single triad $(2,5,19)$, whereas the FTBS–FSTS selects a model with two triads $(2,5,19),(2,5,13)$. FTBS–FSR is preferred by BIC whereas FTBS–FSTS is preferred by AIC although the differences are rather small. For the simplicity of the exposition we shall refer to the first of these models as the selected (or final) model.
The final preference profile is estimated by
${\widehat{\operatorname{\boldsymbol{\nu}}}}_{s}=\operatorname{\boldsymbol{B}}{\widehat{\operatorname{\boldsymbol{\mu}}}}_{s}+\boldsymbol{c}_{(2,5,19)}{\widehat{{\gamma}}}_{s}$
where $\widehat{{\operatorname{\boldsymbol{\mu}}}}_{s}$ and $\widehat{\boldsymbol{\gamma}}_{s}$ are estimated as described in Theorem 4.2. The full preference profile ${\widehat{\operatorname{\boldsymbol{\nu}}}}_{s}$ is displayed in Table 7 in the Supplementary Materials. The merits associated with the selected model are displayed in column four of Table 5. Further note that in the full model $(\widehat{\nu}_{2,5},\widehat{\nu}_{2,19},\widehat{\nu}_{5,19})=(1.15,-0.35,2.05)$ and in the selected model $({\widehat{\nu}}_{s,2,5},{\widehat{\nu}}_{s,2,19},{\widehat{\nu}}_{s,5,19})=(0.54,-1.14,1.86)$ and moreover $\widehat{{\gamma}}_{s,2,5,19}=1.18$. Thus the triad $(2,5,19)$ exhibits cyclicality even after fitting the selected model. This indicates that the true preference profile among the teams $\{2,5,19\}$ is truly cyclical. It is also interesting to note that a cyclical triad may include teams in very different positions. For example the triad $(2,5,19)$ comprised of Aston Villa, Brighton and West Ham are located in positions $7$, $6$ and $14$. The cycle among them induces a a complicated web of preference in the EPL.
A complete ranking in the presence of true cyclicalities is challenging. Columns six and seven in Table 5 display the dominance scores and ranks for all teams in the EPL, respectively. The dominance score of $\{2,11,18,19\}$ are initially equal so a tie braking procedure, see Section 3.3, is used. Nevertheless, the resulting ranking can not be easily accepted.
Figure 1 depicts the preference profile among the aforementioned teams. An arrow from $i$ pointing at $j$ means that $i\succ j$. The relation between teams $11$ and $18$ is of equality. Figure 1 shows two cycles, i.e., $2\succ 18\sim 11\succ 2$ and $2\succ 18\succ 19\succ 2$. Any global method of ranking will not honor such cycles and is therefore open to challeng. One way of addressing and partially resolving such issues is to define a ranking with respect to each item $i$. Define the sets: (i) $A_{i}=\{j:\widehat{\nu}_{ij}<0\}$; (ii) $B_{i}=\{j:\widehat{\nu}_{ij}>0\}$; and (iii) $E_{i}=\{j:\widehat{\nu}_{ij}=0\}$. Items in $A_{i}$ and $B_{i}$ are ranked above and below $i$ respectively, and items in $E_{i}$ are equivalent to $i$. See Table 8 in Supplementary material for a rank sets of the leading teams five teams in the EPL.
2222111111111818181819191919
Figure 1: Preferences among teams with indices $\{2,11,18,19\}$
Remark 6.2
There are some differences between the FTSB–FSR selected model, analyzed above, and the FTSB–FSTS selected model (which we do not analyze for brevity) although these are generally minor.
7 Discussion
This paper presents a principled and unified approach to the modelling and analysis of PCD in the presence of cyclicality and intransitively. We start by contrasting the reduced, transitive model, indexed by $\operatorname{\boldsymbol{\mu}}\in\mathcal{M}\equiv\mathbb{R}^{K}$ with the full model indexed by $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}\equiv\mathbb{R}^{K(K-1)/2}$ which imposes no restrictions on the preference profile. Clearly models indexed by $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ capture a much broader range of preference relations compared to models indexed by $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$. Modelling cyclicalities allows for a more nuanced understanding of the preference profile. In some settings, e.g., our illustrative example, the analysis provided will naturally result further efforts attepmting to understand and explain the pattren of observed cyclicalities.
The key to modelling is the decomposition of the parameter space $\mathcal{N}$ into a linear transitive subspace and cyclical subspace, i.e., expressing $\operatorname{\boldsymbol{\nu}}$ as (20). This form captures all possible preferences relations and allows for a compact representation in situations in which cyclicalities are limited to a subset of the items, that is when the minimal model is of a relatively small size. It is worth noting that ANOVA type decomposition of the form
$\nu_{ij}=\mu_{i}-\mu_{j}+\gamma_{ij}$
where $\sum_{j}\gamma_{ij}=0$ for all $i$ has been proposed by Scheffe (1952) and studied by several other authors. Although the ANOVA decomposition yields a saturated model, i.e., any $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ can be consistently estimated, it does not capitalize on the intrinsic structure of PCD and lacks the algebraic interpretation of the current model, i.e., Scheffe’s merits and interaction parameters do not correspond to linear transitivity and cyclicality. Saari (2014), expanded on in Saari (2021), was the first to obtain bases for $\mathcal{M}$ and $\mathcal{N}\setminus\mathcal{M}$. His objective was a obtaining a global ranking rather than the modeling of the preference relation $\operatorname{\boldsymbol{\nu}}$. For example, he did not attempt a decomposition of $\operatorname{\boldsymbol{\nu}}$ nor did he consider the notions of a minimal model or the tick table $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$. In addition he did not consider or discuss incomplete PCGs. Most importantly, Saari (2014) and Saari (2021) did not develop any tools for conducting statistical inference for such models.
This paper shows that in scenarios which are likely in practice, cf. Theorems 3.1 and 3.2, PCGs can be analyzed satisfactorily, that is, the sources of cyclicalities and intransitivities can be fully identified, estimated and addressed within the framework of minimal models. We propose a new tailor–made test–based model selection method which we refer to as FTSB. Simulations indicate that the proposed method, which is computationally efficient, performs well in practice especially for larger graphs. Our illustrative example, which is of independent interest, shows that model selection may be a challenge in situations in the number of comparisons between any two items is small. This of course, is not surprising.
Many open problems, both theoretical and applied, have not been explored. To start, our simulation study has not investigated many interesting and important issues. For example, we have not explored very large comparison graphs with either sparse or dense degree of cyclicality. Nor have we attempted to optimize FTSB. Recall that FTSB is a block–wise method, i.e., it selects blocks of columns rather than a single columns at a time as is done by FSR and FSTS. It is clear that by reducing the block size by partitioning existing block into sub–blocks (using some criteria such as their level of significance of their individual components) one could reduce the expected size of the model, i.e., $\mathbb{E}(|\widehat{\mathcal{S}}|/|\mathcal{S}|)$ without compromising on the probability of correct selection. One approach to this problem is to combine FTSB with methods such as FSR or FSTS as demonstrated in Section 6. Applying model selection criteria such as AIC or BIC (Clasken and Hojrt, 2008) can further improve the selected model. Furthermore, we believe that by optimizing the significance level at which tests are performed at various stages one may be able to attain additional benefits.
We have applied the proposed methodology to the ranking of teams from the EPL. We have found that linear transitive model does not fit these data. Under the selected model some triads exhibit cyclicality, a fact that complicates ranking. We are eager to apply the proposed methodology to a wide variety of other sports, chess, and consumer preferences and to a variety of relevant data sets explored by other researchers. Such results will be of great interest.
Numerous theoretical challenges remain. For example, we have not addressed the issue of long cycles, cf. Equation (5). In general a cycle among the items $i_{1},\ldots,i_{r}$ can be expressed by the vector $\boldsymbol{c}_{\boldsymbol{\lambda}}\in\mathcal{N}\setminus\,\mathcal{M}$ where
$c_{ij}({\boldsymbol{\lambda}})=\mathbb{I}((i\succ j)\in\boldsymbol{\lambda})-\mathbb{I}((j\succ i)\in\boldsymbol{\lambda}),$
and $\boldsymbol{\lambda}=(i_{1},i_{2},\ldots,i_{r})$. Note that $\boldsymbol{c}_{\boldsymbol{\lambda}}$ can be expressed as a linear combination of vectors in $\{\boldsymbol{c}_{(i,j,k)}\}_{1\leq i<j<k\leq K}$. For example, the longest possible cycle, i.e., $1\succ 2\succ\ldots\succ K\succ 1$, associated with the vector $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,\ldots,K)}$ can be expressed the sum of $K-2$ linearly independent triads, e.g., $\boldsymbol{c}_{(1,\ldots,K)}=\sum_{j=2}^{K-1}\boldsymbol{c}_{(1,j,j+1)}$. There are, of course, many other ways of writing $\boldsymbol{c}_{(1,\ldots,K)}$. Thus cycles are not associated with a unique minimal model. Nevertheless, it is clear that long cycles can be identified from the data and incorporated into our modelling framework. In this communication we have focused on complete PCGs. In practice, however, many PCGs are incomplete. Some aspects thereof, mainly pertaining to modelling, have been discussed in Section 3.2 but a statistical analysis has not been pursued here. We plan to do so elsewhere. Another important area requiring further research is the relationship between the proposed framework and the elucidation of transitivity relations. For example can tests for and against weak and/or strong transitivity relations be developed? can estimators of the model parameters designed to honor such relations be developed? what properties are to be expected. Finally, the approach outlined here can be readily extended to binary and other types of paired comparison data. For example, in binary PCD the probability that item $i$ is preferred over item $j$ is typically modelled as $p_{ij}=F(\mu_{i}-\mu_{j})$ where $F$ is some distribution function. If $F$ is the logistic distribution function then the resulting model is the well known Bradley–Terry (BTL) model (Bradley and Terry 1952; Wu et al. 2022). The classic BTL model can be extended to model cyclicality if it is rewritten as $p_{ij}=F(\nu_{ij})$ where the vector $\operatorname{\boldsymbol{\nu}}$ is modeled as in (20). Thus our framework can be readily extended beyond cardinal PCD.
Acknowledgments
The work of Rahul Singh was conducted while a post doctoral fellow at the University of Haifa. The work of Ori Davidov was partially supported by the Israeli Science Foundation Grant No. 2200/22 and gratefully acknowledged.
Appendix
Appendix A Proofs
Proof of Theorem 2.1:
Proof. Let $\boldsymbol{U}_{n}$ be the random vector of dimension $|\operatorname{\boldsymbol{\nu}}|$ whose $(i,j)^{th}$ components is $U_{ij}=\sqrt{n_{ij}}(S_{ij}/n_{ij}-(\widehat{\mu}_{i}-\widehat{\mu}_{j}))$ when $(i,j)\in\mathcal{E}_{n}$ and 00 otherwise. Note that
$\widehat{\mu}_{i}-\widehat{\mu}_{j}=(\boldsymbol{e}_{i}-\boldsymbol{e}_{j})^{\top}\widehat{\operatorname{\boldsymbol{\mu}}}=(\boldsymbol{e}_{i}-\boldsymbol{e}_{j})^{\top}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{S}}$
where $\operatorname{\boldsymbol{S}}$, defined in Section 2.1, is a function of all $S_{ij}$’s. Let $\widetilde{\operatorname{\boldsymbol{S}}}_{\rm ALL}$ be the $|\operatorname{\boldsymbol{\nu}}|$ vector whose $(i,j)^{th}$ element is equal to $S_{ij}/n_{ij}$’s if $(i,j)\in\mathcal{E}_{n}$ and 00 otherwise. Now, using some algebra it can be shown that
$\boldsymbol{U}_{n}=\boldsymbol{D}^{1/2}(\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{D}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{D})\widetilde{\operatorname{\boldsymbol{S}}}_{\rm ALL},$
where $\boldsymbol{D}$ is a $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ diagonal matrix whose $(i,j)^{th}$ diagonal element is $n_{ij}$ if $(i,j)\in\mathcal{E}_{n}$ and 00 otherwise. Furthernote,
$\boldsymbol{U}_{n}=$ $~{}\boldsymbol{D}^{1/2}(\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{D}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{D})(\boldsymbol{D}^{+})^{1/2}\boldsymbol{D}^{1/2}\widetilde{\operatorname{\boldsymbol{S}}}_{\rm ALL}$
$=$ $~{}(\boldsymbol{D}/n)^{1/2}(\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}(\boldsymbol{D}/n)\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}(\boldsymbol{D}/n))((\boldsymbol{D}/n)^{+})^{1/2}\boldsymbol{D}^{1/2}\widetilde{\operatorname{\boldsymbol{S}}}_{\rm ALL}$
$=$ $~{}\boldsymbol{M}\boldsymbol{D}^{1/2}\widetilde{\operatorname{\boldsymbol{S}}}_{\rm ALL},$
where $\boldsymbol{M}=\boldsymbol{D}^{1/2}(\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{D}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{D})(\boldsymbol{D}^{+})^{1/2}$. Condition 2.3 guarantees that as $n\rightarrow\infty$ we have $\boldsymbol{M}\to\boldsymbol{\Xi}^{1/2}(\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi})(\boldsymbol{\Xi}^{+})^{1/2}$. In addition by the CLT we have $\boldsymbol{D}^{1/2}(\widetilde{\operatorname{\boldsymbol{S}}}_{\rm ALL}-\operatorname{\boldsymbol{\nu}})\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\sigma^{2}\boldsymbol{H})$, where $\boldsymbol{H}$ is a $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ diagonal matrix whose diagonal elements are $1$ if $(i,j)\in\mathcal{E}_{n}$ and 00 otherwise. It follows by Slutzky’s Theorem that $\boldsymbol{U}_{n}\Rightarrow\boldsymbol{U}$ where $\boldsymbol{U}$ follows a $\mathcal{N}_{|\operatorname{\boldsymbol{\nu}}|}(0,\boldsymbol{\Psi})$ with
$\boldsymbol{\Psi}=\sigma^{2}\boldsymbol{\Xi}^{1/2}(\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi})\boldsymbol{\Xi}^{+}(\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi})^{\top}\boldsymbol{\Xi}^{1/2}.$
Next observe that
$R_{n}=\sum_{(i,j)\in\mathcal{E}}n_{ij}(\overline{S}_{ij}-(\widehat{\mu}_{i}-\widehat{\mu}_{j}))^{2}=\boldsymbol{U}_{n}^{\top}\boldsymbol{U}_{n}\Rightarrow\boldsymbol{U}^{\top}\boldsymbol{U}.$
[…]
Remark A.1
Algebraically it is difficult to find rank of matrix $\boldsymbol{\Psi}$. However, the rank of $\boldsymbol{\Psi}$ can be computed using a statistical argument. To do so suppose that the errors $\epsilon_{ijk}$ are IID $\mathcal{N}(0,\sigma^{2})$ RVs. It is well known that the LRT for testing $H_{0}:\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ versus $H_{1}:\operatorname{\boldsymbol{\nu}}\notin\mathcal{M}$ follows, under the null a chi–square distribution with $|\mathcal{E}_{n}|-(K-1)$ degrees of freedom. Furthermore, the corresponding Wald test is of the form $\boldsymbol{U}_{n}^{\top}{\boldsymbol{\Psi}}_{n}^{+}\boldsymbol{U}_{n}$ and has the same asymptotic distribution as the LRT and consequently by Brown and Cribari-Neto (1992) the matrix ${\boldsymbol{\Psi}}_{n}$ must have rank $|\mathcal{E}_{n}|-(K-1)$ for all large $n$ and therefore also in the limit. This established the claim above.
Proof of Theorem 2.3:
Proof. By assumption $\operatorname{\boldsymbol{Y}}\sim{\mathcal{N}}_{n}(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}},\sigma^{2}\boldsymbol{I})$, where matrix $\operatorname{\boldsymbol{L}}$ is defined in Section 4.1. Moreover if $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ then $\nu_{ij}=\mu_{i}-\mu_{j}$ and for brevity will be denoted it by $\operatorname{\boldsymbol{\nu}}_{\mu}$. Let $\boldsymbol{f}_{1}$ and $\boldsymbol{f}_{2}$ be the densities corresponding to $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ and $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$, respectively. The Kullback–Leibler divergence between models indexed by $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ and $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ is
$\mathrm{KL}(\mathcal{N},\mathcal{M})=$ $~{}\mathbb{E}_{\boldsymbol{f}_{1}}(\log{\boldsymbol{f}_{1}}-\log{\boldsymbol{f}_{2}})=\frac{1}{2\sigma^{2}}\mathbb{E}_{\boldsymbol{f}_{1}}(-(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}})^{\top}(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}})+(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu})^{\top}(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu}))$
$=$ $~{}\frac{1}{2\sigma^{2}}\mathbb{E}_{\boldsymbol{f}_{1}}(2(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}})^{\top}(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu})+(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu})^{\top}(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu}))$
$=$ $~{}\frac{1}{2\sigma^{2}}(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{\nu}}_{\mu})^{\top}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{L}}(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{\nu}}_{\mu})=\frac{1}{2\sigma^{2}}\sum_{1\leq i<j\leq K}n_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}.$
It follows that $\mathrm{KL}(\mathcal{N},\mathcal{M})$ is minimized by
${\operatorname{\boldsymbol{\mu}}}=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\mathrm{KL}(\mathcal{N},\mathcal{M})=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\sum_{1\leq i<j\leq K}n_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}.$
Next observe that
$\widehat{\operatorname{\boldsymbol{\mu}}}=$ $~{}\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\frac{1}{n}\sum_{1\leq i<j\leq K}\sum_{t=1}^{n_{ij}}(Y_{ijt}-\left(\mu_{i}-\mu_{j}\right))^{2}=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\frac{1}{n}\sum_{1\leq i<j\leq K}\sum_{t=1}^{n_{ij}}(\nu_{ij}-(\mu_{i}-\mu_{j})+\epsilon_{ijt})^{2}$
$=$ $~{}\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\sum_{1\leq i<j\leq K}\frac{n_{ij}}{n}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}-2\sum_{1\leq i<j\leq K}\frac{n_{ij}}{n}(\nu_{ij}-(\mu_{i}-\mu_{j}))\frac{1}{n_{ij}}\sum_{t=1}^{n_{ij}}\epsilon_{ijt}$
$~{}+\frac{1}{n}\sum_{1\leq i<j\leq K}\sum_{t=1}^{n_{ij}}\epsilon_{ijt}^{2}.$
By the law of large numbers the third term above converges to $\sigma^{2}$ which is independent of $\operatorname{\boldsymbol{\mu}}$. Condition 2.2 implies that $n_{ij}/n=\theta_{ij}+o(1)$ and consequently we can replace $n_{ij}/n$ appearing in the first term above with $\theta_{ij}+o(1)$. If $\theta_{ij}=0$ then the second term is $o_{p}(1)$ whereas if $\theta_{ij}>0$ then ${1}/{n_{ij}}\sum_{t=1}^{n_{ij}}\epsilon_{ijt}\to 0$; hence in both cases the second term is $o_{p}(1)$. Putting it all together we find that
$\widehat{\operatorname{\boldsymbol{\mu}}}=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\sum_{1\leq i<j\leq K}\theta_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}+o_{p}(1).$
establishing (12). Further note that
$\sum_{1\leq i<j\leq K}\theta_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}=(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}})^{\top}\boldsymbol{\Xi}(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}).$ (38)
Condition 2.3 implies that there exist a tree $\mathcal{T}$ such that $\theta_{ij}>0$ for all $(i,j)\in\mathcal{T}$; clearly a unique minimizer exist on $\mathcal{T}$ and consequently on every $\mathcal{G}\supset\mathcal{T}$. For a finite sample analogue see Theorem 2.1 in Singh et al. (2024). Further note that if $\operatorname{\boldsymbol{\nu}}\in\mathcal{M}$ then (12) is minimized by choosing $\operatorname{\boldsymbol{\mu}}$ such that $\nu_{ij}=\mu_{i}-\mu_{j}$ in which case (38) is identically zero for all possible choices of $\boldsymbol{\Theta}$. If however $\operatorname{\boldsymbol{\nu}}\notin\mathcal{M}$ then (38) is always a function of $\boldsymbol{\Theta}$. The Larnagian associated with (12) is
$L=(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}})^{\top}\boldsymbol{\Xi}(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}})+\lambda\boldsymbol{v}^{\top}\operatorname{\boldsymbol{\mu}},$
where $\operatorname{\boldsymbol{1}}1^{\top}\boldsymbol{v}\neq 0$. Note that gradient of $L$ is $\nabla L=(\nabla_{\operatorname{\boldsymbol{\mu}}}L,\nabla_{\lambda}L)$ where
$\nabla_{\operatorname{\boldsymbol{\mu}}}L=-\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}+\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}+\lambda\boldsymbol{v}/2$ and $\nabla_{\lambda}L=\boldsymbol{v}^{\top}\operatorname{\boldsymbol{\mu}}$. Following the proof of Theorem 2.1 in Singh et al. (2024) it can be shown that
$\widehat{\operatorname{\boldsymbol{\mu}}}=(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}+[\boldsymbol{I}-(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}]\boldsymbol{\alpha},$
where $\boldsymbol{\alpha}\in\mathbb{R}^{K}$ is arbitrary which reduces to
$\widehat{\operatorname{\boldsymbol{\mu}}}=(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}$ (39)
when $\boldsymbol{v}=\operatorname{\boldsymbol{1}}1$.
Proof of Proposition 2.1:
Proof. Under the stated conditions
$U_{ij}=\sqrt{n_{ij}}(\overline{S}_{ij}-(\widehat{\mu}_{i}-\widehat{\mu}_{j}))=\sqrt{n_{ij}}(\overline{S}_{ij}-(\widehat{\mu}_{i}-\widehat{\mu}_{j}-n^{-1/2}{\delta}_{ij})+n^{-1/2}{\delta}_{ij})=U_{ij}^{*}+\sqrt{\frac{n_{ij}}{n}}\,{\delta}_{ij},$
where $U_{ij}^{*}=\sqrt{n_{ij}}(\overline{S}_{ij}-(\widehat{\mu}_{i}-\widehat{\mu}_{j}-n^{-1/2}{\delta}_{ij}))$. Consequently $\boldsymbol{U}_{n}\Rightarrow\boldsymbol{U}$ where $\boldsymbol{U}$ is a $\mathcal{N}_{|\operatorname{\boldsymbol{\nu}}|}(\boldsymbol{\Xi}^{1/2}\,\boldsymbol{\delta},\boldsymbol{\Psi})$ RV. It follows that
[…]
$=$ $~{}\boldsymbol{V}^{\top}\boldsymbol{\Lambda}\boldsymbol{V},$
[…]
$R_{n}=\sum_{i=1}^{r}\lambda_{i}(Z_{i}+\gamma_{i})^{2},$
where $\gamma_{i}$ is the element of $\boldsymbol{\gamma}$ corresponding with the $i^{th}$ nonzero diagonal element of $\boldsymbol{\Lambda}$ denoted by $\lambda_{i}$.
Proof of Theorem 3.1:
Proof. By Definition if 3.1 $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ and consequently $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\cap{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Next suppose that ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. It immediately follows that $\mathcal{C}_{\boldsymbol{\gamma}_{1}}$ and $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$ have same set of items, say $\{1,\ldots,L\}$. For if $i$ is present in $\mathcal{C}_{\boldsymbol{\gamma}_{1}}$ and absent in $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$ then there exist a triad $\boldsymbol{c}_{(i,j,k)}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$ and $\boldsymbol{c}_{(i,j,k)}\notin{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ for all $(j,k)\notin{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Hence ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\neq{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Next, suppose that $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$. Then there exists a $k\leq L$ such that $\boldsymbol{c}_{(i,j,k)}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$ and consequently $\boldsymbol{c}_{(i,j,k)}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Therefore if the spans are equal so are the supports.
Let $\boldsymbol{\gamma}_{1}$ and $\boldsymbol{\gamma}_{2}$ be two minimal models. By assumption $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$ implies $\nu_{{\rm cyclic},ij}\neq 0$, therefore it also follows that $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ and consequently ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\subseteq{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Now if ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\subset{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$, there exists some $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ and $(i,j)\notin{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$. Thus there is some triad $(i,j,t)$ which belongs to $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$ but not $\mathcal{C}_{\boldsymbol{\gamma}_{1}}$ consequently $s(\boldsymbol{\gamma}_{2})>s(\boldsymbol{\gamma}_{1})$ and therefore $\boldsymbol{\gamma}_{2}$ can not be a minimal model. Thus ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ and completing the proof of third part.
Finally suppose that any two triads in a minimal model $\boldsymbol{\gamma}$ spanned by $\mathcal{C}_{\boldsymbol{\gamma}}$ share at most one index, i.e., if $\boldsymbol{c}_{(i,j,k)},\boldsymbol{c}_{(u,v,w)}\in\mathcal{C}_{\boldsymbol{\gamma}}$ then |$|\{i,j,k\}\cap\{u,v,w\}|\leq 1$. Using (17) it is easily verified that such triads are associated with orthogonal columns of $\operatorname{\boldsymbol{C}}$. It is also immediate that if $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ then $(i,j)$ belongs to a single triad in $\mathcal{C}_{\boldsymbol{\gamma}}$. By definition
$\operatorname{\boldsymbol{\nu}}_{{\rm cyclic}}=\sum_{\boldsymbol{c}_{(i,j,k)}\in\mathcal{C}_{\boldsymbol{\gamma}}}\boldsymbol{c}_{(i,j,k)}\boldsymbol{\gamma}_{ijk}.$ (40)
It is clear that if $(s,t)\notin{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ then $\nu_{{\rm cyclic},st}=0$, whereas if $(s,t)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ then $(s,t)$ belongs to a single triad $(i,j,k)$, say, and $\nu_{{\rm cyclic},st}={c}_{(i,j,k),st}{\gamma}_{ijk}\neq 0$. We conclude that $\nu_{{\rm cyclic},ij}\neq 0$ if and only if $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$. Also $\sum_{1\leq i<j\leq K}\mathbb{I}(\nu_{{\rm cyclic},ij}\neq 0)=3s$, where $s$ is size of the model. Next suppose there exist a minimal model $\boldsymbol{\gamma}^{\prime}$ for which $\boldsymbol{\gamma}\neq\boldsymbol{\gamma}^{\prime}$.
It follows from the discussion above that if $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ij}\neq 0$ then $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}^{\prime}})$. These pairs, however, are in ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ and therefore ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})\subseteq{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}^{\prime}})$. Remark A.2 implies that $s({\boldsymbol{\gamma}})<s({\boldsymbol{\gamma}^{\prime}})$, i.e., the size of the model $\boldsymbol{\gamma}^{\prime}$ is larger than the size of the model $\boldsymbol{\gamma}$. This leads to a contradiction and the conclusion that $\boldsymbol{\gamma}=\boldsymbol{\gamma}^{\prime}$.
Remark A.2
If a minimal model has to be constructed using a given support, then we seek a minimum number of triads using the support, such that these triads are linearly independent. That is, if $(i,j),(i,k),(j,k)$ are present in the support then $\boldsymbol{c}_{(i,j,k)}$ is a viable triad and we find such independent triads until support is covered. Now if triads in a minimal model $\boldsymbol{\gamma}$ share at most one index then any comparison appears in only one triad of $\mathcal{C}_{\boldsymbol{\gamma}}$. Precisely there is a one-to-one mapping between comparisons and triads in $\boldsymbol{\gamma}$. Consequently only triads in $\mathcal{C}_{\boldsymbol{\gamma}}$ can be made using ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$, therefore the (minimum) number of triads made using ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ is $s({\boldsymbol{\gamma}})$.
Proof of Theorem 3.2:
Proof. Suppose that a minimal model $\boldsymbol{\gamma}$ has a single one 00-tick triad $(i,j,k)$. It follows that all other triads in $\mathcal{C}_{\boldsymbol{\gamma}}$ have at least one tick and that $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ij}=\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ik}=\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},jk}=0$. Now by (40) for any pair $(s,t)$
$\nu_{{\rm cyclic},st}=\sum_{(u,v,w)\in\mathcal{C}_{\boldsymbol{\gamma}}}{c}_{(u,v,w),st}{\gamma}_{uvw}.$ (41)
Therefore if $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ij}=0$ then there must be some other triad in which the comparison $(i,j)$ appears. Similarly for the pairs $(j,k)$ and $(i,k)$. Also notice that any two distinct triads can have at most one comparison in common and therefore the minimal model will have at least three more triads with two indices in common with $(i,j,k)$ and having at least $1$–tick. Assume that there are exactly three such triads. These must be $\boldsymbol{c}_{(i,j,s)},\boldsymbol{c}_{(j,k,t)},\boldsymbol{c}_{(i,k,u)}$ for some $s,t$ and $u$. By (41), $\nu_{{\rm cyclic},ij}={c}_{(i,j,k),ij}{\gamma}_{ijk}+{c}_{(i,j,s),ij}{\gamma}_{ijs}=0$ implying that ${\gamma}_{ijk}=-{\gamma}_{ijs}$. A similarly relation holds for $\nu_{{\rm cyclic},jk}=0$ and $\nu_{{\rm cyclic},ik}=0$ and therefore ${\gamma}_{ijk}=-{\gamma}_{jkt}$ and ${\gamma}_{ijk}={\gamma}_{iku}$. Combining these equalities we find that ${\gamma}_{ijk}=-{\gamma}_{ijs}=-{\gamma}_{jkt}={\gamma}_{iku}$. It is now clear that the elements of the vector $\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s)}-\boldsymbol{c}_{(j,k,t)}+\boldsymbol{c}_{(i,k,u)}$ which correspond to comparisons among the items $\{i,j,k\}$ are 00. This implies by (41) that
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}+\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(-ijk)},$
where $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}={\gamma}_{ijk}(\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s)}-\boldsymbol{c}_{(j,k,t)}+\boldsymbol{c}_{(i,k,u)})$ is the linear combination of the triads involving comparisons of the items in $\{i,j,k\}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(-ijk)}=\operatorname{\boldsymbol{\nu}}_{\rm cyclic}-\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}$.
Since all other triads in the minimal model have at least one tick we assume without any loss of generality that $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},js},\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},kt}\neq 0$. Using (18) we find
$\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s)}=\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s)}+\boldsymbol{c}_{(i,k,t)}-\boldsymbol{c}_{(i,k,t)}=\boldsymbol{c}_{(j,k,s)}-\boldsymbol{c}_{(i,k,t)}.$
Thus by using the above relation, we can rewrite $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}$ as
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}={\gamma}_{ijk}(\boldsymbol{c}_{(j,k,s)}-\boldsymbol{c}_{(i,k,t)}-\boldsymbol{c}_{(j,k,t)}+\boldsymbol{c}_{(i,k,u)}).$
Hence we obtain another minimal model, say $\boldsymbol{\gamma}^{\prime}$, in which there are no triads with 00 ticks. If there are more than three triads with two indices in common with $\{i,j,k\}$ then we can write $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}=\alpha_{1}(\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s_{1})}-\boldsymbol{c}_{(j,k,t_{1})}+\boldsymbol{c}_{(i,k,u_{1})})+\ldots+\alpha_{l}(\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s_{l})}-\boldsymbol{c}_{(j,k,t_{l})}+\boldsymbol{c}_{(i,k,u_{l})})$. Applying the preceding arguments to each term appearing in $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}$ separately we arrive at the same conclusion, i.e., there exist an alternative minimal model with no 00–tick triads. Finally, if there are $l>1$ 00–tick triads in a minimal model then we repeat the procedure outline above for each one of them. This completes the proof of the first part of Theorem 3.2.
Suppose that any two triads in a minimal model $\boldsymbol{\gamma}$ share at most one index. The proof of Theorem 3.1 shows that if $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ then $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ij}\neq 0$. This immediately implies that all triads in $\boldsymbol{\gamma}$ have $3$–ticks. Next it is shown, by example, that a triad $(i,j,k)$ may have $3$-ticks yet not be a member of any minimal model. Let
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\lambda_{1}\boldsymbol{c}_{(i,j,s)}+\lambda_{2}\boldsymbol{c}_{(j,k,t)}+\lambda_{3}\boldsymbol{c}_{(i,k,u)},$
where $|\lambda_{l}|$’s are distinct. It is not hard to see that the minimal model is unique. Also note that the triads $(i,j,s),~{}(j,k,t)$ and $(i,k,u)$ all have $3$ ticks. However so does the triad $(i,j,k)$. This completes the proof of the second part of the theorem.
Finally consider $\alpha\in\mathbb{R}\setminus\{0\}$ and set
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,2,3)}-\alpha\boldsymbol{c}_{(1,2,4)}-(1-\alpha)\boldsymbol{c}_{(1,2,5)}.$
It is easy to verify that the minimal model for $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ above is unique. It is straightforward to verify that all triads in the minimal model have two ticks. This completes the proof.
Proof of Theorem 3.3:
Proof. If $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}({\rm Tri}(\mathcal{G}))$ then there exists a $\boldsymbol{\gamma}\in\mathbb{R}^{\binom{K}{3}}$ such that $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{C}}\boldsymbol{\gamma}$ where $\boldsymbol{\gamma}$ has non-zero elements corresponding to triads present in ${\rm Tri}(\mathcal{G})$. Therefore retaining only the columns of $\operatorname{\boldsymbol{C}}$ corresponding to triads in ${\rm Tri}(\mathcal{G})$ we obtain $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{C}}_{\mathcal{G}}\boldsymbol{\gamma}_{\mathcal{G}}$, where $\boldsymbol{\gamma}_{\mathcal{G}}\in\mathbb{R}^{{\rm rank}({\rm span}({\rm Tri}(\mathcal{G})))}$. Next by Theorem 3.1 if the triads in ${\rm Tri}(\mathcal{G})$ share at most one index then the minimal model is unique and the conclusion follows immediately.
Proof of Theorem 3.4:
Proof. Assume there are $K$ items. Pick a triplet $(1,2,3)$, say. Recall, cf. Example 3.2, that when $K=3$ the relevant means satisfy
$\begin{pmatrix}\nu_{12}\\ \nu_{13}\\ \nu_{23}\end{pmatrix}=\begin{pmatrix}\mu_{1}-\mu_{2}+\gamma\\ \mu_{1}-\mu_{3}-\gamma\\ \mu_{2}-\mu_{3}+\gamma\end{pmatrix},$
for some $\gamma$. Suppose that $\operatorname{\boldsymbol{\nu}}$ is strongly transitive and without loss of generality assume that $1\succeq 2\succeq 3$. If so
$\nu_{12},\nu_{23}\geq 0\text{ and }\nu_{13}\geq\max\{\nu_{12},\nu_{23}\}.$
If $\gamma<0$, then $\nu_{12}\geq 0$ implies that $\mu_{1}-\mu_{2}>-\gamma>0$ so $\mu_{1}>\mu_{2}$. Similarly $\nu_{23}\geq 0$ implies $\mu_{2}>\mu_{3}$. Together these imply that $\mu_{1}\geq\mu_{2}\geq\mu_{3}$. Next if $\gamma>0$ then strong transitivity implies that $\mu_{1}-\mu_{3}-\gamma\geq\mu_{1}-\mu_{2}+\gamma$, so $\mu_{2}-\mu_{3}>2\gamma>0$. Similarly $\mu_{1}-\mu_{3}-\gamma\geq\mu_{2}-\mu_{3}+\gamma$ implies $\mu_{1}-\mu_{2}>2\gamma>0$. Combining these two relations we find that $\mu_{1}\geq\mu_{2}\geq\mu_{3}$ as required. Finally if $\gamma=0$ then we have $\mu_{1}-\mu_{2}\geq 0$ and $\mu_{2}-\mu_{3}\geq 0$ which immediately imply that $\mu_{1}\geq\mu_{2}\geq\mu_{3}$. Since this argument holds for any triplet we conclude that strong stochastic transitivity is equivalent to the ordering of the merits.
Next it will be shown that if $\operatorname{\boldsymbol{\nu}}$ is weakly transitive the merits may not be ordered. Under weak transitivity
$\nu_{12},\nu_{23}\geq 0\text{ implies that }\nu_{13}\geq 0.$
Set $\gamma>0$ and $\mu_{2}=\mu_{1}+\gamma/2$. If so $\nu_{12},\nu_{23},\nu_{13}\geq 0$ holds but $\mu_{1}<\mu_{2}$. We conclude that under weak transitivity the relations $1\succeq 2\succeq 3$ does not imply that $\mu_{1}\geq\mu_{2}\geq\mu_{3}$. This concludes the proof.
Proof of Proposition 3.1:
Proof. Without any loss of generality assume $\mu_{1}^{**}>\ldots>\mu_{K}^{**}$. Consequently: (i) $\nu_{1j}>0$ for $2\leq j\leq K$, so $1\succ j$ where $2\leq j\leq K$; (ii) for any $i>1$, $\nu_{ij}>0$ for $j>i$ and $\nu_{ij}<0$ for $j<i$. Therefore for any $i$ items $\{1,\ldots,i-1\}$ are preferred over $i$ and $i$ is preferred over items $\{i+1,\ldots,K\}$. Implying that the preference profile is weakly transitive with $1\succ\ldots\succ K$.
Proof of Theorem 4.1:
Proof. The proof follows the proof of Theorem 4.1 in Singh et al. (2024). The normal equation are given by
$\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{C}}\\ \operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{C}}\end{pmatrix}\begin{pmatrix}\operatorname{\boldsymbol{\mu}}\\ \boldsymbol{\gamma}\end{pmatrix}=\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}^{\top}\end{pmatrix}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{Y}}.$
Thus the minimum norm solution to the above system is given by
$\begin{pmatrix}\widehat{\operatorname{\boldsymbol{\mu}}}\\ \widehat{\boldsymbol{\gamma}}\end{pmatrix}=\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{C}}\\ \operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{C}}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}^{\top}\end{pmatrix}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{Y}}.$
Explicit expressions for $\widehat{\operatorname{\boldsymbol{\mu}}}$ and $\widehat{\boldsymbol{\gamma}}$ can be obtained by noting that
$\operatorname{\boldsymbol{Y}}=~{}\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}+\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}\boldsymbol{\gamma}+\boldsymbol{\epsilon}=~{}{\operatorname{\boldsymbol{L}}}{\operatorname{\boldsymbol{B}}}\boldsymbol{c}_{1}+\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}\boldsymbol{c}_{2}+\boldsymbol{\epsilon},$
where $\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}=(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}$ and
$\boldsymbol{c}_{2}$ $~{}=\boldsymbol{\gamma},$
$\boldsymbol{c}_{1}$ $~{}=\operatorname{\boldsymbol{\mu}}-(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}\boldsymbol{\gamma}.$
Note that $\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}$ is projection of $\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}$ onto the orthogonal complement of $\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}$, and consequently in terms of $(\boldsymbol{c}_{1},\boldsymbol{c}_{2})$ the normal equations are
$\begin{pmatrix}\operatorname{\boldsymbol{N}}&\boldsymbol{0}\\ \boldsymbol{0}&\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}^{\top}\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}\end{pmatrix}\begin{pmatrix}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\end{pmatrix}=\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{Y}}\\ \widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}^{\top}\operatorname{\boldsymbol{Y}}\end{pmatrix}.$
Here $(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})$ is an $n\times n$ idempotent matrix. Therefore the solution are
$\widehat{\operatorname{\boldsymbol{\mu}}}=\operatorname{\boldsymbol{N}}^{+}(\operatorname{\boldsymbol{S}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}\widehat{\boldsymbol{\gamma}})$
and
$\widehat{\boldsymbol{\gamma}}=(\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{L}}^{\top}(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}})^{+}\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{L}}^{\top}(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{Y}}.$
Proof of Theorem 4.2:
Proof. The proof mimics the proof of Theorem 4.2 in Singh et al. (2024) and is therefore omitted.
Proof of Theorem 4.3:
Proof. A bit of algebra shows that
$(\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}-\widehat{\operatorname{\boldsymbol{\nu}}}_{s})=(\boldsymbol{I}-(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}(\operatorname{\boldsymbol{D}}/n)\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}(\operatorname{\boldsymbol{D}}/n)\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}(\operatorname{\boldsymbol{D}}/n)\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}(\operatorname{\boldsymbol{D}}/n)\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\end{pmatrix}(\operatorname{\boldsymbol{D}}/n))\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}.$
By the CLT we have $\operatorname{\boldsymbol{D}}^{1/2}(\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}-{\operatorname{\boldsymbol{\nu}}}_{s})\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\sigma^{2}\boldsymbol{H}_{s})$, where $\boldsymbol{H}_{s}$ is a $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ diagonal matrix whose diagonal elements are $1$ if $(i,j)\in\mathcal{E}_{n}$ and 00 otherwise. It follows by Slutzky’s theorem that $\boldsymbol{U}_{n,\mathcal{S}}=\operatorname{\boldsymbol{D}}^{1/2}(\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}-\widehat{\operatorname{\boldsymbol{\nu}}}_{s})\Rightarrow\boldsymbol{U}_{\mathcal{S}}$ where $\boldsymbol{U}_{\mathcal{S}}$ follows a $\mathcal{N}_{|\operatorname{\boldsymbol{\nu}}|}(0,\boldsymbol{\Psi}_{\mathcal{S}})$ and ${\boldsymbol{\Psi}}_{\mathcal{S}}=\sigma^{2}\,\boldsymbol{\Xi}^{1/2}\boldsymbol{M}\,\boldsymbol{\Xi}^{+}\boldsymbol{M}^{\top}\boldsymbol{\Xi}^{1/2}$, where
$\boldsymbol{M}=\boldsymbol{I}-(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\end{pmatrix}\boldsymbol{\Xi}.$
Therefore $R_{n,\mathcal{S}}=\boldsymbol{U}_{n,\mathcal{S}}^{\top}\boldsymbol{U}_{n,\mathcal{S}}\Rightarrow\boldsymbol{U}_{\mathcal{S}}^{\top}\boldsymbol{U}_{\mathcal{S}}$. Next proceeding similarly as in the proof of Theorem 2.1 we get
$\boldsymbol{U}_{\mathcal{S}}^{\top}\boldsymbol{U}_{\mathcal{S}}=\sum_{i=1}^{{\rm rank}({\boldsymbol{\Psi}}_{\mathcal{S}})}{\lambda}_{i}Z_{i}^{2},$
where $Z_{1},\ldots,Z_{t}$ are independent $\mathcal{N}(0,1)$ RVs and ${\lambda}_{1},\ldots,{\lambda}_{t}$ are the non–zero eigenvalues of the matrix ${\boldsymbol{\Psi}}_{\mathcal{S}}$ which by Remark A.1 satisfies $t={\rm rank}({\boldsymbol{\Psi}}_{\mathcal{S}})=\binom{K}{2}-(K+r-1)$.
Proof of Proposition 4.1:
Proof. Under $H_{0}$, using Theorem 2.2 asymptotic normality of $T_{n}$ is immediate. Next note that $\nu_{ij}+\nu_{jk}+\nu_{ki}=(\operatorname{\boldsymbol{C}}_{ij}+\operatorname{\boldsymbol{C}}_{jk}-\operatorname{\boldsymbol{C}}_{ik})\boldsymbol{\gamma}$, so under $H_{1}$ $T_{n}$ converges in probability to $\delta$ as $n_{ij},n_{jk},n_{ik}\to\infty$.. Consequently using Theorem 2.2 again under $H_{1}$ we get $T_{n}\Rightarrow\mathcal{N}(\delta,\sigma^{2})$ as $n_{ij},n_{jk},n_{ik}\to\infty$.
Proof of Theorem 4.4:
We start with an auxiliary Lemma.
Lemma A.1
Let $\mathcal{S}=\rm{span}(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})$ for some matrix $\operatorname{\boldsymbol{C}}_{s}$ satisfying $\rm{span}(\operatorname{\boldsymbol{C}}_{s})\subset\rm{span}(\operatorname{\boldsymbol{C}})$. Consider testing $H_{0}:\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$ versus $H_{1}:\operatorname{\boldsymbol{\nu}}\notin\mathcal{S}$ using the statistic $R_{n,\mathcal{S}}$ defined in Theorem 4.3. Assuming $\operatorname{\boldsymbol{\nu}}\notin\mathcal{S}$, $n_{ij}\rightarrow\infty$ for all $1\leq i<j\leq K$ while $n_{ij}=O(n)$, we then have
$\mathbb{P}(R_{n,\mathcal{S}}<c_{\alpha})\rightarrow 0$ (42)
where $c_{\alpha}=\mathbb{P}_{H_{0}}(R_{n,\mathcal{S}}\geq c_{\alpha})$ is the $\alpha$ level critical value for $R_{n,\mathcal{S}}$. If, in addition, the errors are subgaussian then for all large $n$ there exists constants $C_{1},C_{2}>0$ such that
$\mathbb{P}(R_{n,\mathcal{S}}<c_{\alpha})\leq C_{1}\exp(-nC_{2}).$ (43)
Remark A.3
Equation (42) shows that under the alternative the power of tests for lack of fit approaches unity as the sample size increases to $\infty$. In fact, this assertion has already been established see the text following the statement of Theorem 4.3. Therefore we only prove (43).
Proof. Observe that
$R_{n,\mathcal{S}}=\sum_{1\leq i<j\leq K}{U}_{n,\mathcal{S},ij}^{2}=\sum_{1\leq i<j\leq K}n_{ij}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})^{2},$ (44)
where $\overline{S}_{{\rm ALL},ij}$ and $\widehat{\nu}_{s,ij}$ are $ij^{th}$ elements of $\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}$ and $\widehat{\operatorname{\boldsymbol{\nu}}}_{s}$, respectively. Notice that both $\overline{S}_{{\rm ALL},ij}$ and $\widehat{\nu}_{s,ij}$ are linear functions of the errors. Therefore $(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})$ is also a subgaussian random variable with mean $\delta_{ij}$ and parameter $\tau_{ij}^{2}\leq\infty$. Note that under the null, i.e., when $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$, we have $\delta_{ij}=\mathbb{E}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})=0$ for all pairs $(i,j)$. However, under the alternative $\delta_{ij}\neq 0$ for some pairs $(i,j)$.
Continuing,
$\mathbb{P}(|\sqrt{n_{ij}}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})-\sqrt{n_{ij}}\delta_{ij}|\geq\epsilon)\leq 2\exp(-\epsilon^{2}n_{ij}/2\tau_{ij}^{2})$
In other words
$\mathbb{P}(\sqrt{n_{ij}}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})\notin(\sqrt{n_{ij}}\delta_{ij}-\epsilon,\sqrt{n_{ij}}\delta_{ij}+\epsilon))\leq 2\exp(-\epsilon^{2}n_{ij}/2\tau_{ij}^{2}),$
so, for all large $n$
$\mathbb{P}({n_{ij}}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})^{2}\notin({n_{ij}}\delta_{ij}^{2}-\epsilon_{ij},{n_{ij}}\delta_{ij}^{2}+\epsilon_{ij}))\leq 2\exp(-\epsilon^{2}n_{ij}/2\tau_{ij}^{2})$
for some $\epsilon_{ij}>0$ where $\epsilon_{ij}\rightarrow 0$ when $\epsilon\rightarrow 0$. Next observe that
$\cup_{1\leq i<j\leq K}\{{U}_{n,\mathcal{S},ij}^{2}\in(n_{ij}\delta_{ij}^{2}-\epsilon_{ij},n_{ij}\delta_{ij}^{2}+\epsilon_{ij})\}\subset\{R_{n,\mathcal{S}}\in(\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}-\widetilde{\epsilon},\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}+\widetilde{\epsilon})\}$
where $\widetilde{\epsilon}=\sum_{1\leq i<j\leq K}\epsilon_{ij}$. It follows that for all large $n$
$~{}\mathbb{P}(R_{n,\mathcal{S}}\notin(\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}-\widetilde{\epsilon},\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}+\widetilde{\epsilon}))$
$\leq$ $~{}\mathbb{P}(\cup_{1\leq i<j\leq K}\{{U}_{n,\mathcal{S},ij}^{2}\notin(n_{ij}\delta_{ij}^{2}-\epsilon_{ij},n_{ij}\delta_{ij}^{2}+\epsilon_{ij})\})\leq\sum_{1\leq i<j\leq K}2\exp(-\epsilon^{2}n_{ij}/2\tau_{ij}^{2})$
$\leq$ $~{}C_{1}\exp(-n\epsilon^{2}C_{2}),$
for $C_{1}=2\binom{K}{2}$ and $C_{2}=\min\{n_{ij}/(2n\tau_{ij}^{2}):1\leq i<j\leq K\}$. Therefore
$~{}\mathbb{P}(R_{n,\mathcal{S}}\in(\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}-\widetilde{\epsilon},\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}+\widetilde{\epsilon}))\geq 1-C_{1}\exp(-n\epsilon^{2}C_{2}).$
However for any $\epsilon$ and all large $n$, $\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}-\widetilde{\epsilon}>c_{\alpha}$ so for all large $n$
$\mathbb{P}(R_{n,\mathcal{S}}\geq c_{\alpha})\geq 1-C_{1}\exp(-nC_{2}).$
Taking complements completes the proof.
Remark A.4
Lemma A.1 shows that under subgaussian errors the probability of not rejecting the null when applying ${\rm FTBS}$ it is false is exponentially small.
We now continue with the proof of Theorem 4.4.
Proof. FTBS is applied by $(i)$ testing $\binom{K}{2}$ hypotheses $H_{0}:\nu_{{\rm cyclic},ij}=0$, ${1\leq i<j\leq K}$ and generating an estimated tick–table $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}}$; and $(ii)$ then sequentially testing for lack of fit using Theorems 2.1 and 4.3. At most $3$ such tests are applied. The set of selected triads obtained after applying $(i)$ and $(ii)$ above is denoted by $\widehat{\mathcal{S}}$.
By Theorem 3.2 there exists a minimal model $\mathcal{S}$ in which all triads have at least one tick in $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cycllic}}$. Further note that $\mathcal{S}\subseteq\mathcal{S}_{b}$ for some $b\in\{1,\ldots,4\}$. Now, if there is no cyclicality in $\mathcal{S}$ then $\mathcal{S}\subseteq\mathcal{S}_{1}$; if all triads in $\mathcal{S}$ have three ticks then $\mathcal{S}\subseteq\mathcal{S}_{2}$; if some triads in $\mathcal{S}$ have two ticks then $\mathcal{S}\subseteq\mathcal{S}_{3}$; and if some triads in $\mathcal{S}$ have one tick then $\mathcal{S}\subseteq\mathcal{S}_{4}$. Thus if $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cycllic}}$ was known then applying Theorems 2.1 and 4.3 would guarantee that FTBS would select $\widehat{\mathcal{S}}\supset\mathcal{S}$.
Now, the event $\{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\notin{\rm span}(\widehat{\mathcal{S}})\}$ will occur if either $\{\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cycllic}}\not\subset\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cycllic}^{*}}\}$ or if for some $a\in\{1,\ldots,3\}$ smaller than $b$ the event $\{R_{n,\mathcal{S}_{a}}<c_{\alpha,a}\}$ occurs. Also note that if $\{\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cycllic}}\subset\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cycllic}^{*}}\}$ then $\mathcal{S}_{a}\subset\mathcal{S}_{a}^{*}$, moreover $\mathcal{S}_{b}\subset\mathcal{S}_{a}^{*}$ may hold for some $b>a$. It follows that:
$\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\notin{\rm span}(\widehat{\mathcal{S}}))=$ ($~{}\mathbb{P}(\{{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}}\}\cup_{\{a:\,\mathcal{S}_{b}\not\subset\mathcal{S}_{a}^{*}\}}(\{R_{n,\mathcal{S}_{a}}<c_{\alpha,a}\}\cap\{{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}}\})))$
$\leq$ $~{}\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})+\sum_{\{a:\,\mathcal{S}_{b}\not\subset\mathcal{S}_{a}^{*}\}}\mathbb{P}(\{R_{n,\mathcal{S}_{a}^{*}}<c_{\alpha,a}\}).$ (45)
By Lemma A.1 if $\mathcal{S}_{b}\not\subset\mathcal{S}_{a}^{*}$ then with high probability we will reject the hypothesis that $\mathcal{S}\subset\mathcal{S}_{a}^{*}$. Therefore the sum in (45) is over set $\{a:\mathcal{S}_{b}\not\subset\mathcal{S}_{a}^{*}\}$.
Next we bound the individual probabilities in (45). First note that if $\nu_{{\rm cyclic},ij}\neq 0$ and the procedure described in Remark 4.1 is followed then $\mathbb{P}(\phi_{ij}(\widehat{\operatorname{\boldsymbol{\nu}}}_{{\rm cyclic}}=1)\rightarrow 1$, i.e., the null $H_{0}:\nu_{{\rm cyclic},ij}=0$ is rejected with probability tending to one. Since $K$ is finite all elements of $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ which are different than 00 will be identified with probability tending to one. Furthermore approximately $\alpha(\binom{K}{2}-\sum_{1\leq i<j\leq K}\mathbb{I}({\nu_{{\rm cyclic},ij}}\neq 0))$ other elements of $\operatorname{\boldsymbol{\nu}}$ will be falsely flagged as non–zero. We have just proved that
$\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})\to 0.$ (46)
Consequently $\mathcal{S}_{a}\subset\mathcal{S}_{a}^{*}$ for $a\in\{1,\ldots,4\}$ with probability tending to $1$ as $n\to\infty$. Next using Lemma A.1 for $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}_{b}$ and $b>a$, we have
($\mathbb{P}(\{R_{n,\mathcal{S}_{a}^{*}}<c_{\alpha,a}\})\to 0.$ (47)
Therefore (46) and (47) imply that
$\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\notin{\rm span}(\widehat{\mathcal{S}}))\to 0,$
establishing the first stated claim.
[…]
$\mathbb{P}(|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\geq\epsilon)\leq 2e^{-\epsilon^{2}/({\tau}_{ij}^{2}/n)}=2e^{-\epsilon^{2}n/{\tau}_{ij}^{2}}.$ (48)
Next using union bound and (48) we have
$\mathbb{P}(\cup_{1\leq i<j\leq K}\{|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\geq\epsilon\})$ $~{}\leq\sum_{1\leq i<j\leq K}\mathbb{P}(\{|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\geq\epsilon\})$
$~{}\leq\sum_{1\leq i<j\leq K}2e^{-\epsilon^{2}n/{\tau}_{ij}^{2}}.$
Let ${\tau}_{{\rm cyclic}}^{2}=\min\{{\tau}_{ij}^{2}:1\leq i<j\leq K\}$. Then
$\mathbb{P}(\cup_{1\leq i<j\leq K}\{|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\leq\epsilon\})\geq 1-\sum_{1\leq i<j\leq K}2e^{-\epsilon^{2}n/{\tau}_{ij}^{2}}\geq 1-\binom{K}{2}2e^{-\epsilon^{2}n/{\tau}_{{\rm cyclic}}^{2}}.$
Since $K$ is finite, there exist $\bar{C}_{1},\bar{C}_{2}>0$ such that
$\mathbb{P}(\cup_{1\leq i<j\leq K}\{|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\geq\epsilon\})\leq\bar{C}_{1}e^{(-n\bar{C}_{2})}.$
Therefore not rejecting $H_{0}:\nu_{{\rm cyclic},ij}=0$ when $H_{0}$ is false has probability
$\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})\leq\bar{C}_{1}e^{(-n\bar{C}_{2})}.$ (49)
Next using Lemma A.1 for $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}_{b}$ and $b>a$, we have
($\mathbb{P}(\{R_{n,\mathcal{S}_{a}^{*}}<c_{\alpha,a}\})\leq C_{a1}e^{(-nC_{a2})}.$ (50)
Thus (45), (49) and (50) imply that
$\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\notin{\rm span}(\widehat{\mathcal{S}}))\leq$ $~{}\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})+\sum_{1\leq a<b}\mathbb{P}(\{R_{n,\mathcal{S}_{a}^{*}}<c_{\alpha,a}\})$
$\leq$ $~{}\bar{C}_{1}e^{(-n\bar{C}_{2})}+\sum_{1\leq a<b}C_{a1}e^{(-nC_{a2})}\leq C_{1}e^{(-nC_{2})}$
for some finite $C_{1},C_{2}>0$. Therefore there exists $C_{1},C_{2}>0$ such that
$\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))\geq 1-C_{1}e^{(-nC_{2})}.$
Thus establishing the second claim.
By Theorem 3.2 if triads in the minimal model share at most index then the minimal model is unique, moreover only triads which are part of the minimal model will have three ticks. Condition (36) guarantees that $H_{0}:{\nu}_{{\rm cyclic},ij}=0$ is rejected with probability tending to one if and only if ${\nu}_{{\rm cyclic},ij}\neq 0$. Thus $\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}={\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})\to 1$. Further note that under the stated condition $\mathcal{S}=\mathcal{S}_{2}$ and $\mathcal{S}_{2}=\mathcal{S}_{2}^{*}$ with probability tending to $1$. Since the hypotheses in Step 2 of FTBS are tested at level $\alpha_{n}\rightarrow 0$ the algorithm FTBS will terminate at $a=2$ with probability tending to one. In other words $\mathbb{P}(\mathcal{S}=\widehat{\mathcal{S}})\to 1$, concluding the proof.
Appendix B Supplementary Material
B.1 Tick Tables for Simulation Scenarios
Scenario I
123 124 125 126 134 135 136 145 146 156 234 235 236 245 246 256 345 346 356 456
12
13
14
15
16
23
24
25
26
34
35
36
45
46
56
Scenario II
123 124 125 126 134 135 136 145 146 156 234 235 236 245 246 256 345 346 356 456
12
13
14
15
16
23
24
25
26
34
35
36
45
46
56
Scenario III
123 124 125 126 134 135 136 145 146 156 234 235 236 245 246 256 345 346 356 456
12
13
14
15
16
23
24
25
26
34
35
36
46
56
B.2 EPL data
Figure 2: Triads and their p-values
Table 6: First $100$ most significant triads with their p-values
Serial Triad p-value Serial Triad p-value Serial Triad p-value Serial Triad p-value
1 (2,5,19) 0.0001056949 26 (11,13,17) 0.0033286082 51 (4,12,19) 0.0070912725 76 (1,6,8) 0.014200921
2 (6,11,12) 0.0002864972 27 (3,7,9) 0.0038915299 52 (5,6,12) 0.0070912725 77 (1,7,14) 0.014200921
3 (6,12,14) 0.0002864972 28 (10,14,20) 0.0038915299 53 (11,16,19) 0.0070912725 78 (3,11,19) 0.014200921
4 (10,11,13) 0.0004193156 29 (1,14,16) 0.0038915299 54 (5,8,19) 0.0081880626 79 (8,14,17) 0.014200921
5 (2,5,13) 0.0004193156 30 (8,13,19) 0.0038915299 55 (2,4,5) 0.0081880626 80 (2,3,11) 0.0161972879
6 (1,2,5) 0.0007283604 31 (2,8,13) 0.0045382968 56 (3,4,19) 0.0081880626 81 (2,3,13) 0.0161972879
7 (6,14,15) 0.0007283604 32 (4,10,13) 0.0045382968 57 (4,6,12) 0.0081880626 82 (3,8,16) 0.0161972879
8 (2,5,6) 0.0008710999 33 (5,6,15) 0.0045382968 58 (14,16,17) 0.0081880626 83 (3,14,16) 0.0161972879
9 (7,11,13) 0.0010391678 34 (10,13,14) 0.0045382968 59 (3,6,12) 0.0094311784 84 (4,14,15) 0.0161972879
10 (5,17,19) 0.0012365207 35 (11,13,15) 0.0045382968 60 (6,7,11) 0.0094311784 85 (6,12,17) 0.0161972879
11 (4,5,15) 0.0012365207 36 (1,7,9) 0.0052793771 61 (6,14,17) 0.0094311784 86 (10,16,20) 0.0161972879
12 (14,15,20) 0.0012365207 37 (2,4,13) 0.0052793771 62 (3,5,19) 0.0108363111 87 (12,14,20) 0.0161972879
13 (1,6,12) 0.0014676313 38 (2,4,19) 0.0052793771 63 (6,7,10) 0.0108363111 88 (13,14,15) 0.0161972879
14 (4,16,19) 0.0014676313 39 (4,11,15) 0.0052793771 64 (9,14,17) 0.0108363111 89 (3,15,16) 0.0184294035
15 (5,16,19) 0.0014676313 40 (7,9,11) 0.0052793771 65 (12,14,16) 0.0108363111 90 (4,10,19) 0.0184294035
16 (11,13,16) 0.0014676313 41 (11,12,13) 0.0052793771 66 (14,15,16) 0.0108363111 91 (5,6,7) 0.0184294035
17 (1,11,13) 0.0020518965 42 (1,2,12) 0.0052793771 67 (8,10,13) 0.0108363111 92 (5,14,16) 0.0184294035
18 (4,9,19) 0.0020518965 43 (2,5,9) 0.0052793771 68 (11,13,14) 0.0108363111 93 (6,10,12) 0.0184294035
19 (7,9,19) 0.0024170324 44 (1,2,4) 0.0061262151 69 (1,2,16) 0.0124202582 94 (8,9,13) 0.0184294035
20 (1,2,8) 0.0028399943 45 (3,11,13) 0.0061262151 70 (1,13,14) 0.0124202582 95 (8,11,16) 0.0184294035
21 (1,8,14) 0.0028399943 46 (13,14,17) 0.0061262151 71 (3,6,11) 0.0124202582 96 (14,17,20) 0.0184294035
22 (4,5,17) 0.0028399943 47 (2,11,13) 0.0061262151 72 (1,8,9) 0.0124202582 97 (11,14,17) 0.0184294035
23 (1,2,3) 0.0033286082 48 (4,9,15) 0.0061262151 73 (3,14,15) 0.0124202582 98 (11,14,19) 0.0184294035
24 (1,2,17) 0.0033286082 49 (6,11,15) 0.0070912725 74 (4,11,14) 0.0124202582 99 (1,8,11) 0.0209183218
25 (6,10,13) 0.0033286082 50 (2,5,11) 0.0070912725 75 (6,12,16) 0.0124202582 100 (1,8,19) 0.0209183218
Table 7: Components of $\widehat{\operatorname{\boldsymbol{\nu}}}_{s}$
Serial $(i,j)$ $\widehat{\nu}_{s,ij}$ Serial $(i,j)$ $\widehat{\nu}_{s,ij}$ Serial $(i,j)$ $\widehat{\nu}_{s,ij}$ Serial $(i,j)$ $\widehat{\nu}_{s,ij}$ Serial $(i,j)$ $\widehat{\nu}_{s,ij}$
1 (1,2) 7.90E-01 39 (3,5) -1.22E+00 77 (5,12) 4.50E-02 115 (8,11) -1.83E-01 153 (11,19) -2.28E-01
2 (1,3) 1.37E+00 40 (3,6) -5.65E-01 78 (5,13) -5.75E-01 116 (8,12) -1.05E+00 154 (11,20) 2.60E-01
3 (1,4) 5.68E-01 41 (3,7) -4.13E-01 79 (5,14) 1.55E-01 117 (8,13) -1.67E+00 155 (12,13) -6.20E-01
4 (1,5) 1.48E-01 42 (3,8) -1.30E-01 80 (5,15) -2.20E-01 118 (8,14) -9.35E-01 156 (12,14) 1.10E-01
5 (1,6) 8.03E-01 43 (3,9) -2.03E-01 81 (5,16) 1.20E+00 119 (8,15) -1.31E+00 157 (12,15) -2.65E-01
6 (1,7) 9.55E-01 44 (3,10) -1.45E-01 82 (5,17) 1.17E+00 120 (8,16) 1.13E-01 158 (12,16) 1.16E+00
7 (1,8) 1.24E+00 45 (3,11) -3.13E-01 83 (5,18) 3.90E-01 121 (8,17) 7.75E-02 159 (12,17) 1.12E+00
8 (1,9) 1.17E+00 46 (3,12) -1.18E+00 84 (5,19) 1.86E+00 122 (8,18) -7.00E-01 160 (12,18) 3.45E-01
9 (1,10) 1.22E+00 47 (3,13) -1.80E+00 85 (5,20) 1.17E+00 123 (8,19) -4.10E-01 161 (12,19) 6.35E-01
10 (1,11) 1.06E+00 48 (3,14) -1.07E+00 86 (6,7) 1.53E-01 124 (8,20) 7.75E-02 162 (12,20) 1.12E+00
11 (1,12) 1.93E-01 49 (3,15) -1.44E+00 87 (6,8) 4.35E-01 125 (9,10) 5.75E-02 163 (13,14) 7.30E-01
12 (1,13) -4.28E-01 50 (3,16) -1.75E-02 88 (6,9) 3.63E-01 126 (9,11) -1.10E-01 164 (13,15) 3.55E-01
13 (1,14) 3.03E-01 51 (3,17) -5.25E-02 89 (6,10) 4.20E-01 127 (9,12) -9.73E-01 165 (13,16) 1.78E+00
14 (1,15) -7.25E-02 52 (3,18) -8.30E-01 90 (6,11) 2.53E-01 128 (9,13) -1.59E+00 166 (13,17) 1.74E+00
15 (1,16) 1.35E+00 53 (3,19) -5.40E-01 91 (6,12) -6.10E-01 129 (9,14) -8.63E-01 167 (13,18) 9.65E-01
16 (1,17) 1.32E+00 54 (3,20) -5.25E-02 92 (6,13) -1.23E+00 130 (9,15) -1.24E+00 168 (13,19) 1.26E+00
17 (1,18) 5.38E-01 55 (4,5) -4.20E-01 93 (6,14) -5.00E-01 131 (9,16) 1.85E-01 169 (13,20) 1.74E+00
18 (1,19) 8.28E-01 56 (4,6) 2.35E-01 94 (6,15) -8.75E-01 132 (9,17) 1.50E-01 170 (14,15) -3.75E-01
19 (1,20) 1.32E+00 57 (4,7) 3.88E-01 95 (6,16) 5.48E-01 133 (9,18) -6.28E-01 171 (14,16) 1.05E+00
20 (2,3) 5.78E-01 58 (4,8) 6.70E-01 96 (6,17) 5.13E-01 134 (9,19) -3.38E-01 172 (14,17) 1.01E+00
21 (2,4) -2.23E-01 59 (4,9) 5.98E-01 97 (6,18) -2.65E-01 135 (9,20) 1.50E-01 173 (14,18) 2.35E-01
22 (2,5) 5.41E-01 60 (4,10) 6.55E-01 98 (6,19) 2.50E-02 136 (10,11) -1.68E-01 174 (14,19) 5.25E-01
23 (2,6) 1.25E-02 61 (4,11) 4.88E-01 99 (6,20) 5.13E-01 137 (10,12) -1.03E+00 175 (14,20) 1.01E+00
24 (2,7) 1.65E-01 62 (4,12) -3.75E-01 100 (7,8) 2.83E-01 138 (10,13) -1.65E+00 176 (15,16) 1.42E+00
25 (2,8) 4.48E-01 63 (4,13) -9.95E-01 101 (7,9) 2.10E-01 139 (10,14) -9.20E-01 177 (15,17) 1.39E+00
26 (2,9) 3.75E-01 64 (4,14) -2.65E-01 102 (7,10) 2.68E-01 140 (10,15) -1.30E+00 178 (15,18) 6.10E-01
27 (2,10) 4.33E-01 65 (4,15) -6.40E-01 103 (7,11) 1.00E-01 141 (10,16) 1.28E-01 179 (15,19) 9.00E-01
28 (2,11) 2.65E-01 66 (4,16) 7.83E-01 104 (7,12) -7.63E-01 142 (10,17) 9.25E-02 180 (15,20) 1.39E+00
29 (2,12) -5.98E-01 67 (4,17) 7.48E-01 105 (7,13) -1.38E+00 143 (10,18) -6.85E-01 181 (16,17) -3.50E-02
30 (2,13) -1.22E+00 68 (4,18) -3.00E-02 106 (7,14) -6.53E-01 144 (10,19) -3.95E-01 182 (16,18) -8.13E-01
31 (2,14) -4.88E-01 69 (4,19) 2.60E-01 107 (7,15) -1.03E+00 145 (10,20) 9.25E-02 183 (16,19) -5.23E-01
32 (2,15) -8.63E-01 70 (4,20) 7.48E-01 108 (7,16) 3.95E-01 146 (11,12) -8.63E-01 184 (16,20) -3.50E-02
33 (2,16) 5.60E-01 71 (5,6) 6.55E-01 109 (7,17) 3.60E-01 147 (11,13) -1.48E+00 185 (17,18) -7.78E-01
34 (2,17) 5.25E-01 72 (5,7) 8.08E-01 110 (7,18) -4.18E-01 148 (11,14) -7.53E-01 186 (17,19) -4.88E-01
35 (2,18) -2.53E-01 73 (5,8) 1.09E+00 111 (7,19) -1.28E-01 149 (11,15) -1.13E+00 187 (17,20) 1.91E-15
36 (2,19) -1.15E+00 74 (5,9) 1.02E+00 112 (7,20) 3.60E-01 150 (11,16) 2.95E-01 188 (18,19) 2.90E-01
37 (2,20) 5.25E-01 75 (5,10) 1.08E+00 113 (8,9) -7.25E-02 151 (11,17) 2.60E-01 189 (18,20) 7.78E-01
38 (3,4) -8.00E-01 76 (5,11) 9.08E-01 114 (8,10) -1.50E-02 152 (11,18) -5.18E-01 190 (19,20) 4.88E-01
Table 8: Rank sets
Team $A_{i}$ $B_{i}$ $E_{i}$
1 12,13 2,3,4,5,6,7,8,9,10,11,14,15,16,17,18,19,20
2 1,6,7,11,12,13,14,15,19 3,4,5,8,10,16,17,18,20 9
3 1,2,4,5,6,7,10,12,13,14,15,16,18,19,20 8,9,11,17
4 1,2,5,11,12,13 3,6,7,8,9,10,14,15,16,17,18,19,20
5 1,2,13,15 3,4,6,7,8,9,10,11,12,14,16,17,18,19,20
6 1,4,5,10,11,13,14,15,16,19 2,3,7,8,9,12,17,18,20
7 1,4,5,6,8,9,12,13,14,15,17,18,20 2,3,10,11,16,19
8 1,2,3,4,5,6,9,11,12,13,14,15,18,19 7,10,16,17,20
9 1,3,4,5,6,11,12,13,14,15,18,19 7,8,10,16,17,20 2
10 1,2,4,5,7,8,9,11,12,13,14,15,18,19,20 3,6,16,17
11 1,3,5,7,12,14,15,16,17 2,4,6,8,9,10,13,19,20 18
12 5,6,13 1,2,3,4,7,8,9,10,11,14,15,16,17,18,19,20
13 11 1,2,3,4,5,6,7,8,9,10,12,14,15,16,17,18,19,20
14 1,4,5,12,13,15,17 2,3,6,7,8,9,10,11,16,18,20 19
15 1,4,12,13 2,3,5,6,7,8,9,10,11,14,16,17,18,19,20
16 1,2,4,5,7,8,9,10,12,13,14,15,17,18,19 3,6,11,20
17 1,2,3,4,5,6,8,9,10,12,13,15,18,19 7,11,14,16,20
18 1,2,4,5,6,12,13,14,15 1,2,4,5,6,12,13,14,15 11
19 1,4,5,7,11,12,13,15,18 2,3,6,8,9,10,16,17,20 14
20 1,2,4,5,6,8,9,11,12,13,14,15,16,17,18,19 3,7,10