### v2

# The analysis of paired comparison data in the presence of cyclicality and intransitivity

Abstract
A principled approach to cyclicality and intransitivity in paired comparison data is developed. The proposed methodology enables more precise estimation of the underlying preference profile and facilitates the identification of all cyclic patterns and potential intransitivities. Consequently, it improves upon existing methods for ranking and prediction, including enhanced performance in betting and wagering systems. Fundamental to our development is a detailed understanding and study of the parameter space that accommodates cyclicality and intransitivity. It is shown that identifying cyclicality and intransitivity reduces to a model selection problem, and a new method for model selection employing geometrical insights, unique to the problem at hand, is proposed. The large sample properties of the estimators and guarantees on the selected model are provided. Thus, it is shown that in large samples all cyclical relations and consequent intransitivities can be identified. The method is exemplified using simulations and analysis of an illustrative example.
Key–Words: Cyclicality, Goodness of fit, Intransitivity, Large sample properties, Least square ranking, Model selection.
1 Introduction
There are many situations in which a preference profile, or more restrictively a ranking, among a set of items is desired. Examples include the evaluation of political candidates, marketing, sports, information retrieval, and a variety of modern internet–related applications. There is an extensive literature on the theory, methodology, and applications of ranking methods spanning a variety of disciplines, including mathematics (Langville and Meyer, 2012), social choice theory (Sen, 1986; Slutzki and Volij, 2005), machine learning (Ailon et al., 2008; Fürnkranz and Hüllermeier, 2010), psychology (Regenwetter et al., 2011; Davis–Stober et al., 2016), and statistics (e.g., Marden, 1995; Alvo and Yu, 2014).
This paper addresses ranking from paired comparison data (PCD) (David, 1988). To fix ideas, suppose that there are $K$ items labeled $1,\ldots,K$. Let $Y_{ijk}$ denote the outcome of the $k^{th}$ comparison among items $i$ and $j$. The random variable (RV) $Y_{ijk}$ may be binary, ordinal, or cardinal. In this paper, we focus on cardinal, that is, continuous, paired comparison data (PCD); the applicability of the proposed methodology to binary PCD will also be discussed. Concretely, we shall assume that the observations $Y_{ijk}$ for $1\leq i\neq j\leq K$ and $k=1,\ldots,n_{ij}$ satisfy
$Y_{ijk}=\nu_{ij}+\epsilon_{ijk},$ (1)
where $\nu_{ij}=\mathbb{E}(Y_{ijk})$ and the errors $\epsilon_{ijk}$ are independent identically distributed RVs with zero mean and variance $\sigma^{2}$. In cardinal PCD $Y_{ijk}=-Y_{jik}$ and $\nu_{ij}=-\nu_{ji}$ (Jiang et al. 2011). Thus, the model (1) is indexed by a parameter $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}\equiv\mathbb{R}^{K(K-1)/2}$ where $\operatorname{\boldsymbol{\nu}}=(\nu_{12},\ldots,\nu_{1K},\nu_{23},\ldots,\nu_{2K},\ldots,\nu_{K-1,K})$ is a vector of means, arranged in lexicographic order. We refer to $\operatorname{\boldsymbol{\nu}}$ as the preference profile. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denote the graph whose vertices are the items $1,\ldots,K$ and the edges are the pairs $(i,j)$ for which $n_{ij}>1$. With each edge $(i,j)$ we associate a random sample $\mathcal{Y}_{ij}=(Y_{ij1},\ldots,Y_{ijn_{ij}})$ of $n_{ij}$ comparisons; the set of all samples is denoted by $\mathcal{Y}$. The pair $(\mathcal{G},\mathcal{Y})$ is called a pairwise comparison graph (PCG). Typically, it is assumed that
$\nu_{ij}=\mu_{i}-\mu_{j}\quad\mbox{for all}\quad 1\leq i\neq j\leq K,$ (2)
where $\mu_{1},\ldots,\mu_{K}$ are known as scores or merits. Note that when the restriction (2) holds
$\operatorname{\boldsymbol{\nu}}^{\top}=(\mu_{1}-\mu_{2},\ldots\mu_{1}-\mu_{K},\mu_{2}-\mu_{3},\ldots,\mu_{2}-\mu_{K},\ldots,\mu_{K-1}-\mu_{K})$ (3)
for some $\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}$. The collection of vectors satisfying (3) will be denoted by $\mathcal{L}$. It is well known that the parameterization (2) results in a non–identifiable model. Therefore, it is common to further assume that the merit vector satisfies a linear constraint such as $\mu_{1}+\cdots+\mu_{K}=0$. Consequently $\mathcal{L}$ can be further identified with $\mathbb{R}^{K-1}$. The least squares estimator (LSE) for model (1), under the assumption (2), has a long history and has been studied in a variety of fields (see, for example, Mosteller 1951; Kwiesielewicz 1996; Csató 2015). For a recent comprehensive account, see Singh et al. (2025). However, despite its great popularity, the assumption that $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$ is highly restrictive as it imposes a very strong form of transitivity as explained in the following.
We write $i\succeq j$ when $\nu_{ij}\geq 0$. The symbol $\succeq$ is commonly used to denote preference relations. The term preference is rooted in mathematical economics and psychology and is widely used in the ranking literature. A strict preference, that is, $i\succ j$ is associated with a strict inequality $\nu_{ij}>0$. The magnitude of $\nu_{ij}$ measures the degree of preference. Thus, as stated earlier, $\operatorname{\boldsymbol{\nu}}$ encodes a preference relations among the items $1,\ldots,K$. A preference relation between the items $i$, $j$ and $k$ is said to be transitive if $i\succeq j$ and $j\succeq k$ imply $i\succeq k$. If transitivity holds for all triplets, then a global ranking exists and $\operatorname{\boldsymbol{\nu}}$ is said to be transitive. It is emphasized that a global ranking exists only if preferences are transitive (Vriend, 1996). It is well known that the axioms of rationality imply transitivity; their violation would lead to the so–called money pump (Gustafsson 2022).
There are several types of stochastic transitivity relations, originally defined in the context of binary PCD. Recall that in binary PCD $Y_{ijk}\in\{0,1\}$ and $i\succeq j$ whenever $p_{ij}\geq 1/2$. Weak stochastic transitivity (WST) holds if, whenever $p_{ij}\geq 1/2$ and $p_{jk}\geq 1/2$, it follows that $p_{ik}\geq 1/2$ whereas strong stochastic transitivity (SST) requires that if $p_{ij}\geq 1/2$ and $p_{jk}\geq 1/2$, then $p_{ik}\geq\max\{p_{ij},p_{jk}\}$. An even stronger form of transitivity, which we refer to as linear stochastic transitivity (LST), requires that $p_{ij}=F(\mu_{i}-\mu_{j})$ for some symmetric distribution function $F$ and constants $\mu_{1},\ldots,\mu_{K}$. See Fishburn (1973) and Oliveira et al. (2018) for further details. The analogy to cardinal models is immediate: WST and SST can be defined by replacing $p_{ij}$ with $\nu_{ij}$ and $1/2$ with 0, while LST reduces to the relation (2). A concrete example illustrating WST, SST, and LST in the context of cardinal PCD appears in Example 2.2. It should also be noted that under LST, $i\succeq j$ whenever $\mu_{i}\geq\mu_{j}$; therefore, the vector of scores $\operatorname{\boldsymbol{\mu}}=(\mu_{1},\ldots,\mu_{K})$ uniquely determines the ranking among all items.
Despite the fact that transitivity underlies much of decision theory, numerous authors have advocated for relaxations of transitivity, e.g., Fishburn and LaValle (1988), Schmeidler (1989), Sen (1997) and Klimenko (2015). In fact, it is well known that not all preference relations are transitive. For example, among three items $\{i,j,k\}$, it is possible to observe the cyclical preference relation $i\succeq j\succeq k\succeq i$. Such situations arise in certain toy examples, such as the rock–paper–scissor game, as well as in more subtle contexts like Efron’s intransitive dice (Gardner, 1970). More importantly, there is substantial empirical evidence of intransitivity in some real–world domains. Tversky (1969), using binary choices among monetary lotteries, demonstrated that individual choices may violate transitivity. Many papers extending and further formalizing his work have been put forth, e.g., Birnbaum and Gutierrez (2007), Birnbaum and Schmidt (2008), Regenwetter et al. (2011), Cavagnaro and Davis-Stober (2014), Baillon et al. (2015), McCausland et al. (2020) and Alós–Ferrer and Garagnani (2021) and Alós–Ferrer et al. (2022) among many others. The aforementioned papers provide both analytical methods for testing intransitivity and experimental data supporting it. We note that Iverson and Falmagne (1985), who reanalyzed Tversky (1969) data, argued that the evidence in that data set is compatible with transitivity. Similar criticisms directed at other authors are also known, e.g., Loomes et al. (1992). Birnbaum and Schmidt (2010) argue that one reason for these conflicting views is that researchers in the field were unable to discriminate between a systematic violation of transitivity and violations due to random sampling. We believe that these controversies are primarily the result of the lack of a suitable statistical framework and models to analyze cyclical and intransitive PCD.
To date, a principled data–analytical treatment of intransitivity and cyclicality in the context of PCD is lacking. Csato (2019) adopts an axiomatic framework and proves some impossibility theorems. He further discusses restrictions on the preference functions that result in transitivity. Jiang et al. (2011) develop a framework for least squares ranking by viewing cardinal PCGs as edge flows on a comparison graph. Using discrete Hodge theory (Lim, 2020) they argue that edge flows, our $\operatorname{\boldsymbol{\nu}}$, can be decomposed into three orthogonal components which they refer to as the gradient flow, the curl flow and the harmonic flow. Saari (2014) and Saari (2021) develop a linear algebraic formulation and discuss an orthogonal decomposition of $\operatorname{\boldsymbol{\nu}}$ into linear and cyclical components. The aforementioned papers, which provide a conceptual foundation for reasoning about cyclicality, lie outside the statistical literature. However, the phenomenon of cyclicality has long been recognized within the statistical literature as well. Kendall and Smith (1940) developed a coefficient of consistency for detecting cyclicality using combinatorial counting arguments. Their statistical ideas were further developed by Moran (1947), Alway (1962), Knezek et al. (1998), Iida (2009), whereas combinatorial aspects were pursued by Bezembinder (1981), Siraj et al. (2015) and Kulkowski (2018) among many others. It is emphasized that these papers provide methods for detecting, but not for addressing, cyclicality and intransitivity. The literature on statistical models that explicitly account for intransitivity and cyclicality is relatively recent and sparse. We mention the papers of Causeur and Husson (2005), Chen and Joachims (2016), Makhijani and Ugander (2019) and Spearing et al. (2023) who all developed limited parametric models for possibly intransitive paired comparison data.
This paper aims to fill this gap by proposing a new methodology for analyzing PCD with possible cyclicalities and intransitivities. We capitalize on Jiang et al. (2011) topological view and Saari’s (2014) and Saari (2021) linear algebraic insights, upon which we expand. Our main contributions are:
  1. 1.
    A detailed structural analysis of the parameter space $\mathcal{N}$ is provided, including an orthogonal decomposition and explicit spanning sets for both the linear and cyclical subspaces. This yields a unified modeling framework capable of accommodating cyclicality and intransitivity. The concept of a minimal (parsimonious) model is introduced and studied, along with related notions such as the tick–table.
  2. 2.
    A novel model selection technique is proposed. The method leverages $(i)$ geometric insights derived from the study of tick–tables; together with $(ii)$ an understanding of the properties of estimators for models accommodating cyclicality and intransitivity and their accompanying lack–of–fit tests. Minimal models are shown to be both identifiable and consistently selected. Unlike many existing methods, the proposed approach scales quadratically with the number of items, making it well suited for large comparison graphs.
Items (1)–(2) above provide the foundation for a unified methodology for detecting and modeling cyclicality and intransitivity. Consequently, the proposed approach yields more accurate estimation of the underlying preference profile, leading to improved ranking, prediction, and enhanced performance in betting and wagering systems.
The paper is organized as follows. Section 2 presents a decomposition and rigorous analysis of the parameter space $\mathcal{N}$, yielding a modeling framework that accommodates cyclicality and intransitivity. Some remarks on ranking are provided in Section 2.1. Estimators and tests for such models, along with a new model selection strategy, are developed in Section 3. Section 4 contains simulation results, and Section 5 discusses an illustrative example. Section 6 concludes with a summary and discussion. All proofs are collected in the Supplement. The R-scripts for reproducing all experiments are available at github.com/rahulstats/GLM-II.
2 Cyclicality, intransitivity and ranking
The purpose of this section is to provide an understanding of the structure of the parameter space $\mathcal{N}$. This is achieved by decomposing $\mathcal{N}$ into orthogonal components referred to as the linear and cyclical components. Meaningful spanning sets are provided for these components. The recognition that these spanning sets enable the modeling and identification of intransitivity and cyclicality is the key to all further developments. We assume that the comparison graph $\mathcal{G}$ is complete and, consequently, that the full paired preference profile, that is, the entire vector $\operatorname{\boldsymbol{\nu}}$, is available. The relaxation of this assumption is discussed in Section 6.
Let $\operatorname{\boldsymbol{B}}$ be a $\binom{K}{2}\times K$ matrix with columns $\boldsymbol{b}_{1},\ldots,\boldsymbol{b}_{K}$. The $(s,t)^{th}$ lexicographically ordered element of $\boldsymbol{b}_{k}$, denoted $b_{k}(s,t)$, is defined by the relation
$b_{k}(s,t)=\mathbb{I}(s=k)-\mathbb{I}(t=k).$ (4)
It is easily verified that if $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$ it can be expressed as
$\operatorname{\boldsymbol{\nu}}=(\mu_{1}-\mu_{2},\ldots,\mu_{K-1}-\mu_{K})^{\top}=\mu_{1}\boldsymbol{b}_{1}+\ldots+\mu_{K}\boldsymbol{b}_{K}=\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}},$ (5)
and therefore $\mathcal{L}=\mathrm{span}(\boldsymbol{b}_{1},\ldots,\boldsymbol{b}_{K})$. Consequently, $\mathcal{L}$ is a linear subspace of $\mathcal{N}$ for which $\{\boldsymbol{b}_{1},\ldots,\boldsymbol{b}_{K}\}$ is a natural spanning set, see Saari (2021). It is also easy to verify that $\sum_{k=1}^{K}\boldsymbol{b}_{k}=\boldsymbol{0}$ and that the dimension of $\mathcal{L}$ is K1K-1.
Jiang et al. (2011), using graphical combinatorial Hodge theory (see, Lim, 2020), and Saari (2014) and Saari (2021), using the language of linear algebra, established that:
Fact 2.1
If
$\nu_{ij}+\nu_{jk}+\nu_{ki}=0$ (6)
for all $(i,j,k)\in\{1,2,\ldots,K\}^{3}$ then $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$.
The preference vector $\operatorname{\boldsymbol{\nu}}$ is said to be consistent on the triad, or triangle, $(i,j,k)$ if (6) holds, and globally consistent if (6) holds on every triad. Global consistency is equivalent to LST, i.e., the parametrization (2), see Jiang et al. (2011). If (6) does not hold, then a cyclical preference relation between the items $i,j$ and $k$ is obtained. For example, if $(\nu_{ij},\nu_{ik},\nu_{jk})=(1,-1,1)$ then (6) is obviously violated, moreover $i\succ j\succ k\succ i$. The latter preference relationship is captured by a cyclical component of $\mathcal{N}$ that is outside of $\mathcal{L}$. Next, denote the orthogonal complement of $\mathcal{L}$ in $\mathcal{N}$ by $\mathcal{C}$. It follows that any $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ can be uniquely expressed as a sum of two orthogonal vectors, $\operatorname{\boldsymbol{\nu}}_{\rm linear}\in\mathcal{L}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in\mathcal{C}$ for which
$\operatorname{\boldsymbol{\nu}}=\operatorname{\boldsymbol{\nu}}_{\rm linear}+\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ (7)
where $\operatorname{\boldsymbol{\nu}}_{\rm linear}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ capture, respectively, the linear and cyclic aspects of the preference relations (Saari, 2021).
Example 2.1
Suppose that $K=3$ and $\operatorname{\boldsymbol{\nu}}=(1,-1,1)$. Clearly $\nu_{12}+\nu_{23}+\nu_{31}=3$ so $\operatorname{\boldsymbol{\nu}}\notin\mathcal{L}$. By (4) that $\boldsymbol{b}_{1}=(1,1,0)$, $\boldsymbol{b}_{2}=(-1,0,1)$ and $\boldsymbol{b}_{3}=(0,-1,-1)$. Clearly $\operatorname{\boldsymbol{\nu}}^{\top}\boldsymbol{b}_{i}=0$ for all $i$ so $\operatorname{\boldsymbol{\nu}}\in\mathcal{C}$ and consequently $\operatorname{\boldsymbol{\nu}}_{\rm linear}=\operatorname{\boldsymbol{0}}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{\nu}}$. Since $\nu_{12}=1$ we have $1\succ 2$; similarly $\nu_{23}=1$ and $\nu_{13}=-1$ imply that $2\succ 3$ and $3\succ 1$ respectively. Combining these preference relations we find that $1\succ 2\succ 3\succ 1$, i.e., the preference relation is not transitive but cyclical; it is also equivalent to $2\succ 3\succ 1\succ 2$ and $3\succ 1\succ 2\succ 3$, depending on the item is used to anchor the cycle. A cyclic relation with three items is called a cyclic triad.
The next example illuminates the relationship between cyclicality and various form of transitivity.
Example 2.2
Let $K=3$, $\operatorname{\boldsymbol{\nu}}_{\rm linear}=(1,2,1)\in\mathcal{L}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=(1,-1,1)\in\mathcal{C}$. Clearly $\operatorname{\boldsymbol{\nu}}_{\rm linear}$ is associated with the preference relation $1\succ 2\succ 3$ whereas $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ corresponds to the cyclic relation $1\succ 2\succ 3\succ 1$, as demonstrated in Example 2.1. Next, let
$\operatorname{\boldsymbol{\nu}}=\alpha\operatorname{\boldsymbol{\nu}}_{\rm linear}+\beta\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$
where $\alpha,\beta\geq 0$. A bit of algebra shows that (1) is linearly transitive (or LST) if and only if $\beta=0$, strongly transitive (or SST) if $\alpha>2\beta$ and weakly transitive (or WST) if $\alpha>\beta/2$. Moreover, if $\alpha=\beta/2$ then $1\succ 2$, $2\succ 3$ but $1\sim 3$, i.e., $\nu_{13}=0$, whereas if $\alpha<\beta/2$ then the cyclical component dominates the linear component, i.e., $1\succ 2\succ 3\succ 1$. The model becomes purely cyclic if and only if $\alpha=0$.
Example 2.1 shows that when $K=3$ the cyclic relation $1\succ 2\succ 3\succ 1$ can be expressed by the vector (1,1,1)(1,-1,1) which is orthogonal to $\mathcal{L}$ and completes the basis for $\mathcal{N}$. Similarly, for any $K$ the cyclic relation $i\succeq j\succeq k\succeq i$ can be expressed by the vector $\boldsymbol{c}_{(i,j,k)}\in\mathcal{C}$ whose $(s,t)^{th}$ element, in lexicographical order, is
$c_{(i,j,k)}(s,t)=\mathbb{I}((s,t)\in\{(i,j),(j,k),(k,i)\})-\mathbb{I}((s,t)\in\{(j,i),(k,j),(i,k)\}).$ (8)
Thus the $(i,j)^{th}$ and $(j,k)^{th}$ elements of $\boldsymbol{c}_{(i,j,k)}$ are $1$, the $(i,k)^{th}$ element is 1-1 and all other elements are 0. Equation (8) is significant, as it demonstrates that cyclical relations admit a linear algebraic representation. It is easy to verify that
$\boldsymbol{c}_{(i,j,k)}=\boldsymbol{c}_{(j,k,i)}=\boldsymbol{c}_{(k,i,j)}=-\boldsymbol{c}_{(j,i,k)}=-\boldsymbol{c}_{(i,k,j)}=-\boldsymbol{c}_{(k,j,i)},$
hence there are only two types of cycles among any three items; i.e., $i\succeq j\succeq k\succeq i$ or its reverse $k\succeq j\succeq i\succeq k$. Both cycles can be represented by a single vector, i.e., they lie in the same one dimensional linear space. As a consequence, only one element among those in the display above is necessary for modeling cyclicality among $\{i,j,k\}$; for convenience, we shall choose $\boldsymbol{c}_{(i,j,k)}$ with $i<j<k$. Further note that a cyclic relation on $(i,j,k)$ can be expressed as superposition of the cyclic relations $(s,i,j),(s,j,k)$ and $(s,i,k)$ for some $s\notin\{i,j,k\}$, i.e.,
$\boldsymbol{c}_{(i,j,k)}=\boldsymbol{c}_{(s,i,j)}+\boldsymbol{c}_{(s,j,k)}-\boldsymbol{c}_{(s,i,k)}.$ (9)
Note that the cyclic relations on the right-hand side of (9) honor the preference relation in $(i,j,k)$. Equation (9) plays an important role in later developments.
The following was established by Saari (2014) and Saari (2021),
Fact 2.2
The set $\{\boldsymbol{c}_{(i,j,k)}\}_{1\leq i<j<k\leq K}$, whose cardinality is $\binom{K}{3}$, spans $\mathcal{C}$. Furthermore,
$\dim(\mathcal{C})=\binom{K}{2}-(K-1)=\binom{K-1}{2}$.
Consequently, for all $K\geq 4$ the spanning set is larger, often much larger, than the dimension of $\mathcal{C}$. Let $\operatorname{\boldsymbol{C}}$ be a $\binom{K}{2}\times\binom{K}{3}$ matrix whose columns are $\{\boldsymbol{c}_{(i,j,k)}\}_{1\leq i<j<k\leq K}$. Note that the rows of $\operatorname{\boldsymbol{C}}$ are indexed by the edges $(i,j)$ and the columns are indexed by the triads $(i,j,k)$. It is now clear that if $\operatorname{\boldsymbol{\nu}}\in\mathcal{C}$ we have
$\operatorname{\boldsymbol{\nu}}=\sum_{1\leq i<j<k\leq K}\gamma_{ijk}\,\boldsymbol{c}_{(i,j,k)}=\operatorname{\boldsymbol{C}}\boldsymbol{\gamma},$ (10)
where $\gamma_{ijk}$ is the coefficient associated with the cyclical triad (i,j,k){(i,j,k)} and $\boldsymbol{\gamma}\in\mathbb{R}^{\binom{K}{3}}$. Hence, for any $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ there exist $\operatorname{\boldsymbol{\mu}}$ and $\boldsymbol{\gamma}$ such that
$\operatorname{\boldsymbol{\nu}}=\operatorname{\boldsymbol{\nu}}_{\rm linear}+\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}+\operatorname{\boldsymbol{C}}\boldsymbol{\gamma}.$ (11)
Since $\mathcal{L}\perp\!\!\!\perp\mathcal{C}$, the linear and cyclical components of $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ are given by
$\operatorname{\boldsymbol{\nu}}_{\rm linear}=\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{\nu}}~\mbox{and}~~\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{C}}(\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{C}})^{+}\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{\nu}}.$ (12)
Thus, the linear and cyclical components of $\operatorname{\boldsymbol{\nu}}$ can always be identified. Further note that when $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{0}$ then $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$ is LST, i.e., it is transitive; however, if $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\neq\boldsymbol{0}$ then $\operatorname{\boldsymbol{\nu}}$ may or may not be transitive, cf. Example 2.2. Before proceeding, we provide a remark concerning an alternative parameterization.
Remark 2.1
Jiang et al. (2011) propose an alternative, equivalent, matrix–based parameterization. In their formulation, the unknown parameter is not a vector $\operatorname{\boldsymbol{\nu}}\in\mathbb{R}^{\binom{K}{2}}$ but a $K\times K$ skew–symmetric matrix $\boldsymbol{\mathcal{V}}=(\nu_{ij})$ whose elements satisfy $\nu_{ij}=-\nu_{ji}$ whenever $i\neq j$ and 0 otherwise. Let $\mathfrak{L},\,\mathfrak{C}$ and $\mathfrak{N}$ be the matrix equivalents of $\mathcal{L},\,\mathcal{C}$ and $\mathcal{N}$ respectively. It is not hard to see that $\mathfrak{N}$ is spanned by the matrices $\{\boldsymbol{E}_{ij}\}_{1\leq i,\,j\leq K}$, $\boldsymbol{E}_{ij}=\boldsymbol{e}_{i}\boldsymbol{e}_{j}^{\top}-\boldsymbol{e}_{j}\boldsymbol{e}_{i}^{\top}$ where $\boldsymbol{e}_{k}$ is the $k^{th}$ standard basis vector for $\mathbb{R}^{K}$. Moreover, $\{\boldsymbol{E}_{ij}\}_{1\leq i<j\leq K}$ forms the standard basis of $\mathfrak{N}$. Further note that the subspace $\mathfrak{L}\subset\mathfrak{N}$ is the space of matrices of the form $\boldsymbol{1}\operatorname{\boldsymbol{\mu}}^{\top}-\operatorname{\boldsymbol{\mu}}\boldsymbol{1}^{\top}$ with $\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}$. Observe that $\{\boldsymbol{B}_{i}\}_{1\leq i\leq K-1}$ with $\boldsymbol{B}_{i}=\boldsymbol{1}\boldsymbol{e}_{i}^{\top}-\boldsymbol{1}\boldsymbol{e}_{K}^{\top}-\boldsymbol{e}_{i}\boldsymbol{1}^{\top}+\boldsymbol{e}_{K}\boldsymbol{1}^{\top}$ forms a basis for $\mathfrak{L}$, cf. (5). Next, with some algebra, it can be shown that $\mathfrak{C}$ is spanned by $\{\boldsymbol{F}_{ij}:1\leq i,\,j\leq K-1\}$ where $\boldsymbol{F}_{ij}=\boldsymbol{E}_{ij}-\boldsymbol{E}_{iK}+\boldsymbol{E}_{jK}$. Moreover, a basis is given by $\{\boldsymbol{F}_{ij}:1\leq i<j\leq K-1\}$. Note that $\boldsymbol{B}_{i}$ and $\boldsymbol{F}_{ij}$ are matrix–based equivalences of $\boldsymbol{b}_{i}$ and $\boldsymbol{c}_{(i,j,k)}$. The two parameterizations are equivalent. However, its vectorized form naturally fits within the linear model framework, as analyzed in Section 3, thus facilitating the identification of cyclicalities and possible intransitivities.
We continue exploring the parameterization (11). First note that $\operatorname{\boldsymbol{\mu}}$ is identifiable provided that a constraint, such as $\sum_{i=1}^{K}\mu_{i}=0$, is imposed, see Singh et al. (2025). Next, observe that Equation (9) and Fact 2.2 imply that whenever $K\geq 4$ the set $\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}=\{\boldsymbol{\gamma}\in\mathbb{R}^{\binom{K}{3}}:\,\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{C}}\boldsymbol{\gamma}\}$ has a cardinality greater than one for every $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in\mathcal{C}$. This implies that $\boldsymbol{\gamma}$ cannot be identified in the usual sense.
Definition 2.1
For $\boldsymbol{\gamma}\in\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ let $s(\boldsymbol{\gamma})=\sum\mathbb{I}(\gamma_{ijk}\neq 0)$ denote the size of the model associated with $\boldsymbol{\gamma}$. A model $\boldsymbol{\gamma}\in\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ is said to be minimal if $s(\boldsymbol{\gamma})\leq s(\boldsymbol{\gamma}^{\prime})$ for all $\boldsymbol{\gamma}^{\prime}\in\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$.
Note that in the context of Definition 2.1 the terms model, representation, and parameter are used interchangeably. To illustrate the concept of a minimal model, or minimal representation, consider the case where $\operatorname{\boldsymbol{\nu}}_{\rm linear}=(1,2,3,1,2,1)\in\mathcal{L}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=(0,0,0,1,-1,1)\in\mathcal{C}$. Clearly $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(2,3,4)}$ so $\boldsymbol{\gamma}_{1}=(0,0,0,1)\in\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$, by (9) $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,2,3)}+\boldsymbol{c}_{(1,3,4)}-\boldsymbol{c}_{(1,2,4)}$ so $\boldsymbol{\gamma}_{2}=(1,1,-1,0)\in\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$. Clearly $\rm{dim}(\mathcal{C})=6$ so the set $\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ contains many other elements. Since $s(\boldsymbol{\gamma}_{1})=1$ and $s(\boldsymbol{\gamma})>1$ for all $\boldsymbol{\gamma}\neq\boldsymbol{\gamma}_{1}$ and $\boldsymbol{\gamma}\in\Gamma_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ it follows that $\boldsymbol{\gamma}_{1}$ is the unique minimal model.
Minimal models are parsimonious, that is, they include only those cyclical triads that are necessary for explaining the observed preference relations. We shall always assume that the data–generating mechanism is associated with a minimal model; consequently, non–minimal models, which incorporate some unnecessary triads, are not of interest to us. The assumption that the observed data is generated by a minimal model is reasonable and useful but generally not testable. We also note that the construction of parsimonious models requires that all possible cyclical triads in the set $\{\boldsymbol{c}_{(i,j,k)}\}_{1\leq i<j<k\leq K}$ be considered. We associate each minimal model with the set $\mathcal{C}_{\boldsymbol{\gamma}}$ of columns of $\operatorname{\boldsymbol{C}}$ corresponding to the non–zero coefficients of $\boldsymbol{\gamma}$.
Definition 2.2
The span and support of a minimal model are
${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}})={\rm span}(\{\boldsymbol{c}_{(i,j,k)}:\,\gamma_{ijk}\neq 0\})~~\mbox{and}~~{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})=\{(i,j):\boldsymbol{c}_{(i,j,k)}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}})\}.$
Ideally, minimal models would have a unique support and be identifiable from their span. Unfortunately, this is not always the case.
Example 2.3
For any $K\geq 6$ let $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,2,3)}-\boldsymbol{c}_{(1,2,4)}+\boldsymbol{c}_{(1,3,6)}-\boldsymbol{c}_{(2,3,5)}$. We refer to this model as $\mathcal{C}_{\boldsymbol{\gamma}_{1}}$. Using (9) we also express $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ as $\boldsymbol{c}_{(1,3,4)}+\boldsymbol{c}_{(1,3,6)}+\boldsymbol{c}_{(2,3,4)}-\boldsymbol{c}_{(2,3,5)}$ which we refer to as $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$. Thus, the minimal model is not unique. Moreover, ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\neq{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ and ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\neq{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ so minimal models may be associated with different linear subspaces and supports.
Example 2.3 seems to suggest that minimal models, as well as the associated linear spaces and supports, can not be identified. This view, however, is overly pessimistic. Example 2.3 is possible due to two special features: $(i)$ the vector $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ is a linear combination of triads with high degree of overlap, such as $(1,2,3)$ and $(1,2,4)$ and; $(ii)$ the coefficients of the linear combinations generating $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ are all of the same magnitude (equal to unity). Consequently, elements in the linear expansions $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ are “zeroed out” precluding a unique minimal model. In practice, these conditions are not expected to hold simultaneously.
Theorem 2.1
For minimal models $\boldsymbol{\gamma}_{1}$ and $\boldsymbol{\gamma}_{2}$ we have $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\cap{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Moreover, if ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ then ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Furthermore, suppose the following conditions are satisfied:
  • (i)
    For all $(i,j)\in\operatorname{supp}(\mathcal{C}_{\boldsymbol{\gamma}})$, we have $\nu_{\text{cyclic},ij}\neq 0$;
  • (ii)
    If $\operatorname{supp}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})=\operatorname{supp}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$, then $\operatorname{span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})=\operatorname{span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$,
then $\mathcal{C}_{\boldsymbol{\gamma}_{1}}$ and $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$ span the same subspace. Finally, if any two triads in a minimal model share at most one index then the minimal model is unique.
Theorem 2.1 indicates that in practice minimal models are likely to be unique and hence identifiable from their support. Moreover, when a unique minimal model does not exist, the span of all possible minimal models is often the same.
Note that $\nu_{{\rm cyclic},ij}$, the $(i,j)^{th}$ component of $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$, is given by
$\nu_{{\rm cyclic},ij}=\sum_{1\leq s<t<u\leq K}c_{(s,t,u),ij}\gamma_{stu}.$ (13)
The condition that $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ implies that $\nu_{{\rm cyclic},ij}\neq 0$ and means that the sum in (13) is non–zero for $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$. This will happen if the absolute values of the nonzero elements of $\boldsymbol{\gamma}$ are all distinct, as are all of their partial sums. This condition is very likely to hold in practice.
Next, we explore the relationship between $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ and minimal models. It can be easily verified that for any triad $(i,j,k)$
$\nu_{ij}+\nu_{jk}+\nu_{ki}=\nu_{{\rm cyclic},ij}+\nu_{{\rm cyclic},jk}+\nu_{{\rm cyclic},ki}.$
Thus, cyclic triads, i.e., triads violating (6), can also be identified using $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$. It is emphasized that some triads violating (6) may not be part of any minimal model. For example, if $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,2,3)}$, then any triad $(i,j,k)$ sharing two indices with the triad $(1,2,3)$, e.g., $(1,2,4)$, will be a cyclic triad, but not be part of any minimal model.
Let $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ denote a table with rows and columns indexed by the edges and triads as the matrix $\operatorname{\boldsymbol{C}}$. If the $(s,t)^{th}$ element of $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ is non–zero we place a tick mark in every column in which the combination $(s,t)$ occurs, i.e., in any column indexed by $(s,t,u)$, for some $u\notin\{s,t\}$, or possible permutations thereof. Consequently the columns of $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$, referred to as the tick–table, may have $0,1,2$ or $3$ ticks. For example, let $K=4$ and set $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,2,3)}=(1,-1,0,1,0,0)$, then $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ is given by:
123 124 134 234
12
13
14
23
24
34
Note that in this tick table only the $(1,2,3)^{th}$ column, associated with the minimal model, while all other columns have one tick each.
Theorem 2.2
If there is a minimal model $\boldsymbol{\gamma}$ with some 0–tick triads then there exists another minimal model $\boldsymbol{\gamma}^{\prime}$ for which no triad has 0-ticks. Additionally, if any two triads in a minimal model share at most one index then only triads in the minimal model will have 33–ticks. However not all 33–tick triads are members of a minimal model. Finally, if the minimal model is unique and some of its triads share two indices then some of the triads may have less than 33–ticks.
Theorem 2.2 shows that when the minimal model is unique it can be identified using the tick–table $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$, whereas if the minimal model is not unique then some minimal model can always be identified.
2.1 Ranking
As noted in Section 1, if there are no cyclicalities, i.e., $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$ then $i\succeq j$ if and only if $\mu_{i}\geq\mu_{j}$. As expected, ranking is more complicated when $\operatorname{\boldsymbol{\nu}}\notin\mathcal{L}$.
Theorem 2.3
Suppose (11) holds and let $\operatorname{\boldsymbol{\nu}}$ be transitive. If so $i\succeq j$ implies that $\mu_{i}\geq\mu_{j}$ if $\operatorname{\boldsymbol{\nu}}$ is strongly transitive but not necessarily if $\operatorname{\boldsymbol{\nu}}$ is weakly transitive.
We conclude that if $\operatorname{\boldsymbol{\nu}}$ is strongly transitive, then the merits can be used for ranking; but not so under weak stochastic transitivity. Nevertheless, items can always be ranked even under weak transitivity.
Next consider intransitive models. Saari (2014) and Saari (2021) addressed this issue by treating cyclicalities as noise. For $i=1,\ldots,K$ he proposed to compute the pseudo merit of the $i^{th}$ item by
$\mu_{i}^{*}=\sum_{j=1}^{K}\nu_{ij}$ (14)
and rank $i$ above $j$ if the pseudo–merits satisfy $\mu_{i}^{*}\geq\mu_{j}^{*}$. This approach is a relative of the well–known row–sum method (Huber, 1963). Further note that for complete and balanced comparison graphs ranking based on $\mu_{i}$ and $\mu_{i}^{*}$ are identical; see the proof of Theorem 5.1 in Singh et al. (2025). In general, the pseudo–merits can be defined by $\sum_{j=1}^{K}\theta_{ij}\nu_{ij}$. In the same spirit one can compute
$\mu_{i}^{**}=\sum_{j=1}^{K}\mathbb{I}(\nu_{ij}>0)$ (15)
which we refer to as the dominance score. Notice that the dominance score is reminiscent of the well known Borda–Count advocated among some voting theorists (e.g., Saari 2023). Using the dominance score rank $i$ above $j$ if $\mu_{i}^{**}>\mu_{j}^{**}$. Ties among the dominance score can be further broken by repeatedly computing (15) on sets of the form $\{k:\mu_{i}^{**}=\mu_{j}^{**}\}$. The following proposition suggests that dominance scores can be used for ranking.
Proposition 2.1
If $\mu_{i}^{**}\neq\mu_{j}^{**}$ for all $i\neq j$ then $\operatorname{\boldsymbol{\nu}}$ is weakly transitive.
Equations (14) and (15) map $\mathcal{N}\mapsto P_{n}$, where $P_{n}$ is the space of all permutations of $\{1,\ldots,n\}$. Such mappings are of course not completely satisfactory. In fact, as noted in Section 1, cyclicalities are often an important aspect of the data and should not be averaged out and discarded.
Finally we note that global non–transitivity may allow in some settings for partially transitivity, i.e., the set $\{1,\ldots,K\}$ may be partitioned into $g$ groups satisfying
$\langle i_{1,1},\ldots,i_{1,K_{1}}\rangle\succeq\langle i_{2,1},\ldots,i_{2,K_{2}}\rangle\succeq\ldots\succeq\langle i_{g,1},\ldots,i_{g,K_{g}}\rangle,$ (16)
where $\langle i_{1,1},\ldots,i_{1,K_{1}}\rangle\succeq\langle i_{2,1},\ldots,i_{2,K_{2}}\rangle$ means $i_{1,l_{1}}\succeq i_{2,l_{2}}$ for any $1\leq l_{1}\leq K_{1}$ and $1\leq l_{2}\leq K_{2}$. If strong stochastic transitivity holds among any three items inhabiting three different groups, then by Theorem 2.3 the merits of these items can be used to rank them. If (16) holds with $G=1$ then $\operatorname{\boldsymbol{\nu}}$ is strongly intransitive although subsets of transitive items exist.
3 Estimation and model selection
In this section, a statistical methodology that accommodates intransitivity and cyclicality is developed. It is assumed that for all pairs $1\leq i\neq j\leq K$ we have $n_{ij}>0$, i.e., the PCG is complete.
3.1 Preliminaries
If $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$, we say that the reduced model holds. If so $\boldsymbol{\gamma}=\boldsymbol{0}$ and it is well known that the least square estimator (LSE) of $\operatorname{\boldsymbol{\mu}}$ assuming $\sum_{i=1}^{K}\mu_{i}=0$ is given by $\widehat{\operatorname{\boldsymbol{\mu}}}=\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{S}}$ where $\operatorname{\boldsymbol{N}}^{+}$ is the Moore–Penrose inverse of $\operatorname{\boldsymbol{N}}$, the Laplacian of the graph $\mathcal{G}$ and $\operatorname{\boldsymbol{S}}=(S_{1},\ldots,S_{K})^{\top}$, is a vector of sums where $S_{i}=\sum_{j\neq i}S_{ij}$ and $S_{ij}=\sum_{k=1}^{n_{ij}}Y_{ijk}$. It can be further shown that under the conditions stated in Singh et al. (2025) when $n\to\infty$ we have
$\sqrt{n}(\widehat{\operatorname{\boldsymbol{\mu}}}-\operatorname{\boldsymbol{\mu}})\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\sigma^{2}\boldsymbol{\Theta}^{+}),$
where $\Rightarrow$ denotes convergence in distribution, $\boldsymbol{\Theta}=\lim_{n}\operatorname{\boldsymbol{N}}/n$ is the scaled limit of the Laplacian, and $n=\sum_{1\leq i<j\leq K}n_{ij}$ is the total number of paired comparisons. When $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$, without any further restrictions, we say that the full model holds. It is clear that the LSE for the full model is $\widehat{\operatorname{\boldsymbol{\nu}}}=\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}$ where $\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}$ is a $|\operatorname{\boldsymbol{\nu}}|$ vector whose $(i,j)^{th}$ element is $\overline{S}_{ij}=S_{ij}/n_{ij}$ if $n_{ij}>0$ and set, arbitrarily, to be 0 otherwise. Furthermore, if $n_{ij}=O(n)$ for all $1\leq i<j\leq K$ then
$\sqrt{n}(\widehat{\operatorname{\boldsymbol{\nu}}}-\operatorname{\boldsymbol{\nu}})\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\sigma^{2}\,\boldsymbol{\Xi}^{+}),$ (17)
where $\boldsymbol{\Xi}$ be the $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ diagonal matrix whose $(i,j)^{th}$ diagonal element is $\theta_{ij}=\lim{n_{ij}/n}$. For more on the relationship between the reduced and full model see Appendix B in the Supplement. In the next subsection, we develop and study an estimation procedure that that employs the parameterization (11); although the method is valid for the full model it is advantageous especially for intermediate models. An intermediate model is a model in which $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$ where $\mathcal{L}\subsetneq\mathcal{S}\subsetneq\mathcal{N}$ and $\subsetneq$ denotes strict inclusion. Thus, $\mathcal{S}={\rm span}({\operatorname{\boldsymbol{B}}},{\operatorname{\boldsymbol{C}}}_{s})$ where $\operatorname{\boldsymbol{C}}_{s}$ is a sub–matrix of $\operatorname{\boldsymbol{C}}$ with at most $\binom{K-1}{2}-1$ columns.
3.2 Estimation
Let $\operatorname{\boldsymbol{Y}}$ be the $n\times 1$ vector of outcomes arranged lexicographically, i.e.,
$\operatorname{\boldsymbol{Y}}=(Y_{121},\ldots,Y_{12n_{12}},Y_{131},\ldots,Y_{13n_{13}},\ldots,Y_{K-1,K1},\ldots,Y_{K-1,Kn_{K-1,K}})^{\top}.$
Let $\operatorname{\boldsymbol{L}}$ be the $n\times|\operatorname{\boldsymbol{\nu}}|$ matrix whose elements $L_{ij}\in\{0,1\}$ are such that if the $s^{th}$ element of $\operatorname{\boldsymbol{Y}}$ is $Y_{ijk}$ and $t^{th}$ element of $\operatorname{\boldsymbol{\nu}}$ is $\nu_{ij}$ then $L_{st}=1$ and $L_{su}=0$ for all $u\neq t$. It is easy to verify that $\mathbb{E}(\operatorname{\boldsymbol{Y}})=\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}$. It follows that we can rewrite (1) using (11) as
$\operatorname{\boldsymbol{Y}}=(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}},\,\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}})\begin{pmatrix}\operatorname{\boldsymbol{\mu}}\\ \boldsymbol{\gamma}\end{pmatrix}+\boldsymbol{\epsilon}.$ (18)
Clearly (18) is a full model. Let $Q(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma})=(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{\nu}}(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}))^{\top}(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{\nu}}(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}))$ denote the sum of squares associated with (18). As noted, the model (18) is not identifiable with respect to $(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma})$, precluding a unique LSE. We circumvent this problem by defining
$(\widehat{\operatorname{\boldsymbol{\mu}}},\widehat{\boldsymbol{\gamma}})=\arg\min_{\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}}\{Q(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}):\min\|(\operatorname{\boldsymbol{\mu}}^{\top},\boldsymbol{\gamma}^{\top})^{\top}\|\},$ (19)
[…]
Theorem 3.1
Suppose that $n_{ij}>0$ for all $1\leq i<j\leq K$. Then the LSE (19) is unique and given by
[…]
and
$\widehat{\boldsymbol{\gamma}}=(\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{L}}^{\top}(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}})^{+}\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{L}}^{\top}(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{Y}}.$ (21)
[…]
Example 3.1
Consider a PCG on which $\operatorname{\boldsymbol{\nu}}=\operatorname{\boldsymbol{\nu}}_{\rm linear}+\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ with
$\operatorname{\boldsymbol{\nu}}_{\rm linear}=(1,2,3,4,1,2,3,1,2,1)^{\top}\mbox{ and }\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=(1,-1,0,0,1,0,0,0,0,0)^{\top},$
i.e., the merits are $(-2,-1,0,1,2)^{\top}$ and the minimal model for $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ is comprised of a single vector $\boldsymbol{c}_{(1,2,3)}$ with coefficient $\gamma_{123}=1$ and therefore norm $1$. Further note that if $n_{ij}=m\to\infty$ for all $1\leq i,j\leq K$ then the LSE (21) converges to (0.6,0.2,0.2,0.2,0.2,0,0.2,0.2,0,0)(0.6,0.2,0.2,-0.2,-0.2,0,0.2,0.2,0,0) whose norm is $\sqrt{0.6}$. Thus the minimum norm solution tends to assign a non–zero value to too many elements of $\boldsymbol{\gamma}$ and consequently will not recover the minimal model.
Next, consider the situation in which the minimal model is known to be an intermediate model and consists of the triads $(i_{1},j_{1},k_{1}),\ldots,(i_{r},j_{r},k_{r})$ with $r\leq\binom{K-1}{2}-1$. Let $\operatorname{\boldsymbol{C}}_{s}$ denote the matrix whose columns are $\boldsymbol{c}_{(i_{1},j_{1},k_{1})},\ldots,\boldsymbol{c}_{(i_{r},j_{r},k_{r})}$ and let $\boldsymbol{\gamma}_{s}$ be the corresponding vector of coefficients. In this case, we fit the model (18) replacing $\operatorname{\boldsymbol{C}}$ with $\operatorname{\boldsymbol{C}}_{s}$ and $\boldsymbol{\gamma}$ with $\boldsymbol{\gamma}_{s}$. The LSE is given by
$(\widehat{\operatorname{\boldsymbol{\mu}}},\widehat{\boldsymbol{\gamma}}_{s})=\arg\min_{\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}_{s}}\{Q(\operatorname{\boldsymbol{\mu}},\boldsymbol{\gamma}_{s}):\operatorname{\boldsymbol{1}}1^{\top}\operatorname{\boldsymbol{\mu}}=0\}.$ (22)
Theorem 3.2
Assume that the errors $\epsilon_{ijk}$ are IID with zero mean and finite variance $\sigma^{2}$. Then:
  1. 1.
    If $n_{ij}>0$ for all $1\leq i<j\leq K$ then the LSEs for $\operatorname{\boldsymbol{\mu}}$ and $\boldsymbol{\gamma}_{s}$ are given by substituting $\operatorname{\boldsymbol{C}}_{s}$ for $\operatorname{\boldsymbol{C}}$ in (20) and (21). The resulting estimators are unbiased.
  2. 2.
    If $n_{ij}\to\infty$ for $1\leq i<j\leq K$, then $(\widehat{\operatorname{\boldsymbol{\mu}}},\widehat{\boldsymbol{\gamma}}_{s})\to(\operatorname{\boldsymbol{\mu}},{\boldsymbol{\gamma}}_{s})$ in probability as $n\to\infty$.
  3. 3.
    If $n_{ij}\to\infty$ and $n_{ij}=O(n)$ for all $1\leq i<j\leq K$ then
    $\sqrt{n}\begin{pmatrix}\widehat{\operatorname{\boldsymbol{\mu}}}-\operatorname{\boldsymbol{\mu}}\\ \widehat{\boldsymbol{\gamma}}_{s}-\boldsymbol{\gamma}_{s}\end{pmatrix}\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\boldsymbol{\Sigma}_{\mu,\gamma_{s}}),$
    where
    $\boldsymbol{\Sigma}_{\mu,\gamma_{s}}=\sigma^{2}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\end{pmatrix}\boldsymbol{\Xi}^{3}(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}$.
Theorem 3.2 shows that when the minimal model is known, the LSE $(\widehat{\operatorname{\boldsymbol{\mu}}},\widehat{\boldsymbol{\gamma}}_{s})$ is unbiased, consistent, and asymptotically normal.
Next, a lack–of–fit test is investigated. Consider testing $H_{0}:~\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$ against $H_{1}:~\operatorname{\boldsymbol{\nu}}\notin\mathcal{S}$. Motivated by Jiang et al. (2011) we define
$R_{n,\mathcal{S}}=\boldsymbol{U}_{n,\mathcal{S}}^{\top}\boldsymbol{U}_{n,\mathcal{S}},$
where $\boldsymbol{U}_{n,\mathcal{S}}=\operatorname{\boldsymbol{D}}^{1/2}(\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}-\widehat{\operatorname{\boldsymbol{\nu}}}_{s})$, and $\widehat{\operatorname{\boldsymbol{\nu}}}_{s}=\operatorname{\boldsymbol{B}}\widehat{\operatorname{\boldsymbol{\mu}}}+\operatorname{\boldsymbol{C}}_{s}\widehat{\boldsymbol{\gamma}}_{s}$ is the LSE under $H_{0}$. Thus $R_{n,\mathcal{S}}$ is the squared norm of the residual.
Theorem 3.3
Assume that $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$ and the condition of Theorem 3.2 hold. If so as $n\to\infty$
$R_{n,\mathcal{S}}\Rightarrow\sum_{i=1}^{t}{\lambda}_{i}Z_{i}^{2},$
where $t={\rm rank}({\boldsymbol{\Psi}}_{\mathcal{S}})=\binom{K}{2}-(K+r-1)$, $r$ is the number of columns in $\operatorname{\boldsymbol{C}}_{s}$, $Z_{1},\ldots,Z_{t}$ are independent $\mathcal{N}(0,1)$ RVs and ${\lambda}_{1},\ldots,{\lambda}_{t}$ are the non–zero eigenvalues of the $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ matrix ${\boldsymbol{\Psi}}_{\mathcal{S}}=\sigma^{2}\,\boldsymbol{\Xi}^{1/2}\boldsymbol{M}\,\boldsymbol{\Xi}^{+}\boldsymbol{M}^{\top}\boldsymbol{\Xi}^{1/2}$, where
$\boldsymbol{M}=\boldsymbol{I}-(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\end{pmatrix}\boldsymbol{\Xi}.$
It can be further shown that under local alternatives, see Proposition A.1 in the Supplement, we have
$R_{n,\mathcal{S}}\Rightarrow\sum_{i=1}^{t}{\lambda}_{i}(Z_{i}+\gamma_{i})^{2},$
where $\gamma_{1},\ldots,\gamma_{t}$ are the elements of the vector $\boldsymbol{\gamma}_{s}=\boldsymbol{O}_{s}(\boldsymbol{\Psi}_{s}^{+})^{1/2}\boldsymbol{\Xi}^{1/2}\boldsymbol{\delta}$ which correspond to the nonzero eigenvalues of $\boldsymbol{\Psi}$. Here $\boldsymbol{O}_{s}$ is the orthonormal matrix whose columns are the eigenvectors of $\boldsymbol{\Psi}_{s}$. We further note that, as an alternative to $R_{n,\mathcal{S}}$ a Wald–type tests of the form
$W_{n,\mathcal{S}}=\boldsymbol{U}_{n,\mathcal{S}}^{\top}\boldsymbol{\Psi}_{n,\mathcal{S}}^{+}\boldsymbol{U}_{n,\mathcal{S}}\Rightarrow\chi^{2}(\binom{K}{2}-(K+r-1)),$
can also be employed.
3.3 Model selection
[…]
In this section, two new methods for model selection, tailored to the problem at hand, are proposed. The first is motivated by Fact 2.1.Observe that cyclical triads can be identified by testing
$H_{0}:\nu_{ij}+\nu_{jk}+\nu_{ki}=0\text{ against }H_{1}:\nu_{ij}+\nu_{jk}+\nu_{ki}\neq 0.$ (23)
A natural test statistic for (23) is
$T_{n}=T_{n}(i,j,k)=\frac{1}{\sqrt{1/n_{ij}+1/n_{jk}+1/n_{ki}}}(\widehat{\nu}_{ij}+\widehat{\nu}_{jk}+\widehat{\nu}_{ki}),$ (24)
for which:
Proposition 3.1
Let $n_{ij},n_{jk},n_{ik}\to\infty$ at the same rate. Then, under $H_{0}$, $T_{n}\Rightarrow\mathcal{N}(0,\sigma^{2})$ where $\sigma^{2}$ can be consistently estimated by $\widehat{\sigma}^{2}=\sum_{1\leq i<j\leq K}\sum_{k=1}^{n_{ij}}(Y_{ijk}-\widehat{\nu}_{ij})^{2}/n$. Under local alternatives of the form $\nu_{ij}+\nu_{jk}+\nu_{ki}=\delta\sqrt{1/n_{ij}+1/n_{jk}+1/n_{ki}}$ we have $T_{n}\Rightarrow\mathcal{N}(\delta,\sigma^{2})$.
Let $P_{(i,j,k)}$ denote the $p$-value associated with the statistic $T_{n}(i,j,k)$. Apply the test (24) to all triads and arrange these pp–value according to their significance. Suppose the order is $(i_{1},j_{1},k_{1}),\ldots,(i_{r},j_{r},k_{r})$. Let $\mathcal{S}_{1}=\mathcal{L}$, and for $a=2,\ldots,r$ let $\mathcal{S}_{a+1}=\{(i_{1},j_{1},k_{1}),\ldots,(i_{a},j_{a},k_{a})\}$. We now introduce a model selection method referred to as forward step–wise triad selection (FSTS).
Forward Stepwise Triad Selection ($\rm FSTS$)
1:Test whether $H_{0}:\mathcal{S}=\mathcal{S}_{1}$ holds. If the null is not rejected set $\widehat{\mathcal{S}}=\mathcal{S}_{1}$ and stop. Otherwise, set $a=2$.
2:Test whether $H_{0}:\mathcal{S}=\mathcal{S}_{a}$ holds. If the null is not rejected set $\widehat{\mathcal{S}}=\mathcal{S}_{a}$ and stop. Otherwise,
3:Set $a=a+1$ and repeat Step 2.
FSTS requires $\binom{K}{3}$ simple tests. Additionally, at each stage of the model fitting process, a test for lack of fit, as prescribed in Theorem 3.3, is performed. Clearly FSTS is a relative of forward step–wise regression (FSR); the difference being that the order by which covariates are added to the model is determined by the pp–values $P_{(i,j,k)}$ rather than by a sequence of F–test as in FSR.
Numerical experiments indicate that a small p-value $P_{(i,j,k)}$ does not necessarily imply that the triad $(i,j,k)$ is associated with a minimal model even in large samples. This observation motivates the introduction of a second method which builds on the insights of Section 2, specifically Theorems 2.1 and 2.2. The proposed method is, like FSTS, a relative of FSR, and referred to as forward tick–based selection (FTBS). In contrast with standard methods it does not directly utilize features of the design matrix but instead relies on an estimated tick table. To fix ideas, let $\mathcal{S}\subseteq\mathcal{N}$ denote the smallest subspace which contains all possible minimal models and let $\widehat{\mathcal{S}}$ denote the selected model. To find $\widehat{\mathcal{S}}$ first compute $\widehat{\operatorname{\boldsymbol{\mu}}}$ and $\widehat{\operatorname{\boldsymbol{\nu}}}$, as described in Section 3.1, and $\widehat{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}$ as described in (12). Let $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}}$ be the tick–table corresponding to $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}$ where
[…]
Here $\phi_{ij}(\widehat{\operatorname{\boldsymbol{\nu}}}_{{\rm cyclic}})=1$ if $H_{0}^{(i,j)}:\nu_{{\rm cyclic},ij}=0$ is rejected 0 otherwise. The family of hypotheses $\{H_{0}^{(i,j)}\}_{1\leq i<j\leq K}$ is tested at the level $\alpha$ while controlling either the family–wise error rate (FWER) or the false discovery rate (FDR), see Stoica and Babu (2022) for more details on testing based approachs to model selection.
Remark 3.1
[…]
Let $\mathcal{C}_{0},\ldots,\mathcal{C}_{3}$ denote the collection of columns of $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ with $0,\ldots,3$ ticks respectively. Further let $\mathcal{C}_{1,\rm lin}$, $\mathcal{C}_{2,\rm lin}$ and $\mathcal{C}_{3,\rm lin}$ be sets of linearly independent elements in $\mathcal{C}_{1}$, $\mathcal{C}_{2}$ and $\mathcal{C}_{3}$ respectively. Next let $\mathcal{C}_{2,\rm lin}^{+}$ be the set of elements in $\mathcal{C}_{2,\rm lin}$ which are not spanned by $\mathcal{C}_{3,\rm lin}$ and similarly let $\mathcal{C}_{1,\rm lin}^{+}$ be the set of elements in $\mathcal{C}_{1,\rm lin}$ which are not spanned by $\mathcal{C}_{3,\rm lin}\cup\mathcal{C}_{2,\rm lin}^{+}$. Further define a nested sequence of models $\mathcal{S}_{1}\subset\ldots\subset\mathcal{S}_{4}$ where $\mathcal{S}_{1}=\mathcal{L}$, $\mathcal{S}_{2}=\mathcal{L}\cup\mathcal{C}_{3,\rm lin}$, $\mathcal{S}_{3}=\mathcal{L}\cup\mathcal{C}_{3,\rm lin}\cup\mathcal{C}_{2,\rm lin}^{+}$, $\mathcal{S}_{4}=\mathcal{L}\cup\mathcal{C}_{3,\rm lin}\cup\mathcal{C}_{2,\rm lin}^{+}\cup\mathcal{C}_{1,\rm lin}^{+}$. Let $\mathcal{S}_{1}^{*},\ldots,\mathcal{S}_{4}^{*}$ denote the corresponding random sets constructed from $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}}$. Now:
Forward Tickwise Selection ($\rm FTBS$)
1:Test whether $H_{0}:\mathcal{S}=\mathcal{S}_{1}^{*}$ holds using Theorem 3.3. If the null is not rejected set $\widehat{\mathcal{S}}=\mathcal{S}_{1}^{*}$ and stop. Otherwise, set $a=2$.
2:Test whether $H_{0}:\mathcal{S}=\mathcal{S}_{a}^{*}$ holds using Theorem 3.3. If the null is not rejected set $\widehat{\mathcal{S}}=\mathcal{S}_{a}^{*}$ and stop. Otherwise,
3:Set $a=a+1$ and repeat Step 2.
Algorithm FTBS is computationally simple. It requires $\binom{K}{2}$ simple tests, see (25), followed by the construction of the linearly independent sets $\mathcal{S}_{1}^{*},\ldots,\mathcal{S}_{4}^{*}$ which is, computationally, a simple task. In addition, FTBS requires the fitting and consequent evaluation, by testing for lack of fit, of at most four linear models, see Theorem 3.3. The properties Algorithm $\rm FTBS$ are explored in Theorem 3.4 below.
Theorem 3.4
Fix $\alpha$ and let $\widehat{\mathcal{S}}$ denote the model selected by Algorithm $\rm FTBS$. If $n_{ij}\to\infty$, $1\leq i<j\leq K$} and $n_{ij}=O(n)$ then $\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))\to 1$. Moreover, if the errors are sub–Gaussian RVs, then for some $C_{1},C_{2}\geq 0$ we have
$\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))\geq 1-C_{1}e^{(-C_{2}n)}.$
Additionally, if: $(i)$ all tests within $\rm FTBS$ are performed at a level $\alpha_{n}\rightarrow 0$; (ii) the power of rejecting any of the false nulls is $\pi_{n}\rightarrow 1$ when $(i)$ holds; and if $(iii)$ triads in the minimal model share at most index. Then,
$\mathbb{P}(\mathcal{S}=\widehat{\mathcal{S}})\to 1.$ (26)
Theorem 3.4 shows that Algorithm FTBS outputs a model $\widehat{\mathcal{S}}$ which is a superset of the true model $\mathcal{S}$. Moreover, if the minimal is unique and all tests are applied with a small significance level, then in large samples the correct model will be selected with probability one. Theorem 3.4 is established by showing that for large enough samples $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}}\supset\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$ with high probability. Thus, the sets $\mathcal{S}_{a}^{*}\supset\mathcal{S}_{a}$, which implies that all false null hypotheses tested in Step 2 of FTBS will be rejected. When the errors are subguassian the probability of not selecting $\widehat{\mathcal{S}}\supset\mathcal{S}$ is exponentially small. Incidentally, the assumption of subguassian errors can be replaced with subexponential errors and possibly even further relaxed. In the third part of Theorem 3.4 it is assumed that triads in the minimal model share at most index. This implies a unique minimal model, a one–to–one correspondence between the support and the span and guarantees that
$\mathbb{P}(\mathcal{S}=\widehat{\mathcal{S}})\to 1.$ (27)
If the triads in the minimal model share more than one index then (27) can not be guaranteed. For example, suppose that the unique minimal model is $\mathcal{C}_{\boldsymbol{\gamma}_{1}}=\{\boldsymbol{c}_{(1,4,5)},\boldsymbol{c}_{(2,5,6)},\boldsymbol{c}_{(3,4,6)}\}$. The model $\mathcal{C}_{\boldsymbol{\gamma}_{2}}=\{\boldsymbol{c}_{(1,4,5)},\boldsymbol{c}_{(2,5,6)},\boldsymbol{c}_{(3,4,6)},\boldsymbol{c}_{(4,5,6)}\}$ is such that ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$, but ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\neq{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. It can be verified that all triads in $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$ have $3$-ticks and therefore it is clear that FTBS will select $\boldsymbol{\gamma}_{2}$ and whence $\mathcal{S}\neq\widehat{\mathcal{S}}$.
Remark 3.2
An analogue of Theorem 3.4 applies also to FSTS except for the relation expressed in (26).
4 Simulations
In this section we compare four methods of model selection; two general purpose methods, namely forward stepwise regression (FSR) and the LASSO, see Hastie et al. (2009), and two specialized methods designed for the problem at hand, namely forward tickwise selection (FTBS) and step wise triad based selection (FSTS) described in Section 3.3. LASSO is implemented using cross validation, see the R-package glmnet. All other methods use multi–step testing based methods, in which all tests are carried out at the $0.05$ level. In FTBS, a Bonferroni corrected FWER is used for selecting the support.
Three experimental scenarios are considered, see Table 1. In Scenario I, the minimal model is unique, moreover the triads in the minimal model have three ticks whereas all other triads have either 0 or $1$ ticks in $\mathcal{C}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$. In Scenario II the minimal model is again unique, all triads in the minimal model have three ticks as does the triad $(4,5,6)$ which is not present in the minimal model. Finally, in Scenario III the minimal model is not unique as $\boldsymbol{c}_{(1,2,3)}-\boldsymbol{c}_{(1,2,4)}=-\boldsymbol{c}_{(1,3,4)}+\boldsymbol{c}_{(2,3,4)}$, and thus no triad has three ticks.
Table 1: Simulation scenarios
Scenario $K$ $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$
I 6 $\boldsymbol{c}_{(1,2,3)}-\boldsymbol{c}_{(1,4,5)}$
II 6 $\boldsymbol{c}_{(1,4,5)}-\boldsymbol{c}_{(2,5,6)}+\boldsymbol{c}_{(3,4,6)}$
III 6 $\boldsymbol{c}_{(1,2,3)}-\boldsymbol{c}_{(1,2,4)}$
[…]
Table 2: Outcomes of Scenario I and II with $K=6$
$m=5$ $m=10$ $m=20$
[…]
Scenario I FSR 0.842 1.30 0.987 1.33 1 1.31
LASSO 0.976 3.20 1 3.21 1 3.21
FTBS 0.312 1.33 0.853 1.52 1 1.06
FSTS 0.209 0.66 0.764 0.97 0.993 1.03
Scenario II FSR 0.738 1.17 0.973 1.17 0.998 1.18
LASSO 0.971 2.77 1 2.63 1 2.55
FTBS 0.191 1.70 0.673 1.82 0.999 1.38
FSTS 0.066 0.71 0.568 1.01 0.986 1.08
Table 3: Outcomes Scenario II with $K=10$
$m=5$ $m=10$ $m=20$
[…]
FSR 0.607 1.97 0.962 1.96 0.999 1.90
LASSO 0.855 4.42 0.997 4.63 1 4.43
FTBS 0.086 2.63 0.374 2.40 0.942 1.55
FSTS 0.006 0.49 0.202 0.81 0.896 1.03
Table 4: Outcomes Scenario III with $K=6$
$m=5$ $m=10$ $m=20$
$)\widehat{\mathbb{P}}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}$)) $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$ $)\widehat{\mathbb{P}}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}$)) $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$ $)\widehat{\mathbb{P}}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}$)) $\widehat{\mathbb{E}}(|\widehat{\mathcal{S}}|/|{\mathcal{S}}|)$
FSR 0.743 1.24 0.982 1.27 1 1.26
LASSO 0.784 3.99 0.987 4.44 1 4.54
FTBS 0.228 1.16 0.718 1.55 0.993 1.57
FSTS 0.022 0.61 0.29 1.08 0.926 1.53
[…]
Remark 4.1
In some settings FTBS can be used as a screening tool to which one can apply either FSR or FSTS. Concretely, one may use the tick–table to identify potential triads to be included in the selected model and then apply either FSR or FSTS to this reduced collection of triads. This approach is exemplified in the following section.
Finally, we examine the mean squared errors (MSEs) incurred when estimating the preference profile $\operatorname{\boldsymbol{\nu}}$ in the three scenarios presented in Table 1. We do so in four settings: $(i)$ when the true model is known in advance; $(ii)$ when the true model is unknown but selected using FTBS; $(iii)$ when the true model is unknown and the full model is fit; and $(iv)$ when the true model is unknown and the reduced fully transitive model is fit.
Table 5: Mean squared errors for the LSE of $\operatorname{\boldsymbol{\nu}}$ under Scenarios I–III in the settings described above and $m=20$.
$m=20$
Model Scenario I Scenario II Scenario III
True 0.354 0.405 0.352
Selected 0.369 0.483 0.436
Full 0.745 0.745 0.745
Reduced 6.252 9.252 4.252
As expected, the MSE is minimized when the true model, as specified in Table 1, is known. When the true model is unknown but selected using FTBS, the increase in MSE is modest (on the order of $20\%$). In contrast, a substantial increase in MSE (approximately $100\%$) is observed when fitting the full model (i.e., $\widehat{\operatorname{\boldsymbol{\nu}}}=\bar{\operatorname{\boldsymbol{S}}}_{\rm ALL}$). Finally, when the reduced linearly transitive model is fitted, the MSE increases by nearly $1000\%$; this is because, under the LST assumption, consistent estimation of $\operatorname{\boldsymbol{\nu}}$ is not possible. These results demonstrate that the proposed methodology substantially improves estimation of the underlying preference profile. Consequently, improvements in ranking and prediction are to be expected. An application to betting and wagering is presented in the next section.
5 Illustrative example
Sporting competitions are an important source of interest, study and entertainment for the general public, as well as an active arena in which statistical analyses and ranking methods are passionately applied. To date, methods employed for ranking sports teams have assumed linear transitive models (Sinuany-Stern 1988, Cassady et al. 2005, Langville and Meyer 2012, Barrow et al. 2013). Alternatives allowing for cyclicalities and intransitivities were simply not available.
We apply the framework developed in this communication to analyze data from the English Premier League (EPL), one of the foremost soccer/football leagues in the world. We focus on the 2022–23 season in which $20$ teams competed. The EPL operates on a double round–robin format, i.e., each team plays every other team twice, once on each home field. Thus, $38$ matches are played by each team resulting in a total of $380$ matches. In the following, the outcome of each game is defined as the difference in expected goals (xG) scored; see Roccetti et al. (2024) for a discussion of xG in evaluating soccer performance. The second and third columns of Table 6 report the names of the teams and their positions, that is, their official ranking as it appears on the EPL website www.premierleague.com. Teams are alphabetically listed, and the first column of Table 6 assigns each team an index number by which we shall refer to them.
[…]
Index Name Position $\widehat{\mu}_{s,i}$ $r(\widehat{\mu}_{s,i})$ $\mu_{i}^{**}$ $r(\mu_{i}^{**})$
1 Arsenal 2 0.7325 3 17 2
2 Aston Villa 7 -0.0575 9 10 11
3 Bournemouth 15 -0.635 20 4 19
4 Brentford 9 0.165 8 13 6
5 Brighton 6 0.585 4 15 5
6 Chelsea 12 -0.07 10 9 12
7 Crystal Palace 11 -0.2225 12 6 14
8 Everton 17 -0.505 16 5 15
9 Fulham 10 -0.4325 14 7 13
10 Leeds United 19 -0.49 15 4 17
11 Leicester City 18 -0.3225 13 10 8.5
12 Liverpool 5 0.54 5 16 3
13 Manchester City 1 1.16 1 18 1
14 Manchester Utd 3 0.43 6 12 7
15 Newcastle Utd 4 0.805 2 15 4
16 Nott’ham Forest 16 -0.6175 19 4 18
17 Southampton 20 -0.5825 17 5 16
18 Tottenham 8 0.195 7 10 8.5
19 West Ham 14 -0.095 11 10 10
20 Wolves 13 -0.5825 18 3 20
First, ignoring possible cyclicalities, the data is analyzed assuming the reduced model, i.e., assuming the relation (2) holds. The fourth column of Table 6 reports the estimated merits and the fifth column the merit–based rankings where the rank of the $i^{th}$ team is given by $r(\widehat{\mu}_{i})=\sum_{j\neq i}\mathbb{I}(\widehat{\mu}_{i}>\widehat{\mu}_{j})$. For example, Manchester City is ranked first with a merit of $2.32$. In addition, note that merit–based ranking coincides with both the row–sum ranking method and the method proposed by Saari (2014) and Saari (2021), see Equation (14).
Next, we apply a test for lack of fit, to assess whether the reduced model fits the data. We find that the p-value associated with this test is $10^{-3}$ indicating a strong deviation from the null. A test for cyclicality, see Proposition 3.1, is applied to each of the $1140$ triads. Under the null, $57$ triads are expected to be associated with a p–value smaller than $0.05$; in fact, there are $195$ such triads, providing additional evidence for a departure from LST. Interestingly, the most significant triad with a p-value of $10^{-4}$ is $(2,5,19)$ for which $(\widehat{\nu}_{2,5},\widehat{\nu}_{2,19},\widehat{\nu}_{5,19})=(1.15,-0.35,2.05)$ so $2\succ 5\succ 19\succ 2$.
Moving on to model selection, we first directly apply the methods compared in Section 4. We find that: (a) LASSO, which uses cross–validation, does not select any triads; (b) FSR (with $\alpha=0.05$) selects a model with $16$ triads; (c) FSTS selects the model with the two most significant triads $(2,5,19)$ and $(6,11,12)$; $(d)$ Applying FTBS with a Bonferonni–corrected family wise error rate of $0.1$ we find that the only element of $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ which is different from 0 is $\nu_{{\rm cyclic},2,5}$. Consequently the tick–table $\operatorname{\boldsymbol{C}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}}$ has no $3$ or 22–tick columns and $18$ triads have 11–tick. Since FTBS selects variables in blocks, a naive application of FTBS will select a model with $18$ triads. If the family wise error rate is reduced to $0.05$ then no triads would be chosen.
Clearly, FTBS does not operate satisfactorily in this example. This is likely due to the fact that $m=2$ so the tests in (25) are under–powered and therefore do not detect some pairs in the support of the true minimal model. The correct estimation of the support would likely have led to the detection of true three-tick triads. The performance of FTBS can be modified and enhanced in various ways. In essence, to reduce the expected model size of FTBS, the number of column blocks added at each stage should be minimized. One approach is to use FTBS as a screening tool to identify potential triads for further selection. Methods that select one triad at a time can then be applied to this subset, making FTBS–FSTS or FTBS–FSR hybrids possible.
When applied to the EPL2023 data set FTBS–FSR selects a model with a single triad $(2,5,19)$, whereas the FTBS–FSTS selects a model with two triads $(2,5,19),(6,11,12)$. FTBS–FSR is preferred by BIC whereas FTBS–FSTS is preferred by AIC although the differences are rather small. For simplicity of exposition, we shall refer to the first of these models as the selected (or final) model; it is given by
${\widehat{\operatorname{\boldsymbol{\nu}}}}_{s}=\operatorname{\boldsymbol{B}}{\widehat{\operatorname{\boldsymbol{\mu}}}}_{s}+\boldsymbol{c}_{(2,5,19)}{\widehat{{\gamma}}}_{s}$
where $\widehat{{\operatorname{\boldsymbol{\mu}}}}_{s}$ and $\widehat{\gamma}_{s}$ are estimated as described in Theorem 3.2. The merits associated with the selected model are displayed in column four of Table 6 and ${\widehat{{\gamma}}}_{s}=1.18$. Further note that in the full model $(\widehat{\nu}_{2,5},\widehat{\nu}_{2,19},\widehat{\nu}_{5,19})=(1.15,-0.35,2.05)$ and in the selected model $({\widehat{\nu}}_{s,2,5},{\widehat{\nu}}_{s,2,19},{\widehat{\nu}}_{s,5,19})=(0.54,-1.14,1.86)$. Thus, the preference profile is non transitive even after incorporating the triad $(2,5,19)$ in the model. It is also interesting that a cyclical triad may include teams in very different positions. For example, the triad $(2,5,19)$ consists of Aston Villa, Brighton, and West Ham located in positions $6$, 77 and $14$. The cycle among them induces a complicated web of preferences.
A complete ranking in the presence of true cyclicalities is challenging. Columns six and seven in Table 6 display the dominance scores and ranks for all teams in the EPL, respectively. The dominance score of teams $\{2,11,18,19\}$ are initially equal, so a tie braking procedure, see Section 2.1, is used. However, the resulting ranking cannot be easily accepted. Figure 1 depicts the preference profile among the aforementioned teams. An arrow from $i$ pointing at $j$ means that $i\succ j$. The relation between teams $11$ and $18$ is of equality. Figure 1 shows two cycles, i.e., $2\succ 18\sim 11\succ 2$ and $2\succ 18\succ 19\succ 2$. Any global method of ranking will not honor such cycles and is therefore open to challenge. One way of addressing and partially resolving these issues is to define a ranking with respect to each item $i$. Define the sets: (i) $S_{i}=\{j:\widehat{\nu}_{ij}<0\}$; (ii) $I_{i}=\{j:\widehat{\nu}_{ij}>0\}$; and (iii) $E_{i}=\{j:\widehat{\nu}_{ij}=0\}$. The elements in $S_{i}$ and $I_{i}$ are ranked above (superior) and below (inferior) item $i$, respectively, and the elements in $E_{i}$ are equivalent to $i$. See Table 7 in the Supplementary Materials for the corresponding sets.
22111118181919
Figure 1: Preferences among teams with indices $\{2,11,18,19\}$
Finally, we explore implications for wagers and bets. It is well known that if one believes that $\mathbb{P}(W)=\omega$ for some event $W$, then one should be willing to accept a bet of size $({1-\omega})/{\omega}$ against a bet of one monetary unit (see, e.g., Epstein, 2012). Suppose now that an adversary believes that $\mathbb{P}(W)=\omega$, but in truth $\mathbb{P}(W)=\tau$, and for the moment let $\tau>\omega$. Since our adversary is willing to bet $(1-\omega)/\omega$ for one monetary unit, our expected gain is $\tau({1-\omega})/{\omega}+(1-\tau)(-1)=({\tau-\omega})/{\omega}$. Since $\tau>\omega$, our net gain is positive. Similarly, if $\tau<\omega$, we should instead bet against $W$ (i.e., for $W^{c}$). Repeating the calculation, our net gain is $({\omega-\tau})/({1-\omega})>0$.
Next let $W_{ij}$ denote the event that team $i$ wins against team $j$ in their next encounter. Assuming that $Y_{ij}$ is a $\mathcal{N}(\nu_{ij},\sigma^{2})$ RV we find that
$\mathbb{P}(W_{ij})=1-\Phi\left(\frac{\nu_{ij}}{\sigma}\right),$
where $\Phi$ is the standard normal cumulative distribution function. Let $\omega_{ij}$ denote the probability of $W_{ij}$ assuming LST, i.e., model (3), and let $\tau_{ij}$ denote the probability of $W_{ij}$ under a possibly intransitive model, see (11). Thus, $\omega_{ij}=1-\Phi(\nu_{ij,\mathrm{LST}}/{\sigma}),$ where $\nu_{ij,\mathrm{LST}}$ is the $(i,j)^{th}$ entry of $\operatorname{\boldsymbol{\nu}}_{\mathrm{LST}}$, as given in Theorem B.1. Most gambling systems (e.g., Langville and Meyer, 2012) incorrectly assume that $\operatorname{\boldsymbol{\nu}}_{\mathrm{LST}}$ represents the true value of $\operatorname{\boldsymbol{\nu}}$. In reality, the true probability is $\tau_{ij}=1-\Phi(\nu_{ij,\mathrm{TRU}}/{\sigma}),$ where $\operatorname{\boldsymbol{\nu}}_{\mathrm{TRU}}$ is the true underlying parameter vector.
The expected gain from betting one monetary unit for, or against, the event $W_{ij}$ is then given by
$\mathrm{Win}_{ij}=\begin{cases}\displaystyle\frac{\tau_{ij}-\omega_{ij}}{\omega_{ij}},&\text{if }\tau_{ij}>\omega_{ij},\\[8.61108pt] \displaystyle\frac{\omega_{ij}-\tau_{ij}}{1-\omega_{ij}},&\text{if }\tau_{ij}<\omega_{ij},\\[8.61108pt] 0,&\text{if }\tau_{ij}=\omega_{ij}.\end{cases}$
Accordingly, the total expected gain over a round–robin tournament (where each pair $(i,j)$ meets once) is
$\mathrm{TotalWin}=\sum_{1\leq i<j\leq K}\left[\frac{\tau_{ij}-\omega_{ij}}{\omega_{ij}}\,\mathbb{I}(\tau_{ij}>\omega_{ij})+\frac{\omega_{ij}-\tau_{ij}}{1-\omega_{ij}}\,\mathbb{I}(\tau_{ij}<\omega_{ij})\right].$
This expression quantifies the cost of belief miscalibration due to model misspecification, i.e., assuming transitivity when cyclicality may be present. As an illustration, we estimate $\mathrm{Win}_{ij}$ and $\mathrm{TotalWin}$ using EPL2023 data. Specifically, we set $\nu_{ij,\mathrm{LST}}=\widehat{\mu}_{i}-\widehat{\mu}_{j}$, $\nu_{ij,\mathrm{TRU}}=\widehat{\nu}_{ij,s}$, and $\sigma=\widehat{\sigma}$ as given in Proposition 3.1. Notably, $\mathrm{Win}_{ij}>0$ only if $(i,j)\in\operatorname{support}(\mathcal{C}_{\gamma_{s}})$; hence in the example at hand, $\mathrm{Win}_{ij}>0$ whenever $\{i,j\}\in\{2,5,19\}$, and zero otherwise. In particular, $\mathrm{Win}_{2,5}=2.67,\quad\mathrm{Win}_{2,19}=0.93,\quad\mathrm{Win}_{5,19}=0.23$, so that
$\mathrm{TotalWin}=3.83.$
In summary, modeling cyclicality in paired comparison data can have a considerable impact on betting and wagering systems.
6 Discussion
This paper presents a principled and unified approach to the modeling and analysis of PCD in the presence of cyclicality and intransitivity. The key is the decomposition of the parameter space $\mathcal{N}$ into a linear transitive subspace and cyclical subspace and expressing $\operatorname{\boldsymbol{\nu}}$ as (11). This form captures all possible preferences relations and allows for a compact representation in situations in which cyclicalities are limited to a subset of the items, that is, when the minimal model is of a relatively small size. It is worth noting that an analysis of variance (ANOVA) type decomposition of the form
$\nu_{ij}=\mu_{i}-\mu_{j}+\gamma_{ij}$
where $\sum_{j}\gamma_{ij}=0$ and $\sum_{i}\gamma_{ij}=0$ for all $i$ and $j$ has been proposed by Scheffé (1952) and studied by several other authors. Although the ANOVA decomposition yields a saturated model, that is, any $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ can be consistently estimated, it does not exploit the intrinsic structure of PCD and lacks the algebraic interpretation offered by the current model. In particular, Scheffé’s merits and interaction parameters do not correspond to linear transitivity and cyclicality. Saari (2014), expanded on in Saari (2021), was the first to obtain bases for $\mathcal{L}$ and $\mathcal{C}$. However, his objective was to obtain a global ranking rather than modeling the preference relationship $\operatorname{\boldsymbol{\nu}}$. Moreover, he did not attempt a decomposition of $\operatorname{\boldsymbol{\nu}}$ nor considered the notions of a minimal model or the tick table $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}}$. Most importantly, Saari (2014) and Saari (2021) did not develop any tools for conducting statistical inference for such models.
This paper starts with a rigorous study of the parameter space $\mathcal{N}$ and then shows that in scenarios likely in practice PCD can be satisfactorily analyzed, that is, the sources of cyclicalities and intransitivities can be fully identified, estimated, and addressed within the framework of minimal models. Thus, more accurate estimators of the preference profile are possible yielding improved rankings, predictions and bets. These aspects are well demonstrated by simulation and the illustrative example. We propose a new tailor–made test–based model selection method, which we refer to as FTBS. Simulations indicate that the proposed method, which is computationally efficient, performs well in practice, especially for larger graphs. Our illustrative example, which is of independent interest, shows that when ranking EPL teams the linear transitive model does not fit. In the selected model, some triads exhibit cyclicality, a fact that complicates ranking. We are eager to apply the proposed methodology to a wide variety of other sports, chess, and consumer preferences and to a variety of relevant data sets explored by other researchers. We believe that such studies will be of great interest. The example also demonstrates that model selection may be a challenge in situations where the number of comparisons between two items is small. This outcome, while unsurprising, indicates that some adaptation of existing methods will be required.
Many open problems, both theoretical and applied, have not been explored. Some are listed here. To start with, our simulation study has not investigated many interesting and important issues. For example, we have not explored very large comparison graphs with either sparse or dense degree of cyclicality. Nor have we attempted to optimize FTBS. Recall that FTBS is a block–wise method, i.e., it selects blocks of columns rather than a single columns at a time as is done by FSR and FSTS. It is clear that by reducing the block size by partitioning existing block into sub–blocks (using some criteria such as their level of significance of their individual components), one could reduce the expected size of the model, i.e. $\mathbb{E}(|\widehat{\mathcal{S}}|/|\mathcal{S}|)$ without compromising on the probability of correct selection. One approach to this problem is to combine FTBS with methods such as FSR or FSTS as demonstrated in Section 5. Applying model selection criteria such as AIC or BIC (Clasken and Hojrt, 2008) can further improve the selected model. Furthermore, we believe that by optimizing the significance level at which tests are performed at various stages one may be able to attain additional benefits. In addition, we have not addressed the issue of long cycles of the form $i_{1}\succeq i_{2}\succeq\cdots\succeq i_{r}\succeq i_{1}$ for $r\geq 3$. In general, such cycles can be expressed by the vector $\boldsymbol{c}_{\boldsymbol{\lambda}}\in\mathcal{C}$ where
$c_{ij}({\boldsymbol{\lambda}})=\mathbb{I}((i\succ j)\in\boldsymbol{\lambda})-\mathbb{I}((j\succ i)\in\boldsymbol{\lambda}),$
and $\boldsymbol{\lambda}=(i_{1},i_{2},\ldots,i_{r})$. Note that $\boldsymbol{c}_{\boldsymbol{\lambda}}$ can be expressed as a linear combination of vectors in $\{\boldsymbol{c}_{(i,j,k)}\}_{1\leq i<j<k\leq K}$. For example, the longest possible cycle, i.e., $1\succ 2\succ\ldots\succ K\succ 1$, associated with the vector $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,\ldots,K)}$ can be expressed the sum of K2K-2 linearly independent triads, e.g., $\boldsymbol{c}_{(1,\ldots,K)}=\sum_{j=2}^{K-1}\boldsymbol{c}_{(1,j,j+1)}$. Of course, there are many other ways to write $\boldsymbol{c}_{(1,\ldots,K)}$. Thus, cycles are not associated with a unique minimal model. Nevertheless, it is clear that long cycles can be identified from the data and incorporated into our modeling framework. In this paper, we focus on complete PCGs. However, in practice, PCGs may be incomplete, i.e., not all paired comparisons are present. For clarity and brevity, we omit a detailed analysis of such cases. Another important area that requires further research is the relationship between the proposed framework and the characterization of transitivity relations. For example, can tests be developed for or against weak and/or strong transitivity? Can estimators of the model parameters that respect such relations be constructed, and what properties might these estimators possess?
Finally, although this paper focuses on cardinal PCD it can be extended to binary and other types of paired comparison data. For example, in binary PCD the probability that item $i$ is preferred over item $j$ is typically modeled as $\mathbb{P}(Y_{ijk}=1)=F(\mu_{i}-\mu_{j})$ where $F$ is a distribution function, symmetric about 0, i.e., F(x)+F(x)=1F(x)+F(-x)=1 for all $x\in\mathbb{R}$, and $\mu_{1},\ldots,\mu_{K}$ are the merit parameters, cf. Oliveira et al. (2018). Specifically, when $F$ is the standard logistic distribution we obtain the relation
$\mathbb{P}(Y_{ijk}=1)=\frac{\exp(\mu_{i}-\mu_{j})}{1+\exp(\mu_{i}-\mu_{j})}.$
This is the well–known Bradley–Terry model (BT) (Bradley and Terry 1952, Hunter 2004, Cattelan 2012) whereas when $F$ is the standard normal DF we obtain the so–called Thurstone model (Thurstone 1927, Böckenholt, 2006). Clearly, the classic BT model can be extended. Modelling cyclicalities is straightforward; instead of $\mu_{i}-\mu_{j}$ write $\nu_{ij}$ and conduct inference on $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$. A bit of reflection shows that the modeling framework for cardinal PCD, discussed in Section 2, is applicable. The estimators and corresponding tests will, of course, be different; a detailed analysis is clearly warranted.
Acknowledgments
The work of Rahul Singh was conducted while a post–doctoral fellow at the University of Haifa. The work of Ori Davidov was partially supported by the Israeli Science Foundation Grant No. 2200/22 and gratefully acknowledged.
Appendices
We provide three appendices. Appendix A contains the proofs of all theorems and propositions presented in the main text. Appendix B offers additional results on the relationship between the reduced and full models. Appendix C presents some numerical results related to the illustrative example.
Appendix A Proofs
Appendix A contains the proofs of all theorems and propositions appearing in the main text. Some supporting auxiliary remarks and lemmas are also provided.
Proof of Theorem 2.1:
Proof. By Definition if 2.1 $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ and consequently $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\cap{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Next suppose that ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. It immediately follows that $\mathcal{C}_{\boldsymbol{\gamma}_{1}}$ and $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$ have same set of items, say $\{1,\ldots,L\}$. For if $i$ is present in $\mathcal{C}_{\boldsymbol{\gamma}_{1}}$ and absent in $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$ then there exist a triad $\boldsymbol{c}_{(i,j,k)}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$ and $\boldsymbol{c}_{(i,j,k)}\notin{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ for all $(j,k)\notin{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Hence ${\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\neq{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Next, suppose that $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$. Then there exists a $k\leq L$ such that $\boldsymbol{c}_{(i,j,k)}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$ and consequently $\boldsymbol{c}_{(i,j,k)}\in{\rm span}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Therefore if the spans are equal so are the supports.
Let $\boldsymbol{\gamma}_{1}$ and $\boldsymbol{\gamma}_{2}$ be two minimal models. By assumption $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$ implies $\nu_{{\rm cyclic},ij}\neq 0$, therefore it also follows that $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ and consequently ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\subseteq{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$. Now if ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})\subset{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$, there exists some $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ and $(i,j)\notin{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})$. Thus there is some triad $(i,j,t)$ which belongs to $\mathcal{C}_{\boldsymbol{\gamma}_{2}}$ but not $\mathcal{C}_{\boldsymbol{\gamma}_{1}}$ consequently $s(\boldsymbol{\gamma}_{2})>s(\boldsymbol{\gamma}_{1})$ and therefore $\boldsymbol{\gamma}_{2}$ can not be a minimal model. Thus ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{1}})={\rm support}(\mathcal{C}_{\boldsymbol{\gamma}_{2}})$ and completing the proof of third part.
Finally suppose that any two triads in a minimal model $\boldsymbol{\gamma}$ spanned by $\mathcal{C}_{\boldsymbol{\gamma}}$ share at most one index, i.e., if $\boldsymbol{c}_{(i,j,k)},\boldsymbol{c}_{(u,v,w)}\in\mathcal{C}_{\boldsymbol{\gamma}}$ then $|\{i,j,k\}\cap\{u,v,w\}|\leq 1$. Using (8) it is easily verified that such triads are associated with orthogonal columns of $\operatorname{\boldsymbol{C}}$. It is also immediate that if $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ then $(i,j)$ belongs to a single triad in $\mathcal{C}_{\boldsymbol{\gamma}}$. By definition
$\operatorname{\boldsymbol{\nu}}_{{\rm cyclic}}=\sum_{\boldsymbol{c}_{(i,j,k)}\in\mathcal{C}_{\boldsymbol{\gamma}}}\boldsymbol{c}_{(i,j,k)}\boldsymbol{\gamma}_{ijk}.$ (28)
It is clear that if $(s,t)\notin{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ then $\nu_{{\rm cyclic},st}=0$, whereas if $(s,t)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ then $(s,t)$ belongs to a single triad $(i,j,k)$, say, and $\nu_{{\rm cyclic},st}={c}_{(i,j,k),st}{\gamma}_{ijk}\neq 0$. We conclude that $\nu_{{\rm cyclic},ij}\neq 0$ if and only if $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$. Also $\sum_{1\leq i<j\leq K}\mathbb{I}(\nu_{{\rm cyclic},ij}\neq 0)=3s$, where $s$ is size of the model. Next suppose there exist a minimal model $\boldsymbol{\gamma}^{\prime}$ for which $\boldsymbol{\gamma}\neq\boldsymbol{\gamma}^{\prime}$.
It follows from the discussion above that if $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ij}\neq 0$ then $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}^{\prime}})$. These pairs, however, are in ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ and therefore ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})\subseteq{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}^{\prime}})$. Remark A.1 implies that $s({\boldsymbol{\gamma}})<s({\boldsymbol{\gamma}^{\prime}})$, i.e., the size of the model $\boldsymbol{\gamma}^{\prime}$ is larger than the size of the model $\boldsymbol{\gamma}$. This leads to a contradiction and the conclusion that $\boldsymbol{\gamma}=\boldsymbol{\gamma}^{\prime}$.
Remark A.1
If a minimal model has to be constructed using a given support, then we seek a minimum number of triads using the support, such that these triads are linearly independent. That is, if $(i,j),(i,k),(j,k)$ are present in the support then $\boldsymbol{c}_{(i,j,k)}$ is a viable triad and we find such independent triads until support is covered. Now if triads in a minimal model $\boldsymbol{\gamma}$ share at most one index then any comparison appears in only one triad of $\mathcal{C}_{\boldsymbol{\gamma}}$. Precisely there is a one-to-one mapping between comparisons and triads in $\boldsymbol{\gamma}$. Consequently only triads in $\mathcal{C}_{\boldsymbol{\gamma}}$ can be made using ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$, therefore the (minimum) number of triads made using ${\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ is $s({\boldsymbol{\gamma}})$.
Proof of Theorem 2.2:
Proof. Suppose that a minimal model $\boldsymbol{\gamma}$ has a single one 0-tick triad $(i,j,k)$. It follows that all other triads in $\mathcal{C}_{\boldsymbol{\gamma}}$ have at least one tick and that $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ij}=\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ik}=\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},jk}=0$. Now by (28) for any pair $(s,t)$
$\nu_{{\rm cyclic},st}=\sum_{(u,v,w)\in\mathcal{C}_{\boldsymbol{\gamma}}}{c}_{(u,v,w),st}{\gamma}_{uvw}.$ (29)
Therefore if $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ij}=0$ then there must be some other triad in which the comparison $(i,j)$ appears. Similarly for the pairs $(j,k)$ and $(i,k)$. Also notice that any two distinct triads can have at most one comparison in common and therefore the minimal model will have at least three more triads with two indices in common with $(i,j,k)$ and having at least 11–tick. Assume that there are exactly three such triads. These must be $\boldsymbol{c}_{(i,j,s)},\boldsymbol{c}_{(j,k,t)},\boldsymbol{c}_{(i,k,u)}$ for some $s,t$ and $u$. By (29), $\nu_{{\rm cyclic},ij}={c}_{(i,j,k),ij}{\gamma}_{ijk}+{c}_{(i,j,s),ij}{\gamma}_{ijs}=0$ implying that ${\gamma}_{ijk}=-{\gamma}_{ijs}$. A similarly relation holds for $\nu_{{\rm cyclic},jk}=0$ and $\nu_{{\rm cyclic},ik}=0$ and therefore ${\gamma}_{ijk}=-{\gamma}_{jkt}$ and ${\gamma}_{ijk}={\gamma}_{iku}$. Combining these equalities we find that ${\gamma}_{ijk}=-{\gamma}_{ijs}=-{\gamma}_{jkt}={\gamma}_{iku}$. It is now clear that the elements of the vector $\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s)}-\boldsymbol{c}_{(j,k,t)}+\boldsymbol{c}_{(i,k,u)}$ which correspond to comparisons among the items $\{i,j,k\}$ are 0. This implies by (29) that
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}+\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(-ijk)},$
where $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}={\gamma}_{ijk}(\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s)}-\boldsymbol{c}_{(j,k,t)}+\boldsymbol{c}_{(i,k,u)})$ is the linear combination of the triads involving comparisons of the items in $\{i,j,k\}$ and $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(-ijk)}=\operatorname{\boldsymbol{\nu}}_{\rm cyclic}-\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}$.
Since all other triads in the minimal model have at least one tick we assume without any loss of generality that $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},js},\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},kt}\neq 0$. Using (9) we find
$\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s)}=\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s)}+\boldsymbol{c}_{(i,k,t)}-\boldsymbol{c}_{(i,k,t)}=\boldsymbol{c}_{(j,k,s)}-\boldsymbol{c}_{(i,k,t)}.$
Thus by using the above relation, we can rewrite $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}$ as
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}={\gamma}_{ijk}(\boldsymbol{c}_{(j,k,s)}-\boldsymbol{c}_{(i,k,t)}-\boldsymbol{c}_{(j,k,t)}+\boldsymbol{c}_{(i,k,u)}).$
Hence we obtain another minimal model, say $\boldsymbol{\gamma}^{\prime}$, in which there are no triads with 0 ticks. If there are more than three triads with two indices in common with $\{i,j,k\}$ then we can write $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}=\alpha_{1}(\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s_{1})}-\boldsymbol{c}_{(j,k,t_{1})}+\boldsymbol{c}_{(i,k,u_{1})})+\ldots+\alpha_{l}(\boldsymbol{c}_{(i,j,k)}-\boldsymbol{c}_{(i,j,s_{l})}-\boldsymbol{c}_{(j,k,t_{l})}+\boldsymbol{c}_{(i,k,u_{l})})$. Applying the preceding arguments to each term appearing in $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{(+ijk)}$ separately we arrive at the same conclusion, i.e., there exist an alternative minimal model with no 0–tick triads. Finally, if there are $l>1$ 0–tick triads in a minimal model then we repeat the procedure outline above for each one of them. This completes the proof of the first part of Theorem 2.2.
Suppose that any two triads in a minimal model $\boldsymbol{\gamma}$ share at most one index. The proof of Theorem 2.1 shows that if $(i,j)\in{\rm support}(\mathcal{C}_{\boldsymbol{\gamma}})$ then $\operatorname{\boldsymbol{\nu}}_{{\rm cyclic},ij}\neq 0$. This immediately implies that all triads in $\boldsymbol{\gamma}$ have 33–ticks. Next it is shown, by example, that a triad $(i,j,k)$ may have $3$-ticks yet not be a member of any minimal model. Let
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\lambda_{1}\boldsymbol{c}_{(i,j,s)}+\lambda_{2}\boldsymbol{c}_{(j,k,t)}+\lambda_{3}\boldsymbol{c}_{(i,k,u)},$
where $|\lambda_{l}|$’s are distinct. It is not hard to see that the minimal model is unique. Also note that the triads (i,j,s),(j,k,t)(i,j,s),~(j,k,t) and $(i,k,u)$ all have $3$ ticks. However so does the triad $(i,j,k)$. This completes the proof of the second part of the theorem.
Finally consider $\alpha\in\mathbb{R}\setminus\{0\}$ and set
$\operatorname{\boldsymbol{\nu}}_{\rm cyclic}=\boldsymbol{c}_{(1,2,3)}-\alpha\boldsymbol{c}_{(1,2,4)}-(1-\alpha)\boldsymbol{c}_{(1,2,5)}.$
It is easy to verify that the minimal model for $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ above is unique. It is straightforward to verify that all triads in the minimal model have two ticks. This completes the proof.
Proof of Theorem 2.3:
Proof. Pick a triplet $(1,2,3)$, say. Recall, see Example 2.2, that when $K=3$ the relevant means satisfy
$\begin{pmatrix}\nu_{12}\\ \nu_{13}\\ \nu_{23}\end{pmatrix}=\begin{pmatrix}\mu_{1}-\mu_{2}+\gamma\\ \mu_{1}-\mu_{3}-\gamma\\ \mu_{2}-\mu_{3}+\gamma\end{pmatrix}$,
for some $\gamma$. Suppose that $\operatorname{\boldsymbol{\nu}}$ is strongly transitive and without loss of generality assume that $1\succeq 2\succeq 3$. If so
$\nu_{12},\nu_{23}\geq 0\text{ and }\nu_{13}\geq\max\{\nu_{12},\nu_{23}\}.$
If $\gamma<0$, then $\nu_{12}\geq 0$ implies that $\mu_{1}-\mu_{2}>-\gamma>0$ so $\mu_{1}>\mu_{2}$. Similarly $\nu_{23}\geq 0$ implies $\mu_{2}>\mu_{3}$. Together these imply that $\mu_{1}\geq\mu_{2}\geq\mu_{3}$. Next if $\gamma>0$ then strong transitivity implies that $\mu_{1}-\mu_{3}-\gamma\geq\mu_{1}-\mu_{2}+\gamma$, so $\mu_{2}-\mu_{3}>2\gamma>0$. Similarly $\mu_{1}-\mu_{3}-\gamma\geq\mu_{2}-\mu_{3}+\gamma$ implies $\mu_{1}-\mu_{2}>2\gamma>0$. Combining these two relations we find that $\mu_{1}\geq\mu_{2}\geq\mu_{3}$ as required. Finally if $\gamma=0$ then we have $\mu_{1}-\mu_{2}\geq 0$ and $\mu_{2}-\mu_{3}\geq 0$ which immediately imply that $\mu_{1}\geq\mu_{2}\geq\mu_{3}$. Since this argument holds for any triplet we conclude that strong stochastic transitivity is equivalent to the ordering of the merits.
Next it will be shown that if $\operatorname{\boldsymbol{\nu}}$ is weakly transitive the merits may not be ordered. Under weak transitivity
$\nu_{12},\nu_{23}\geq 0\text{ implies that }\nu_{13}\geq 0.$
Set $\gamma>0$ and $\mu_{2}=\mu_{1}+\gamma/2$. If so $\nu_{12},\nu_{23},\nu_{13}\geq 0$ holds but $\mu_{1}<\mu_{2}$. We conclude that under weak transitivity the relations $1\succeq 2\succeq 3$ does not imply that $\mu_{1}\geq\mu_{2}\geq\mu_{3}$. This concludes the proof.
Proof of Proposition 2.1:
Proof. Without any loss of generality assume $\mu_{1}^{**}>\ldots>\mu_{K}^{**}$. Consequently: (i) $\nu_{1j}>0$ for $2\leq j\leq K$, so $1\succ j$ where $2\leq j\leq K$; (ii) for any $i>1$, $\nu_{ij}>0$ for $j>i$ and $\nu_{ij}<0$ for $j<i$. Therefore for any $i$ items $\{1,\ldots,i-1\}$ are preferred over $i$ and $i$ is preferred over items $\{i+1,\ldots,K\}$. Implying that the preference profile is weakly transitive with $1\succ\ldots\succ K$.
Proof of Theorem 3.1:
Proof. The proof follows the proof of Theorem 4.1 in Singh et al. (2025). The normal equation are given by
$\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{C}}\\ \operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{C}}\end{pmatrix}\begin{pmatrix}\operatorname{\boldsymbol{\mu}}\\ \boldsymbol{\gamma}\end{pmatrix}=\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}^{\top}\end{pmatrix}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{Y}}.$
Thus the minimum norm solution to the above system is given by
$\begin{pmatrix}\widehat{\operatorname{\boldsymbol{\mu}}}\\ \widehat{\boldsymbol{\gamma}}\end{pmatrix}=\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{C}}\\ \operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{D}}\operatorname{\boldsymbol{C}}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}^{\top}\end{pmatrix}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{Y}}.$
Explicit expressions for $\widehat{\operatorname{\boldsymbol{\mu}}}$ and $\widehat{\boldsymbol{\gamma}}$ can be obtained by noting that
$\operatorname{\boldsymbol{Y}}=~\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}+\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}\boldsymbol{\gamma}+\boldsymbol{\epsilon}=~{\operatorname{\boldsymbol{L}}}{\operatorname{\boldsymbol{B}}}\boldsymbol{c}_{1}+\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}\boldsymbol{c}_{2}+\boldsymbol{\epsilon},$
where $\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}=(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}$ and
$\boldsymbol{c}_{2}$ $~=\boldsymbol{\gamma},$
$\boldsymbol{c}_{1}$ $~=\operatorname{\boldsymbol{\mu}}-(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}\boldsymbol{\gamma}.$
Note that $\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}$ is projection of $\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}}$ onto the orthogonal complement of $\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}$, and consequently in terms of $(\boldsymbol{c}_{1},\boldsymbol{c}_{2})$ the normal equations are
$\begin{pmatrix}\operatorname{\boldsymbol{N}}&\boldsymbol{0}\\ \boldsymbol{0}&\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}^{\top}\widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}\end{pmatrix}\begin{pmatrix}\boldsymbol{c}_{1}\\ \boldsymbol{c}_{2}\end{pmatrix}=\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{Y}}\\ \widetilde{\operatorname{\boldsymbol{L}}}_{\operatorname{\boldsymbol{C}}}^{\top}\operatorname{\boldsymbol{Y}}\end{pmatrix}$.
Here $(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})$ is an $n\times n$ idempotent matrix. Therefore the solution are
[…]
and
$\widehat{\boldsymbol{\gamma}}=(\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{L}}^{\top}(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{C}})^{+}\operatorname{\boldsymbol{C}}^{\top}\operatorname{\boldsymbol{L}}^{\top}(\boldsymbol{I}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{B}}^{\top}\operatorname{\boldsymbol{L}}^{\top})\operatorname{\boldsymbol{Y}}.$
Proof of Theorem 3.2:
Proof. The proof mimics the proof of Theorem 4.2 in Singh et al. (2025) and is therefore omitted.
Proof of Theorem 3.3:
Proof. A bit of algebra shows that
$(\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}-\widehat{\operatorname{\boldsymbol{\nu}}}_{s})=(\boldsymbol{I}-(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}(\operatorname{\boldsymbol{D}}/n)\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}(\operatorname{\boldsymbol{D}}/n)\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}(\operatorname{\boldsymbol{D}}/n)\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}(\operatorname{\boldsymbol{D}}/n)\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\end{pmatrix}(\operatorname{\boldsymbol{D}}/n))\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}.$
By the CLT we have $\operatorname{\boldsymbol{D}}^{1/2}(\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}-{\operatorname{\boldsymbol{\nu}}}_{s})\Rightarrow\mathcal{N}(\operatorname{\boldsymbol{0}},\sigma^{2}\boldsymbol{H}_{s})$, where $\boldsymbol{H}_{s}$ is a $|\operatorname{\boldsymbol{\nu}}|\times|\operatorname{\boldsymbol{\nu}}|$ diagonal matrix whose diagonal elements are $1$ if $(i,j)\in\mathcal{E}_{n}$ and 0 otherwise. It follows by Slutzky’s theorem that $\boldsymbol{U}_{n,\mathcal{S}}=\operatorname{\boldsymbol{D}}^{1/2}(\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}-\widehat{\operatorname{\boldsymbol{\nu}}}_{s})\Rightarrow\boldsymbol{U}_{\mathcal{S}}$ where $\boldsymbol{U}_{\mathcal{S}}$ follows a $\mathcal{N}_{|\operatorname{\boldsymbol{\nu}}|}(0,\boldsymbol{\Psi}_{\mathcal{S}})$ and ${\boldsymbol{\Psi}}_{\mathcal{S}}=\sigma^{2}\,\boldsymbol{\Xi}^{1/2}\boldsymbol{M}\,\boldsymbol{\Xi}^{+}\boldsymbol{M}^{\top}\boldsymbol{\Xi}^{1/2}$, where
$\boldsymbol{M}=\boldsymbol{I}-(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}&\operatorname{\boldsymbol{C}}_{s}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{C}}_{s}\end{pmatrix}^{+}\begin{pmatrix}\operatorname{\boldsymbol{B}}^{\top}\\ \operatorname{\boldsymbol{C}}_{s}^{\top}\end{pmatrix}\boldsymbol{\Xi}.$
[…]
$\boldsymbol{U}_{\mathcal{S}}^{\top}\boldsymbol{U}_{\mathcal{S}}=\sum_{i=1}^{{\rm rank}({\boldsymbol{\Psi}}_{\mathcal{S}})}{\lambda}_{i}Z_{i}^{2},$
where $Z_{1},\ldots,Z_{t}$ are independent $\mathcal{N}(0,1)$ RVs and ${\lambda}_{1},\ldots,{\lambda}_{t}$ are the non–zero eigenvalues of the matrix ${\boldsymbol{\Psi}}_{\mathcal{S}}$ which by Remark A.2 satisfies $t={\rm rank}({\boldsymbol{\Psi}}_{\mathcal{S}})=\binom{K}{2}-(K+r-1)$.
Remark A.2
Algebraically it is difficult to find rank of matrix $\boldsymbol{\Psi}_{\mathcal{S}}$. However, the rank of $\boldsymbol{\Psi}_{\mathcal{S}}$ can be computed using a statistical argument. To do so suppose that the errors $\epsilon_{ijk}$ are IID $\mathcal{N}(0,\sigma^{2})$ RVs. It is well known that the LRT for testing $H_{0}:\operatorname{\boldsymbol{\nu}}\in\mathcal{L}\cup\mathcal{C}_{\mathcal{S}}$ versus $H_{1}:\operatorname{\boldsymbol{\nu}}\notin\mathcal{L}\cup\mathcal{C}_{\mathcal{S}}$ follows, under the null a chi–square distribution with $|\mathcal{E}_{n}|-(K+r-1)$ degrees of freedom. Furthermore, the corresponding Wald test is of the form $\boldsymbol{U}_{n}^{\top}{\boldsymbol{\Psi}}_{n}^{+}\boldsymbol{U}_{n}$ and has the same asymptotic distribution as the LRT and consequently by Brown and Cribari-Neto (1992) the matrix ${\boldsymbol{\Psi}}_{n}$ must have rank $|\mathcal{E}_{n}|-(K+r-1)$ for all large $n$ and therefore also in the limit. This established the claim above.
Next, the non–null distribution of $R_{n}$ under local alternatives is given.
Proposition A.1
Assume that $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$ and the condition of Theorem 3.2 hold. Suppose $\operatorname{\boldsymbol{\nu}}_{n}=\operatorname{\boldsymbol{\nu}}+n^{-1/2}\boldsymbol{\delta}$ where $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$ and $\boldsymbol{\delta}\notin\mathcal{S}$ is a fixed vector. If $n\rightarrow\infty$ then
$R_{n}\Rightarrow\sum_{i=1}^{t}\lambda_{i}(Z_{i}+\gamma_{i})^{2},$
where $\lambda_{1},\ldots,\lambda_{t}$ and $Z_{1},\ldots,Z_{t}$ are as in the statement of Theorem 3.3 and $\gamma_{1},\ldots,\gamma_{t}$ are the elements of the vector $\boldsymbol{\gamma}=\boldsymbol{O}(\boldsymbol{\Psi}_{\mathcal{S}}^{+})^{1/2}\boldsymbol{\Xi}^{1/2}\boldsymbol{\delta}$ which correspond to the nonzero eigenvalues of $\boldsymbol{\Psi}$. Here $\boldsymbol{O}$ is the orthonormal matrix whose columns are the eigenvectors of $\boldsymbol{\Psi}_{\mathcal{S}}$.
Proof. Under the stated conditions, the $(i,j)^{th}$ entry of $\boldsymbol{U}_{n,\mathcal{S}}$ can be written as
$U_{n,\mathcal{S},ij}=\sqrt{n_{ij}}(\overline{S}_{ij}-\widehat{\nu}_{\mathcal{S},ij})=\sqrt{n_{ij}}(\overline{S}_{ij}-(\widehat{\nu}_{\mathcal{S},ij}-n^{-1/2}{\delta}_{ij})+n^{-1/2}{\delta}_{ij})=U_{ij}^{*}+\sqrt{\frac{n_{ij}}{n}}\,{\delta}_{ij},$
where $U_{n,ij}^{*}=\sqrt{n_{ij}}\,(\overline{S}_{ij}-(\widehat{\nu}_{\mathcal{S},ij}-n^{-1/2}{\delta}_{ij}))$. Consequently $\boldsymbol{U}_{n,\mathcal{S}}\Rightarrow\boldsymbol{U}_{\mathcal{S}}$ where $\boldsymbol{U}_{\mathcal{S}}$ is a $\mathcal{N}_{|\operatorname{\boldsymbol{\nu}}|}(\boldsymbol{\Xi}^{1/2}\,\boldsymbol{\delta},\boldsymbol{\Psi}_{\mathcal{S}})$ RV. It follows that
[…]
$=$ $~\boldsymbol{V}_{\mathcal{S}}^{\top}\boldsymbol{\Lambda}\boldsymbol{V}_{\mathcal{S}},$
[…]
$R_{n}=\sum_{i=1}^{t}\lambda_{i}(Z_{i}+\gamma_{i})^{2},$
where $\gamma_{i}$ is the element of $\boldsymbol{\gamma}$ corresponding with the $i^{th}$ nonzero diagonal element of $\boldsymbol{\Lambda}$ denoted by $\lambda_{i}$.
Proof of Proposition 3.1:
Proof. If $n_{ij}\to\infty$ and errors are IID with zero mean and finite variance $\sigma^{2}$ then
$\sqrt{n_{ij}}\,(\widehat{\nu}_{ij}-{\nu}_{ij})\Rightarrow\mathcal{N}(0,\sigma^{2}).$
Therefore, under $H_{0}$, asymptotic normality of $T_{n}$ is immediate. Next note that $\nu_{ij}+\nu_{jk}+\nu_{ki}=(\operatorname{\boldsymbol{C}}_{ij}+\operatorname{\boldsymbol{C}}_{jk}-\operatorname{\boldsymbol{C}}_{ik})\boldsymbol{\gamma}$, so under $H_{1}$, $T_{n}$ converges in probability to $\delta$ as $n_{ij},n_{jk},n_{ik}\to\infty$. Consequently using similar algebra again under $H_{1}$ we get $T_{n}\Rightarrow\mathcal{N}(\delta,\sigma^{2})$ as $n_{ij},n_{jk},n_{ik}\to\infty$.
Proof of Theorem 3.4:
We start with an auxiliary Lemma.
Lemma A.1
Let $\mathcal{S}=\rm{span}(\operatorname{\boldsymbol{B}},\operatorname{\boldsymbol{C}}_{s})$ for some matrix $\operatorname{\boldsymbol{C}}_{s}$ satisfying $\rm{span}(\operatorname{\boldsymbol{C}}_{s})\subset\rm{span}(\operatorname{\boldsymbol{C}})$. Consider testing $H_{0}:\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$ versus $H_{1}:\operatorname{\boldsymbol{\nu}}\notin\mathcal{S}$ using the statistic $R_{n,\mathcal{S}}$ defined in Theorem 3.3. Assuming $\operatorname{\boldsymbol{\nu}}\notin\mathcal{S}$, $n_{ij}\rightarrow\infty$ for all $1\leq i<j\leq K$ while $n_{ij}=O(n)$, we then have
$\mathbb{P}(R_{n,\mathcal{S}}<c_{\alpha})\rightarrow 0$ (30)
where $c_{\alpha}=\mathbb{P}_{H_{0}}(R_{n,\mathcal{S}}\geq c_{\alpha})$ is the $\alpha$ level critical value for $R_{n,\mathcal{S}}$. If, in addition, the errors are subgaussian then for all large $n$ there exists constants $C_{1},C_{2}>0$ such that
$\mathbb{P}(R_{n,\mathcal{S}}<c_{\alpha})\leq C_{1}\exp(-nC_{2}).$ (31)
Remark A.3
Equation (30) shows that under the alternative the power of tests for lack of fit approaches unity as the sample size increases to $\infty$. In fact, this assertion has already been established see the text following the statement of Theorem 3.3. Therefore we only prove (31).
Proof. Observe that
$R_{n,\mathcal{S}}=\sum_{1\leq i<j\leq K}{U}_{n,\mathcal{S},ij}^{2}=\sum_{1\leq i<j\leq K}n_{ij}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})^{2},$ (32)
where $\overline{S}_{{\rm ALL},ij}$ and $\widehat{\nu}_{s,ij}$ are $ij^{th}$ elements of $\overline{\operatorname{\boldsymbol{S}}}_{\rm ALL}$ and $\widehat{\operatorname{\boldsymbol{\nu}}}_{s}$, respectively. Notice that both $\overline{S}_{{\rm ALL},ij}$ and $\widehat{\nu}_{s,ij}$ are linear functions of the errors. Therefore $(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})$ is also a subgaussian random variable with mean $\delta_{ij}$ and parameter $\tau_{ij}^{2}\leq\infty$. Note that under the null, i.e., when $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}$, we have $\delta_{ij}=\mathbb{E}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})=0$ for all pairs $(i,j)$. However, under the alternative $\delta_{ij}\neq 0$ for some pairs $(i,j)$.
Continuing,
$\mathbb{P}(|\sqrt{n_{ij}}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})-\sqrt{n_{ij}}\delta_{ij}|\geq\epsilon)\leq 2\exp(-\epsilon^{2}n_{ij}/2\tau_{ij}^{2})$
In other words
$\mathbb{P}(\sqrt{n_{ij}}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})\notin(\sqrt{n_{ij}}\delta_{ij}-\epsilon,\sqrt{n_{ij}}\delta_{ij}+\epsilon))\leq 2\exp(-\epsilon^{2}n_{ij}/2\tau_{ij}^{2}),$
so, for all large $n$
$\mathbb{P}({n_{ij}}(\overline{S}_{{\rm ALL},ij}-\widehat{\nu}_{s,ij})^{2}\notin({n_{ij}}\delta_{ij}^{2}-\epsilon_{ij},{n_{ij}}\delta_{ij}^{2}+\epsilon_{ij}))\leq 2\exp(-\epsilon^{2}n_{ij}/2\tau_{ij}^{2})$
for some $\epsilon_{ij}>0$ where $\epsilon_{ij}\rightarrow 0$ when $\epsilon\rightarrow 0$. Next observe that
[…]
where $\widetilde{\epsilon}=\sum_{1\leq i<j\leq K}\epsilon_{ij}$. It follows that for all large $n$
$)\displaystyle~\mathbb{P}(R_{n,\mathcal{S}}\notin(\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}-\widetilde{\epsilon},\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}+\widetilde{\epsilon}$))
$\leq$ $~\mathbb{P}(\cup_{1\leq i<j\leq K}\{{U}_{n,\mathcal{S},ij}^{2}\notin(n_{ij}\delta_{ij}^{2}-\epsilon_{ij},n_{ij}\delta_{ij}^{2}+\epsilon_{ij})\})\leq\sum_{1\leq i<j\leq K}2\exp(-\epsilon^{2}n_{ij}/2\tau_{ij}^{2})$
$\leq$ $~C_{1}\exp(-n\epsilon^{2}C_{2}),$
for $C_{1}=2\binom{K}{2}$ and $C_{2}=\min\{n_{ij}/(2n\tau_{ij}^{2}):1\leq i<j\leq K\}$. Therefore
$~\mathbb{P}(R_{n,\mathcal{S}}\in(\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}-\widetilde{\epsilon},\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}+\widetilde{\epsilon}))\geq 1-C_{1}\exp(-n\epsilon^{2}C_{2}).$
However for any $\epsilon$ and all large $n$, $\sum_{1\leq i<j\leq K}n_{ij}\delta_{ij}^{2}-\widetilde{\epsilon}>c_{\alpha}$ so for all large $n$
$\mathbb{P}(R_{n,\mathcal{S}}\geq c_{\alpha})\geq 1-C_{1}\exp(-nC_{2}).$
Taking complements completes the proof.
Remark A.4
Lemma A.1 shows that under subgaussian errors the probability of not rejecting the null when applying $\rm FTBS$} it is false is exponentially small.
We now continue with the proof of Theorem 3.4.
Proof. FTBS is applied by $(i)$ testing $\binom{K}{2}$ hypotheses $H_{0}:\nu_{{\rm cyclic},ij}=0$, $1\leq i<j\leq K$} and generating an estimated tick–table $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cyclic}^{*}}$; and $(ii)$ then sequentially testing for lack of fit using Theorem 3.3. At most $3$ such tests are applied. The set of selected triads obtained after applying $(i)$ and $(ii)$ above is denoted by $\widehat{\mathcal{S}}$.
By Theorem 2.2 there exists a minimal model $\mathcal{S}$ in which all triads have at least one tick in $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cycllic}}$. Further note that $\mathcal{S}\subseteq\mathcal{S}_{b}$ for some $b\in\{1,\ldots,4\}$. Now, if there is no cyclicality in $\mathcal{S}$ then $\mathcal{S}\subseteq\mathcal{S}_{1}$; if all triads in $\mathcal{S}$ have three ticks then $\mathcal{S}\subseteq\mathcal{S}_{2}$; if some triads in $\mathcal{S}$ have two ticks then $\mathcal{S}\subseteq\mathcal{S}_{3}$; and if some triads in $\mathcal{S}$ have one tick then $\mathcal{S}\subseteq\mathcal{S}_{4}$. Thus if $\operatorname{\boldsymbol{C}}_{\operatorname{\boldsymbol{\nu}}_{\rm cycllic}}$ was known then applying Theorem 3.3 would guarantee that FTBS would select $\widehat{\mathcal{S}}\supset\mathcal{S}$.
[…]
$)=\displaystyle\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\notin{\rm span}(\widehat{\mathcal{S}}$))= $~\mathbb{P}(\{{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}}\}\cup_{\{a:\,\mathcal{S}_{b}\not\subset\mathcal{S}_{a}^{*}\}}(\{R_{n,\mathcal{S}_{a}}<c_{\alpha,a}\}\cap\{{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}}\})))$
$\leq$ $~\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})+\sum_{\{a:\,\mathcal{S}_{b}\not\subset\mathcal{S}_{a}^{*}\}}\mathbb{P}(\{R_{n,\mathcal{S}_{a}^{*}}<c_{\alpha,a}\}).$ (33)
By Lemma A.1 if $\mathcal{S}_{b}\not\subset\mathcal{S}_{a}^{*}$ then with high probability we will reject the hypothesis that $\mathcal{S}\subset\mathcal{S}_{a}^{*}$. Therefore the sum in (33) is over set $\{a:\mathcal{S}_{b}\not\subset\mathcal{S}_{a}^{*}\}$.
Next we bound the individual probabilities in (33). First note that if $\nu_{{\rm cyclic},ij}\neq 0$ and the procedure described in Remark 3.1 is followed then $\mathbb{P}(\phi_{ij}(\widehat{\operatorname{\boldsymbol{\nu}}}_{{\rm cyclic}}=1)\rightarrow 1$, i.e., the null $H_{0}:\nu_{{\rm cyclic},ij}=0$ is rejected with probability tending to one. Since $K$ is finite all elements of $\operatorname{\boldsymbol{\nu}}_{\rm cyclic}$ which are different than 0 will be identified with probability tending to one. Furthermore approximately $\alpha(\binom{K}{2}-\sum_{1\leq i<j\leq K}\mathbb{I}({\nu_{{\rm cyclic},ij}}\neq 0))$ other elements of $\operatorname{\boldsymbol{\nu}}$ will be falsely flagged as non–zero. We have just proved that
$\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})\to 0.$ (34)
Consequently $\mathcal{S}_{a}\subset\mathcal{S}_{a}^{*}$ for $a\in\{1,\ldots,4\}$ with probability tending to $1$ as $n\to\infty$. Next using Lemma A.1 for $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}_{b}$ and $b>a$, we have
$\mathbb{P}(\{R_{n,\mathcal{S}_{a}^{*}}<c_{\alpha,a}\})\to 0.$ (35)
Therefore (34) and (35) imply that
$\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\notin{\rm span}(\widehat{\mathcal{S}}))\to 0,$
establishing the first stated claim.
[…]
$\mathbb{P}(|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\geq\epsilon)\leq 2e^{-\epsilon^{2}/({\tau}_{ij}^{2}/n)}=2e^{-\epsilon^{2}n/{\tau}_{ij}^{2}}.$ (36)
Next using union bound and (36) we have
$\mathbb{P}(\cup_{1\leq i<j\leq K}\{|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\geq\epsilon\})$ $~\leq\sum_{1\leq i<j\leq K}\mathbb{P}(\{|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\geq\epsilon\})$
$~\leq\sum_{1\leq i<j\leq K}2e^{-\epsilon^{2}n/{\tau}_{ij}^{2}}.$
Let ${\tau}_{{\rm cyclic}}^{2}=\min\{{\tau}_{ij}^{2}:1\leq i<j\leq K\}$. Then
$\mathbb{P}(\cup_{1\leq i<j\leq K}\{|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\leq\epsilon\})\geq 1-\sum_{1\leq i<j\leq K}2e^{-\epsilon^{2}n/{\tau}_{ij}^{2}}\geq 1-\binom{K}{2}2e^{-\epsilon^{2}n/{\tau}_{{\rm cyclic}}^{2}}.$
Since $K$ is finite, there exist $\bar{C}_{1},\bar{C}_{2}>0$ such that
$\mathbb{P}(\cup_{1\leq i<j\leq K}\{|\widehat{\nu}_{{\rm cyclic},ij}-\nu_{{\rm cyclic},ij}|\geq\epsilon\})\leq\bar{C}_{1}e^{(-n\bar{C}_{2})}.$
Therefore not rejecting $H_{0}:\nu_{{\rm cyclic},ij}=0$ when $H_{0}$ is false has probability
$\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})\leq\bar{C}_{1}e^{(-n\bar{C}_{2})}.$ (37)
Next using Lemma A.1 for $\operatorname{\boldsymbol{\nu}}\in\mathcal{S}_{b}$ and $b>a$, we have
$\mathbb{P}(\{R_{n,\mathcal{S}_{a}^{*}}<c_{\alpha,a}\})\leq C_{a1}e^{(-nC_{a2})}.$ (38)
Thus (33), (37) and (38) imply that
$)\displaystyle\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\notin{\rm span}(\widehat{\mathcal{S}}$$)\leq$ $~\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}\not\subset{\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})+\sum_{1\leq a<b}\mathbb{P}(\{R_{n,\mathcal{S}_{a}^{*}}<c_{\alpha,a}\})$
$\leq$ $~\bar{C}_{1}e^{(-n\bar{C}_{2})}+\sum_{1\leq a<b}C_{a1}e^{(-nC_{a2})}\leq C_{1}e^{(-nC_{2})}$
for some finite $C_{1},C_{2}>0$. Therefore there exists $C_{1},C_{2}>0$ such that
$\mathbb{P}(\operatorname{\boldsymbol{\nu}}_{\rm cyclic}\in{\rm span}(\widehat{\mathcal{S}}))\geq 1-C_{1}e^{(-nC_{2})}.$
Thus establishing the second claim.
By Theorem 2.2 if triads in the minimal model share at most index then the minimal model is unique, moreover only triads which are part of the minimal model will have three ticks. Condition (26) guarantees that $H_{0}:{\nu}_{{\rm cyclic},ij}=0$ is rejected with probability tending to one if and only if ${\nu}_{{\rm cyclic},ij}\neq 0$. Thus $\mathbb{P}({\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}}={\operatorname{\boldsymbol{C}}}_{{\operatorname{\boldsymbol{\nu}}}_{\rm cyclic}^{*}})\to 1$. Further note that under the stated condition $\mathcal{S}=\mathcal{S}_{2}$ and $\mathcal{S}_{2}=\mathcal{S}_{2}^{*}$ with probability tending to $1$. Since the hypotheses in Step 2 of FTBS are tested at level $\alpha_{n}\rightarrow 0$ the algorithm FTBS will terminate at $a=2$ with probability tending to one. In other words $\mathbb{P}(\mathcal{S}=\widehat{\mathcal{S}})\to 1$, concluding the proof.
Appendix B The relationship between the reduced and full model
Appendix B explore the relationship between the full and reduced model. In particular we investigate what occurs when model (1) is fit under the assumption that $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$, in circumstances where this assumption may not hold.
Theorem B.1
Assume the errors are IID $\mathcal{N}(0,\sigma^{2})$ RVs. Let $\mathrm{KL}(\mathcal{N},\mathcal{L})$ denote the Kullback–Leibler divergence between the distribution of $\operatorname{\boldsymbol{Y}}$ indexed by $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ and $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$ respectively. Then:
${\operatorname{\boldsymbol{\mu}}}=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\mathrm{KL}(\mathcal{N},\mathcal{L})=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\sum_{1\leq i<j\leq K}n_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}.$
Moreover, if there exists a spanning tree $\mathcal{T}\subset\mathcal{G}$ such that $\min\{n_{ij}:\left(i,j\right)\in\mathcal{T}\}\rightarrow\infty$ as $n\rightarrow\infty$, then the LSE of the reduced model, given by $\widehat{\operatorname{\boldsymbol{\mu}}}=\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{S}}$, will converge to
$\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\{\sum_{1\leq i<j\leq K}\theta_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}:\boldsymbol{v}^{\top}\operatorname{\boldsymbol{\mu}}=0\}.$ (39)
The minimiser of (39) is unique if there exists a spanning tree $\mathcal{T}\subset\mathcal{G}$ such that
$\min\{n_{ij}:(i,j)\in\mathcal{T}\}/n\to c\in(0,\infty)\text{ as }n\to\infty.$
Furthermore, the solution of (39) is independent of $\boldsymbol{\Theta}$ if and only if $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$. Finally if the constraint $\boldsymbol{v}=\operatorname{\boldsymbol{1}}1$ is imposed then the linear relation
$\operatorname{\boldsymbol{\mu}}=(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}.$ (40)
holds.
Proof. By assumption $\operatorname{\boldsymbol{Y}}\sim{\mathcal{N}}_{n}(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}},\sigma^{2}\boldsymbol{I})$, where matrix $\operatorname{\boldsymbol{L}}$ is defined in Section 3.2. Moreover if $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$ then $\nu_{ij}=\mu_{i}-\mu_{j}$ and for brevity will be denoted it by $\operatorname{\boldsymbol{\nu}}_{\mu}$. Let $\boldsymbol{f}_{1}$ and $\boldsymbol{f}_{2}$ be the densities corresponding to $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ and $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$, respectively. The Kullback–Leibler divergence between models indexed by $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$ and $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$ is
$\mathrm{KL}(\mathcal{N},\mathcal{L})=$ $~\mathbb{E}_{\boldsymbol{f}_{1}}(\log{\boldsymbol{f}_{1}}-\log{\boldsymbol{f}_{2}})=\frac{1}{2\sigma^{2}}\mathbb{E}_{\boldsymbol{f}_{1}}(-(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}})^{\top}(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}})+(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu})^{\top}(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu}))$
$=$ $~\frac{1}{2\sigma^{2}}\mathbb{E}_{\boldsymbol{f}_{1}}(2(\operatorname{\boldsymbol{Y}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}})^{\top}(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu})+(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu})^{\top}(\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{L}}\operatorname{\boldsymbol{\nu}}_{\mu}))$
$=$ $~\frac{1}{2\sigma^{2}}(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{\nu}}_{\mu})^{\top}\operatorname{\boldsymbol{L}}^{\top}\operatorname{\boldsymbol{L}}(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{\nu}}_{\mu})=\frac{1}{2\sigma^{2}}\sum_{1\leq i<j\leq K}n_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}.$
It follows that $\mathrm{KL}(\mathcal{N},\mathcal{L})$ is minimized by
${\operatorname{\boldsymbol{\mu}}}=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\mathrm{KL}(\mathcal{N},\mathcal{L})=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\sum_{1\leq i<j\leq K}n_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}.$
Next observe that
$\widehat{\operatorname{\boldsymbol{\mu}}}$= $~\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\frac{1}{n}\sum_{1\leq i<j\leq K}\sum_{t=1}^{n_{ij}}(Y_{ijt}-\left(\mu_{i}-\mu_{j}\right))^{2}=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\frac{1}{n}\sum_{1\leq i<j\leq K}\sum_{t=1}^{n_{ij}}(\nu_{ij}-(\mu_{i}-\mu_{j})+\epsilon_{ijt})^{2}$
$=$ $~\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\sum_{1\leq i<j\leq K}\frac{n_{ij}}{n}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}-2\sum_{1\leq i<j\leq K}\frac{n_{ij}}{n}(\nu_{ij}-(\mu_{i}-\mu_{j}))\frac{1}{n_{ij}}\sum_{t=1}^{n_{ij}}\epsilon_{ijt}$
$~+\frac{1}{n}\sum_{1\leq i<j\leq K}\sum_{t=1}^{n_{ij}}\epsilon_{ijt}^{2}.$
By the law of large numbers the third term above converges to $\sigma^{2}$ which is independent of $\operatorname{\boldsymbol{\mu}}$. By assumption $n_{ij}/n=\theta_{ij}+o(1)$ and consequently we can replace $n_{ij}/n$ appearing in the first term above with $\theta_{ij}+o(1)$. If $\theta_{ij}=0$ then the second term is $o_{p}(1)$ whereas if $\theta_{ij}>0$ then ${1}/{n_{ij}}\sum_{t=1}^{n_{ij}}\epsilon_{ijt}\to 0$; hence in both cases the second term is $o_{p}(1)$. Putting it all together we find that
$\widehat{\operatorname{\boldsymbol{\mu}}}=\arg\min_{\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}}\sum_{1\leq i<j\leq K}\theta_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}+o_{p}(1).$
establishing (39). Further note that
$\sum_{1\leq i<j\leq K}\theta_{ij}(\nu_{ij}-(\mu_{i}-\mu_{j}))^{2}=(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}})^{\top}\boldsymbol{\Xi}(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}).$ (41)
By assumption there exist a tree $\mathcal{T}$ such that $\theta_{ij}>0$ for all $(i,j)\in\mathcal{T}$; clearly a unique minimizer exist on $\mathcal{T}$ and consequently on every $\mathcal{G}\supset\mathcal{T}$. For a finite sample analogue see Theorem 2.1 in Singh et al. (2024). Further note that if $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$ then (39) is minimized by choosing $\operatorname{\boldsymbol{\mu}}$ such that $\nu_{ij}=\mu_{i}-\mu_{j}$ in which case (41) is identically zero for all possible choices of $\boldsymbol{\Theta}$. If however $\operatorname{\boldsymbol{\nu}}\notin\mathcal{L}$ then (41) is always a function of $\boldsymbol{\Theta}$. The Larnagian associated with (39) is
$L=(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}})^{\top}\boldsymbol{\Xi}(\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}})+\lambda\boldsymbol{v}^{\top}\operatorname{\boldsymbol{\mu}},$
where $\operatorname{\boldsymbol{1}}1^{\top}\boldsymbol{v}\neq 0$. Note that gradient of $L$ is $\nabla L=(\nabla_{\operatorname{\boldsymbol{\mu}}}L,\nabla_{\lambda}L)$ where
$\nabla_{\operatorname{\boldsymbol{\mu}}}L=-\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}+\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}+\lambda\boldsymbol{v}/2$ and $\nabla_{\lambda}L=\boldsymbol{v}^{\top}\operatorname{\boldsymbol{\mu}}$. Following the proof of Theorem 2.1 in Singh et al. (2025) it can be shown that
$\widehat{\operatorname{\boldsymbol{\mu}}}=(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}+[\boldsymbol{I}-(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}}]\boldsymbol{\alpha},$
where $\boldsymbol{\alpha}\in\mathbb{R}^{K}$ is arbitrary which reduces to
$\widehat{\operatorname{\boldsymbol{\mu}}}=(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}$ (42)
when $\boldsymbol{v}=\operatorname{\boldsymbol{1}}1$.
Proposition B.1 explores the relationship between $\operatorname{\boldsymbol{\mu}}\in\mathbb{R}^{K}$ and $\operatorname{\boldsymbol{\nu}}\in\mathbb{R}^{K(K-1)/2}$. One of its consequences is that the LSE, assuming the reduce model, given by $\widehat{\operatorname{\boldsymbol{\mu}}}=\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{S}}$, will converge to the LHS of (39). Obviously this limit depends very strongly on the limiting value of the Laplacian, as illustrated by the following example.
Example B.1
Let $K=4$ and suppose that $\operatorname{\boldsymbol{\nu}}=(-2,0,2,0,2,0)$. It is easy to check that $\operatorname{\boldsymbol{\nu}}\notin\mathcal{L}$. By Theorem B.1 the LSE of $\operatorname{\boldsymbol{\mu}}$, will converge to the RHS of (40). If the comparison graph is complete and balanced, i.e., $n_{ij}=m$ for $1\leq i<j\leq K$, then $\boldsymbol{\Xi}=diag(1/6,1/6,1/6,1/6,1/6,1/6)$; whereas if it is a balanced path graph, i.e., $n_{12}=n_{23}=n_{34}=m$, then $\boldsymbol{\Xi}=diag(1/3,0,0,1/3,0,1/3)$. In both cases $\operatorname{\boldsymbol{B}}$ is given by
$\boldsymbol{B}^{\top}=\begin{pmatrix}1&1&1&0&0&0\\ -1&0&0&1&1&0\\ 0&-1&0&-1&0&1\\ 0&0&-1&0&-1&-1\end{pmatrix}$.
Using (40) we find that the the LSE will converge to (0,1,0,1)(0,1,0,-1) for the complete graph and to (1.5,0.5,0.5,0.5)(-1.5,0.5,0.5,0.5) for the path graph. The resulting merits are obviously very different and topology dependent.
It is worth noting that if we substitute the sample versions of $\boldsymbol{\Xi}$ and $\operatorname{\boldsymbol{\nu}}$ on the RHS of (40) then its LHS, i.e., $\operatorname{\boldsymbol{\mu}}$, coincides with the LSE $\widehat{\operatorname{\boldsymbol{\mu}}}=\operatorname{\boldsymbol{N}}^{+}\operatorname{\boldsymbol{S}}$. Therefore (40) provides an alternative way of computing the LSE. Further note that using (40) one can find the value of $\operatorname{\boldsymbol{\nu}}\in\mathcal{L}$, denoted by $\operatorname{\boldsymbol{\nu}}_{\operatorname{\boldsymbol{\mu}}}$, providing the best fit for the model when $\operatorname{\boldsymbol{\nu}}\in\mathcal{N}$. In fact
$\operatorname{\boldsymbol{\nu}}_{\operatorname{\boldsymbol{\mu}}}=\operatorname{\boldsymbol{B}}\operatorname{\boldsymbol{\mu}}=\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{\nu}}$
so $\|\operatorname{\boldsymbol{\nu}}-\operatorname{\boldsymbol{\nu}}_{\operatorname{\boldsymbol{\mu}}}\|\leq\|\boldsymbol{I}-\operatorname{\boldsymbol{B}}(\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\operatorname{\boldsymbol{B}})^{+}\operatorname{\boldsymbol{B}}^{\top}\boldsymbol{\Xi}\|_{F}\|\operatorname{\boldsymbol{\nu}}\|$, where $\|\cdot\|_{F}$ denotes the Frobenius norm. It should also be emphasized that $\operatorname{\boldsymbol{\nu}}_{\operatorname{\boldsymbol{\mu}}}$ is a projection onto $\mathcal{L}$ if and only if $\boldsymbol{\Xi}=a\boldsymbol{I}$ for some $a\in\mathbb{R}$, i.e., if and only if the comparison graph is balanced in which case $n_{ij}=m$ for all pairs $(i,j)$.
Appendix C EPL rank sets
The following table provides the rank sets for the teams participated in EPL season 2022-23.
Table 7: Rank sets
Team $S_{i}$ $I_{i}$ $E_{i}$
1 12,13 2,3,4,5,6,7,8,9,10,11,14,15,16,17,18,19,20
2 1,6,7,11,12,13,14,15,19 3,4,5,8,10,16,17,18,20 9
3 1,2,4,5,6,7,10,12,13,14,15,16,18,19,20 8,9,11,17
4 1,2,5,11,12,13 3,6,7,8,9,10,14,15,16,17,18,19,20
5 1,2,13,15 3,4,6,7,8,9,10,11,12,14,16,17,18,19,20
6 1,4,5,10,11,13,14,15,16,19 2,3,7,8,9,12,17,18,20
7 1,4,5,6,8,9,12,13,14,15,17,18,20 2,3,10,11,16,19
8 1,2,3,4,5,6,9,11,12,13,14,15,18,19 7,10,16,17,20
9 1,3,4,5,6,11,12,13,14,15,18,19 7,8,10,16,17,20 2
10 1,2,4,5,7,8,9,11,12,13,14,15,18,19,20 3,6,16,17
11 1,3,5,7,12,14,15,16,17 2,4,6,8,9,10,13,19,20 18
12 5,6,13 1,2,3,4,7,8,9,10,11,14,15,16,17,18,19,20
13 11 1,2,3,4,5,6,7,8,9,10,12,14,15,16,17,18,19,20
14 1,4,5,12,13,15,17 2,3,6,7,8,9,10,11,16,18,20 19
15 1,4,12,13 2,3,5,6,7,8,9,10,11,14,16,17,18,19,20
16 1,2,4,5,7,8,9,10,12,13,14,15,17,18,19 3,6,11,20
17 1,2,3,4,5,6,8,9,10,12,13,15,18,19 7,11,14,16,20
18 1,2,4,5,6,12,13,14,15 1,2,4,5,6,12,13,14,15 11
19 1,4,5,7,11,12,13,15,18 2,3,6,8,9,10,16,17,20 14
20 1,2,4,5,6,8,9,11,12,13,14,15,16,17,18,19 3,7,10